### v1

# Weak Memory Model Formalisms: Introduction and Survey

Abstract.
Memory consistency models define the order in which accesses to shared memory in a concurrent system may be observed to occur. Such models are a necessity since program order is not a reliable indicator of execution order, due to microarchitectural features or compiler transformations. Concurrent programming, already a challenging task, is thus made even harder when weak memory effects must be addressed. A rigorous specification of weak memory models is therefore essential to make this problem tractable for developers of safety- and security-critical, low-level software.
In this paper we survey the field of formalisations of weak memory models, including their specification, their effects on execution, and tools and inference systems for reasoning about code. To assist the discussion we also provide an introduction to two styles of formal representation found commonly in the literature (using a much simplified version of Intel’s x86 as the example): a step-by-step construction of traces of the system (operational semantics); and with respect to relations between memory events (axiomatic semantics). The survey covers some long-standing hardware features that lead to observable weak behaviours, a description of historical developments in practice and in theory, an overview of computability and complexity results, and outlines current and future directions in the field.
1. Introduction
Programmers of low-level concurrent systems code – e.g., operating systems or library data structures – need to address the behaviour of the target architecture; specifically, how communication occurs via accesses to shared memory. This is nontrivial, as architectures (single- or multi-core) do not guarantee to execute instructions in the order specified, and nor do compilers guarantee to maintain the order of the program statements in the generated code. This can add to the difficulties already associated with concurrency under the standard, sequential interpretation.
For example, consider an Arm or RISC-V processor executing assembly corresponding to the following higher-level code, where one process writes data to a shared variable and then signals this change to a concurrent process (or processes) via a shared “” variable. The reading process spins waiting for the flag to be set and then reads the data from into its local state (variable ).
The processor executing the first process can potentially execute the two instructions out of order, since, from the perspective of a single processor, the order is irrelevant to the final values of the two variables. As such, the flag may be set before is written to , and thus the second process may receive an unintended value. In fact, by the same justification, even if the first process performs its instructions in the order specified, the second processor may execute the load of before the loop is resolved (a type of speculative execution), again resulting in an erroneous value for .
To address this extra complication introduced by architectures and compiler optimisations, processor vendors and language designers define a “memory model”, a minimal set of constraints about the behaviour of interacting concurrent processes, typically with respect to the order of accesses of shared memory. If those constraints are too weak to enforce the logic of an algorithm, programmers may inject “fences” or related mechanisms that can restore the standard, sequential behaviour of a (segment of a) program. For instance, the use of release/acquire constraints is a well-known programming pattern for addressing weak memory model effects [127]. A release fence, ‘’, (in C, atomic_thread_fence(memory_order_release)) can be injected into the first process of (1), . In tandem with an acquire fence after the loop of the second process these ensure the intended value for will be read. However, excessive use of such mechanisms comes with performance penalties, which can be a significant issue for low-level systems code [250].
Since the early days of multiprocessors (a system consisting of multiple processors) to the present age of multicore processors (each processor consisting of multiple cores), the literature on weak memory models has developed in academia, the commercial sector, and standardisation committees. The issue affects hardware designers, compiler writers, application programmers, and language designers [17; 16]. As yet, no consensus on a particular formal representation has been settled upon, leaving programmers – in particular those requiring formal assurance for the code – with many choices. The company Arm has led the way in formalising their architecture and robustly investigating their design, including with publicly accessible tool support. Their formal model is validated on hardware through the use of thousands of small programs called “litmus tests”. However, the mechanisms for inferring properties of programs using their formalisation are complex, even for simple programs; as a result, many other representations have been considered.
In this paper, we survey these diverse formalisms and tools that have been proposed for representing weak memory models. A formal specification of a memory model should preferably match the existing intuitions about processor behaviours and compiler optimisations. Furthermore, it should support an easy enumeration of the consequences: a guide that feeds into the software developers’ effort. Though this paper is primarily aimed at researchers with a background in formal methods, we consider the field from the perspective of a practitioner:
How to specify a memory model? (hardware designers; language designers);
How to understand the effect of a memory model on a program? (compiler writers; application programmers); and
How to formally verify that a program is correct under a particular memory model? (application programmers – specifically for secure or safety-critical systems).
Each of these questions has many potential answers, depending on the level of microarchitectural detail vs. abstraction desired, level of tool support, and program properties under question.
Scope and Limitations.
In this survey we attempt to cover the majority of common formalisations of modern hardware and software weak memory models, with reference to historical developments where there are clear impacts on current practice. We are particularly focussed on formalisations that lead to tools and techniques for inferring properties of programs, i.e., from a developers perspective.
We aim to trace the historical development of processor design from the 1960s, through theoretical developments in the 1990s and up to the present day, which includes regular contributions in major publishing venues and refinements of commercial specifications. As a result, we were unable to find a methodical sequence of search terms of the type used in a recent survey on programming language memory models [219]. Hence, we do not claim full coverage, but intend that the range of works surveyed is indicative of the state of the field.
We do not specifically consider issues related to hardware or compiler construction: we assume that microarchitectural implementations of instruction set architectures faithfully match their specifications, and that compiler transformations correctly preserve their intended software memory model in the underlying architecture. Of course, this is not always the case, with problems being discovered in hardware [36; 38] and potential compiler mappings [284].
Outline.
We first introduce terminology and conventions (Section 1.1), and informally introduce litmus tests (Section 1.2), which are small programs designed to concisely show weak memory effects. We then briefly review aspects of processors that give rise to weak memory effects (Section 2), and give an overview of prominent weak memory models, past and present (Section 3).
As an introduction to formalisations and the range of representations that have been considered we present an operational (Section 4) representation based on write buffers (as defined in [261]), and and an axiomatic (Section 5) view that abstracts from architectural details and instead focusses on properties of traces of behaviours (based on work best exemplified in [36]). We use a representative fragment of Intel’s x86 as a running example of a memory model, and use the store buffering litmus test (Example 1.2) as the running example program. Because examples are traditionally the best way to explain weak memory effects, we keep the them separate from the main text, in highlighted boxes, to assist readers looking for instances over details.
In Section 6 we survey and compare operational, axiomatic, and other representations from the literature. Verification methods building on formalisations are surveyed in Section 7, and tool support (and associated algorithmic problems) is surveyed in Section 8. We suggest ways forward for unifying research into and properties of weak memory models in Section 9, and outline how formalisations intersect with other challenges in software verification in Section 10.
1.1. Terminology and Conventions
We clarify some of the terminology and naming conventions used in this paper, as there is significant overloading of terms in the literature. Our survey covers both low-level and high-level, abstract concepts, but we tend to choose the more abstract terminology.
Processes and Variables. A shared-memory concurrent system consists of a set of processes (covering threads and cores) and a shared memory. The shared memory consists of shared variables (covering locations). Each process can access local variables (covering registers), which are not accessible by other processes.
Instructions and Events. Processes execute instructions such as loads and stores. Instructions may generate reads and writes, which are events (or operations) on the shared memory. Some instructions generate more than one event; these instructions are called atomic primitives or read-modify-writes (RMWs), and examples include the “test-and-set” (TAS) and “compare-and-swap” (CAS) instructions.
By convention we let be a value, typically an integer or byte representation. We write to denote the instruction that stores the value to the shared variable . In the literature, such a store may be written as “,” or in an assembler language such as x86 as “MOV [rx] v”, where rx is a register that holds the location in shared memory associated with a program variable . As above we tend to use the more abstract syntax, which may be straightforwardly converted to other styles. (Similarly, denotes a load of shared variable into register , sometimes written or “MOV r [rx]”.)
Sequential Consistency. The standard, intuitive meaning of a concurrent program is a set of interleavings. Each interleaving is a sequence of all operations, such that if operation i is specified in the program to run before operation j, then i will be executed before j. When a concurrent program is viewed this way, we say that it is executed under Sequential Consistency. As we will see in Section 2, this is not generally guaranteed on real processors or by compilers.
Memory Models. The word “model” is heavily overloaded and potentially ambiguous.111 For example, a “memory model” can mean the structure of the memory address space. This usage, unrelated to concurrency, occurs in microarchitectures [133, p.72] as well as in C programming [189; 213]. In the setting of C, a memory model in this sense is sometimes called a “memory object model.” In this paper, a memory model222A memory model is also called memory consistency, consistency model and memory consistency model [220]. denotes the underlying set of possible behaviours of every concurrent program. Sequential Consistency is a memory model, and in fact the most simple one. When a program exhibits behaviour that is not compatible with Sequential Consistency, then the program is said to run under a weak (or relaxed) memory model.
1.2. Litmus Tests
Weak memory models are traditionally understood using small programs that highlight particular unexpected behaviours, typically called litmus tests. Given some initial state of the shared and local variables, one can “probe” a hardware or software memory model by checking whether a final state is reachable for a given litmus test. Examples 1.2–1.2 give several litmus tests, including the store buffering litmus test, which we use as a running example, and is one of the most common litmus tests referred to in the literature [209; 202]. Table 1 summarises the relationships between the key litmus tests and the common memory models.
| SC | x86 | Arm | RISC-V | C | ||
| SB | (Ex.1.2, p.1.2) | |||||
| LB | (Ex.1.2, p.1.2) | |||||
| MP | (Ex.1.2, p.1.2) | |||||
| IRIW | (Ex.2.4, p.2.4) | |||||
| LB+ctrls | (Ex.3.3, p.3.3) |
[t]
2. The Development of Processors
To contextualise how weak memory behaviours arise we sketch some historical developments in processor design. We begin with one aspect of processor microarchitectures which may lead to weak memory behaviours: instruction-level parallelism within the pipeline (Section 2.1). We then discuss fence instructions (and other mechanisms) that can be used to restore order if necessary (Section 2.2). We also outline other microarchitectural features such as forwarding (Section 2.3) and (the lack of) multicopy atomicity (Section 2.4). Milestones in processor design relevant to the development of weak memory models are outlined in Table 2.
| 1963 | The CDC 6600 [278] significantly improved processor speed by using a “scoreboard” to track which registers and locations are being accessed by instructions in the pipeline; instructions that have no dependencies may be parallelised. |
| 1967 | The first implementation of Tomasulo’s algorithm [279] was Model 91 in IBM’s System/360 family [37], which improved the utilisation of multiple arithmetic units. |
| 1980s | The Stanford DASH multiprocessor [188] brought together much existing research and influenced later processors. |
| 1990s | Speculative execution was introduced by Intel’s Pentium family of processors (a type of x86 architecture). |
| 2001 | IBM’s POWER4 was the first commercially available multicore processor. It includes a cache system with even weaker guarantees than those given by pipeline reordering, and is not multicopy-atomic (Section 2.4). |
| 2011 | RISC-V was released as an openly available specification [249], with a memory model based on Release Consistency (see Section 3.1). |
| 2017 | Based on research outcomes, as described in [36], the Arm specification explicitly became multicopy-atomic [240]. |
2.1. Instruction-Level Parallelism
The execution of a single instruction may involve multiple micro-steps, including to access registers, calculate values, or read and write main memory. These latter micro-steps, which we call “memory operations”, can include complex interactions with the cache system, or across PCI devices. Any one of these or many other aspects of computer hardware [147; 230; 212] could potentially influence the actual order of execution of the original sequence of compiled instructions.
Focussing at the local level, the performance of a processor can be improved in part by executing independent micro-steps in parallel. Techniques in this vein are collectively termed instruction-level parallelism, and have been developed since the early days of computing [246]. Straightforward cases include parallelising loads of distinct locations in memory, and parallelising independent calculations to utilise multiple arithmetic and logic units (ALUs) on a single chip. The decision whether or not to parallelise is guided by the observable effect on the sequential semantics – if the programmer could potentially observe the parallelism on a single chip, then it should not be allowed (see Example 2.1). Early examples include the “scoreboarding” technique of the CDC 6600 computer [278] and Tomasulo’s algorithm [279]. Parallelism can also be employed when accessing distinct locations in memory (see Example 2.1).
[t]
As an example of how microarchitectural features can result in weak memory effects we outline below a standard view of processor pipelines. Pipeline processing can be generally considered in three main phases. First, instructions are fetched from the code base into the pipeline. Second, instructions in the pipeline are issued to local ALUs, registers, or the main memory. Third, when all required values have been retrieved and calculated, the instructions are completed and removed from the pipeline. As different instructions may take different lengths of time in the issue and complete phases, they may be parallelised.
[t]
These “reordering” principles were extended by Intel’s Pentium processors in the 1990s to include speculative execution: while waiting for a branch condition to be resolved to true or false (e.g., while waiting for a value to return from memory), a processor may predict the result and begin transiently executing down the corresponding path. This transient execution may include executing memory operations associated with loads (but not stores) out-of-order with instructions before the branch point. If the prediction is later seen to be correct then the effects of the transient execution are committed; otherwise, the transient calculations are discarded and fetching restarts at the alternative (not-predicted) path. Where processor speeds are much faster than memory latency, speculative execution provides a substantial performance gain. Such optimisations have led to significant security vulnerabilities, exposed by attacks such as Spectre [171] and Meltdown [193], which were only discovered decades after speculative execution and other optimisation mechanisms had become standard.
The out-of-order executions outlined above pose no issue on single-core processors, as they have been designed to preserve the sequential semantics of programs. However, in the concurrent setting, these previously unobservable reorderings become potentially visible. The crucial dependencies between instructions that modify or read local or shared variables can be summarised as follows.
Definition 2.1 (Order dependence of variable accesses).
If instruction i is earlier in the pipeline than j,
then j can be completed before i unless one of the following holds.
j writes to a variable written to by i.
j writes to a variable read by i.
j reads a variable written to by i.333This may not always prevent reordering, due to forwarding (Section 2.3).
j reads a shared variable read by i.
These checks are straightforward for a processor by examining the shared and local variables accessed by instructions in the pipeline. The first three constraints are necessary to preserve sequential semantics on a single processor. The last constraint is not necessary for the sequential semantics (and notably does not apply to accessing registers), but is required to ensure that each process observes the updates to a shared variable in a coherent global ordering. For instance, consider the simple case of polling a (shared) variable that changes over time – it is reasonable to assume that reads of in program order are intended to observe changes to chronologically. The four principles above still apply in the presence of fences and speculative execution, though other concerns – such as preventing writes if there is an earlier unresolved branch – also play a role. Despite the difficulties instruction-level parallelism poses to formalising and reasoning about behaviours and potential security vulnerabilities, these optimisations are likely to remain due to their associated computational efficiency.
2.2. Fences and Memory Ordering Constraints
Because memory operations can be reordered due to a range of factors, programmers need mechanisms to enforce the original order of instructions, if the logic of their algorithm so requires. For example, in many data structures with fine-grained concurrency (i.e., those that do not utilise locks for mutual exclusion), the precise ordering of memory operations is key to the correct interaction of competing processes. To this end, every memory model provides fence instructions types (also called barriers), and/or memory ordering constraints (or access modes) that can adorn existing instructions.
Many different types of fences exist; a full fence ensures that all memory operations before it have been issued, and prevents all later memory operations after it from executing early (see Example 2.2: x86 provides mfence, Arm provides dmb, and RISC-V provides fence.rw.rw). Some processors come with more fine-grained fences; such a specialised fence may only affect specific types of memory operations (loads or stores), or may only affect instructions before or after it. Memory ordering constraints may also be used to enforce ordering between otherwise independent instructions; the classic examples are the release and acquire constraints pioneered by Release Consistency [127] (see Example 2.2 and Section 3.1: as an example, Arm provides both LDAR and LDAPR instructions for acquire loads).
[t]
2.3. Forwarding
Forwarding (or bypassing) is a feature of processors where a calculated value in the pipeline can be used by an instruction later in the pipeline. It manifests in a program such as . According the principles of pipeline reorderings (Section 2.1), must wait for to complete, as it is reading a variable being written by an earlier instruction. Since the value being assigned to is already known, processors will typically simplify to to avoid a potentially costly interaction with the main memory or local registers. After the simplification, the two instructions become independent, and hence can be issued and completed in the reverse order while maintaining behaviour allowed by the sequential semantics.
2.4. Multicopy Atomicity
The majority of modern CPU architectures (x86, Arm, and RISC-V) are said to possess the property of multicopy atomicity – every process observes the modifications to each variable in one consistent order. For instance, suppose Process A modifies and then . On a multicopy-atomic system, both Process B and Process C will observe those modifications in that order. In contrast, on a system that lacks multicopy atomicity, Processes B and C may disagree on the order of modifications, i.e., Process B may observe the modification to followed by the modification to , but Process C can simultaneously observe the modification to before the modification to (see Example 2.4).
[ht]
Currently, IBM’s Power series is the only commercially available CPU architecture that lacks multicopy atomicity (the NVIDIA PTX GPU also lacks multicopy atomicity [224]). This is due to its cache coherence system, which effectively allows individual cores to communicate point-to-point via shared caches, before writes are flushed to the shared memory.
The lack of multicopy atomicity creates an extra level of complexity for describing memory models, as well as understanding their effects on code – both at the assembler-level as well as in high-level languages to be compiled to the hardware. To this end, Arm officially stipulates multicopy atomicity for their architectures from Version 8 in 2017 [240]. (In fact, Arm processors effectively have always been multicopy-atomic. Although multicopy atomicity was not explicitly stipulated by earlier versions of Arm, no manufacturer of Arm processors ever produced an implementation that lacked multicopy atomicity [240].) Similarly, RISC-V insists on multicopy atomicity, and Intel’s x86 architectures are also multicopy-atomic.
In practice, the complexity of dealing with an architecture that lacks multicopy atomicity leads programmers to insert enough fences to restore multicopy atomicity (or indeed Sequential Consistency). We therefore do not consider the lack of multicopy atomicity in this paper; a description of the behaviour of Power is given in [122], and the difficulties of verification under Power’s memory model are discussed in [94].
3. The Development of Memory Models
In this section, we sketch the history of how weak memory models have been described and addressed, and mention the development of significant hardware and software memory models. Table 3 summarises the history with selected references. A reader interested purely in formalisms may skip this section.
| SPARC | Power | x86 | Arm | RISC-V | Java | C | |
| 1990s | [263] | [92] | |||||
| 2000s | [227] | [208] | [55] | ||||
| 2010s | [122] | [249] | [50] | ||||
| 2020s | [27] |
3.1. Early Memory Models
Lamport [184] made the first exploration on the notion of the “intuitive correctness” of concurrent programs in the 1970s, and formalised the concept of what later came to be known as Sequential Consistency. From the late 1980s onwards there were considerable efforts in defining the semantics of various instruction-level parallelism techniques, such as pipelining and out-of-order execution (see Section 2) [65; 145; 160; 262]. This saw the proposal of many theoretical memory models, aiming to categorise and unify observed and potential behaviours. Examples include Weak Consistency [116], PRAM Consistency [194], Release Consistency [127], Cache Consistency [134], Processor Consistency [134], and Causal Consistency [21]. These memory models were unified in 2004 by Steinke and Nutt [272], who used a mathematical formalism to compare these memory models and to arrange them in a lattice. (In 2018, these theoretical models were again unified and compared [256; 257].)
Before the advent of multicore processors, computers that had multiprocessing were mostly large-scale machines at scientific institutions, and only the programmers at these institutions needed to handle weak memory behaviours. Then, as multicore processors rose to the mainstream in the 2000s, weak memory behaviours became visible to more programmers in diverse areas. The focus of the research community thus shifted from defining new theoretical weak memory models to exploring the weak memory behaviour on multicore hardware and related higher-level programming languages.
3.2. Hardware Memory Models
Many processor designers and vendors have attempted to codify and formalise the behaviours of their respective architectures, often with reference to the theoretical developments outlined in Section 3.1. Compared to the memory models in Section 3.1, which are mostly based on reads and writes only, the hardware memory models are more complex due to the specialised fences and atomic primitives of each architecture. We consider several prominent examples below.
SPARC
The Total Store Order (TSO) and Partial Store Order (PSO) memory models were defined for the SPARC architecture of Sun Microsystems [263]. As the name suggests, TSO keeps the writes of each thread in program order, but a read that accesses some location can complete before an earlier write of the same thread to a location different from . PSO relaxes this further to maintain order only on writes to the same variable. SPARC also defines a Relaxed Memory Order (RMO) model, which “places no ordering constraints on memory references beyond those required for processor self-consistency” [271]. The TSO model of SPARC is formalised in Isabelle/HOL [150].
x86
The x86 family of architectures is widely implemented by various companies such as Intel and AMD. Aiming to devise a rigorous specification of the behaviour of x86, Sarkar et al. [255] developed the memory model x86-CC based on Causal Consistency (see Section 3.1). However, x86-CC was later found not to match the actual behaviour of the hardware in all cases. Consequently, Owens et al. [227] developed x86-TSO, after observing similarities between the behaviour of x86 and the TSO memory model of SPARC. To this date, x86-TSO is widely accepted as the correct description of the behaviour of x86 architectures. In [227], x86-TSO was introduced in terms of two formulations, one operational and one axiomatic. The operational formulation uses a write buffer to model the weak memory effects, and is influential on many later formalisations of memory models; we provide a version of it in Section 4. The two key papers [227; 255] that developed x86-TSO are summarised by Sewell et al. [260].
Power
The Power family of processors by IBM has a very relaxed memory model that lacks multicopy atomicity (see Section 2.4), permitting many possible behaviours. An axiomatic formulation for Power’s memory model was first defined by Corella et al. [92], which was improved by Adir et al. [15] a decade later based on observed behaviours of litmus tests. Alglave et al. [33, 35, 36] also studied the Power memory model based on an axiomatic definition. An alternative, operational approach to formalisation of Power was given by Sarkar et al. [254].
Arm
Whereas the desktop and laptop markets are dominated by x86, as of 2025 the mobile device market is dominated by Arm architectures. Arm architectures previously did not forbid non-multicopy-atomic behaviours, but after a major revision in 2017, Arm has mandated multicopy atomicity from Version 8 onwards [240; 27]. There have been numerous works on specifying versions of the Arm memory model using different formalisms, such as [122]. The development history is included in the latest official paper on the Arm memory model [27].
RISC-V
The RISC-V architecture [249], first released in 2010, is an open-source architecture that attempts to remove some of the problems associated with commercial-in-confidence manufacturing. Its memory model is based on experiences at Arm, and includes many recent developments in weak memory theory, including support for release-acquire primitives.
Specialised Hardware
Recent research effort has focussed on GPUs [24; 270; 198] and FPGAs [151], and their combination with CPUs [130]. GPUs in particular focus on significant parallelisation of processing power, and hence would benefit from formalisation; NVIDIA has provided a rigorous description of its memory model [224], which lacks multicopy atomicity (see Section 2.4). Earlier formalisations determined a type of total-store-order model [24], while more recent formalisations suggest a weaker model again [198].
Earlier Hardware Memory Models
Intel’s Itanium processors [14] never became as commercially popular as other processors, and have since ceased production. Similarly, DEC’s Alpha processors were prominent for a time, but have also ceased production.
3.3. Software Memory Models
In the interest of portability, most languages provide their own abstraction and generalisation of the underlying hardware weak memory models, so that programmers need only consider a single set of primitives. This also constrains compilers to certain regular transformations in the presence of instruction dependencies. We consider some of the major languages below with a focus on the development of formal models; we recommend the wider survey of software memory models by Moiseenko et al. [219] for those interested in a comparison of features and limitations.
Java.
Sun Microsystems released the Java programming language in 1996, with multi-threading and shared-memory behaviour informally described in [137].
This was immediately followed by multiple independent efforts to formalise the memory model of Java [43; 132; 148]. There is a formulation that explicitly used local working memories [58], and a more abstract formulation with event structures [76]. In 1999, however, the original prose memory model turned out to contain flaws [238; 239]. Subsequent work [203; 206; 252; 296; 297; 44; 207] attempted to revise the Java memory model, culminating in [208], which describes the new Java memory model as part of Java 5.0.
After the formulation in [208] became accepted as the official Java model, a few alternative specifications continued to emerge [298; 154; 195; 53]. On the compiler-related side, Ševčík and Aspinall [258] identified common optimisations performed Java compilers that do not comply with the Java memory model.
C and C++.
Unlike Java, which included concurrency constructs in its specification from the very start, the C and C++ programming languages were originally designed without built-in mechanisms for concurrent programming. Instead, concurrency was enabled by a separate library of multi-threading constructs. However, as Boehm [54] pointed out in 2005, “threads cannot be implemented as a library”: the correctness of concurrent code cannot be guaranteed when the concurrency mechanisms are designed separately from the core language and its compiler.
Boehm and others worked as part of the International Organisation for Standardisation (ISO) on the formal semantics of concurrent C and C++ programs [55; 56]. The resulting 2011 ISO standards of the C and C++ programming languages were the first time that multi-threading in both languages was formally defined, and the memory model defined by these ISO standards is commonly referred to as the C memory model (or C11 in the literature, owing to its inception in 2011). The definition of the C memory model was then consolidated into a formal axiomatic formulation by Batty et al. [50], and later into a formal operational formulation by Nienhuis et al. [221].
The prominence of the C language and its memory model has drawn much attention in the context of verification, and many recent publications are devoted to program logics [295], model checkers [173], or other problems [266; 265; 86]; more of these are covered in Section 8.
Software memory models are often subject to more potential issues, due to the wider range of transformations open to a compiler, on top of the weak memory behaviours on the target hardware. These issues include the “out of thin air” (see Example 3.3) and “read from untaken branch” behaviours. These have been particularly addressed in the context of C, and the question about how to forbid out-of-thin air (and similar) behaviours without overly constraining standard compiler optimisations is a subject of ongoing debate in the standardisation committees and the academic community [57; 48; 155; 232; 226; 79]. The most recent C reference explicitly gives a version of out-of-thin-air behaviour, stating that it is a theoretically allowed behaviour, but with a caveat that no compiler should ever implement it.
[ht]
With the memory model of Power lacking multicopy atomicity and hence being harder to handle (see Section 3.2), there have been works that focus on the compilation from C to Power while respecting the memory models of both ends [49; 253; 183], and more general compiler optimisations under the C memory model [233; 109; 126].
The Linux Kernel.
Alglave et al. [32] provide a formalisation of the memory model of the Linux kernel. This model is to further codify the complex interaction of the wide variety of hardware memory models and the C language memory model, supporting some of the common patterns specific to Linux kernel programming. The formalisation involved interaction with the developer community, and the effort helped identify and resolve several issues, especially regarding locks and the read-copy-update (RCU) paradigm. As such this work represents a strong argument for the practicality of involving formalisms in non-trivial software development.
Other Programming Languages.
Every programming language that supports shared-memory concurrency has interplay with weak memory models. For programming languages such as JavaScript [135; 290], OCaml [114], OpenCL [167] and Rust [102], there have been works that study their own memory models, or how they are affected by the weak memory models of the languages to which they may be compiled. A potential simplification for this effort for many languages is via a memory model for an intermediate language such as LLVM [78; 186], and through the investigation of general properties related to safe optimisation strategies [136].
4. Introduction to Operational Formalisms
Before giving a general survey of the field we provide an example formalisation: a simplification of the formalisation of x86 given by the influential work by Owens et al. [227; 260]. This formalisation is operational, meaning described by way of individual steps of the system, and as such is often easier to follow than other types of semantics [159]. We also choose it as it is based around a microarchitectural feature, and as such provides a contrast to axiomatic descriptions, which purposefully abstract away from such considerations. In this formalisation, weak memory effects (specifically, loads appearing to take effect before earlier stores) emerge by the use of a write buffer in each core (acting similarly to a pipeline (Section 2.1) or as part of the cache system [212]). We first give an operational semantics for a standard sequential language, as if instructions were executed in the order specified by the program text; after extending that semantics with extra rules covering write buffers the weaker behaviours of x86 (Total Store Order) become apparent. More complex formalisations based on microarchitectural mechanisms are outlined in Section 4.4; see also Section 6.1.
Notation.
Let be a function from domain type to range type . Viewing a function as a set of pairs, we sometime write in place of when . Also, we write to denote the updated function that maps to and maps every other to its original . A typical list is written as , and the empty list is written as . We use ‘’ to append or prepend elements and lists, for example, is the list formed by appending lists and , and if and are elements, .
4.1. Syntax
We begin by introducing an assembly-like programming language. The basic types include values , local variables , and shared variables . By convention, ranges over , ranges over (registers), and range over . An expression can be formed from any of the standard arithmetic, logical operators, and local variables.
The syntax of commands is defined as follows.
A command may be one of three “update” instructions: a local assignment , a store to shared memory , or a load from shared memory . In all cases we assume expression refers only to registers and values. These instructions correspond to different uses of the x86 instruction type mov; for instance, “mov r1,r2” (), “mov [rx],r” (), and “mov r,[rx]” (), where register rx has the value of a location in memory (abstractly corresponding to a shared variable ). In addition to the three “update” instructions, a command could be a (e.g., mfence in x86). Instructions can also be composed sequentially, (though this does not guarantee sequential execution, of course), and a terminated command is .
A multicore system is formed from a single shared state of type , and two or more individual cores, composed in parallel (). Each core is itself formed from a local state of type and a command of type .
For instance, recall the store-buffer (SB) litmus test from Section 1.2. The initial configuration for this system (3) is that shared variables and map to 0 in the global state, and there are exactly two cores, each of which has a single local variable mapping to 0 and containing the corresponding code. To aid readability we write the structures in vertical form.
4.2. Semantics
Each possible execution of the concurrent system is viewed as a sequence of memory events, which may be a read-event , a write-event , a -event, or a silent event . A read observes the value for from shared memory, whereas a write updates the variable to value in shared memory. A -event is issued by a -instruction. A silent event denotes some register-only instruction that is resolved locally.
Local Operational Rules. The behaviour of a process is defined by transitions of the form , where a is formed from a local state and command c, and is the generated memory event. Fig.1 gives the transition rules for the five command types in (2).
Basic commands
System level
TODO
(seq) states that a sequential composition of commands may always take a step of . If has terminated (i.e., is ), then may be executed via (skip).
In (local), a local assignment updates in the local state to , the value of expression evaluated at the local state ; since it is resolved purely in the local configuration, it is a silent transition, having no interaction with shared memory.
In (load), returns some value for from shared memory, and updates the register accordingly. Similarly, in (store), writes the (locally evaluated) value to in shared memory. The final in each rule indicates that the instruction has finished executing, and will eventually be eliminated via (skip).
A -instruction – which has no effect in a sequentially consistent system – simply generates a -event; its influence will be seen in the next section on write buffers.
System-Level Operational Rules. In (par-config), a system with cores operating in parallel may take a step when any single core () takes a step. At the topmost level these generated memory events interact with main memory, . In these rules, is the parallel composition of processes, as in (par-config). In (mem-read), a read-instruction of , where the value of is in shared memory, may proceed, and the shared memory remains unchanged – the precondition for this rule prevents incorrect values being read in (load). In (mem-write), if any process issues a write of the value to the shared variable then the memory is updated accordingly. Finally, (mem-) states that silent events have no effect on main memory; a similar rule holds for instructions.
Termination. A process is said to terminate when it reaches a single command. A system is said to terminate if all processes terminate, and a system state, , is said to be reachable if the system from some starting state can terminate in via a sequence of top-level transitions of (mem-read), (mem-write), and (mem-).
[ht]
4.3. Write Buffers for x86
We now augment the system of Sections 4.1 and 4.2 with an explicit write buffer for each core, following [227]. A write buffer, of type , is a list of pairs , where is a shared variable and is a value being written to , representing a pending write of to . The head (left-hand side) of the write buffer is the oldest write, with new writes being appended to the end (right-hand side) of the write buffer. A core now contains a write buffer, in addition to the local state and command; the system-level structure is otherwise unchanged.
For instance, the starting configuration for SB is similar to that of (3), except that each core also has an (initially empty) buffer.
Fig.2 gives the operational rules of a (formed from a local state and command c) interacting with its local write buffer b. All previous transitions in Fig.1 still apply as-is, with the combination of (x86-write-buf) and (x86-read-mem) allowing loads to take effect before stores.
TODO
[t]
In (x86-write-buf), when a core issues a write via rule (store), rather than interacting directly with the shared memory it is appended to the end of the write buffer. Because this has no effect on the shared memory, it is a silent transition with no effect on the rest of the system.
In (x86-write-mem), the oldest write-event in a write buffer can be “flushed” to the shared memory. This transition is labelled by the write-event that was earlier “delayed” by Rule (x86-write-buf).
In (x86-read-mem), a core attempts to read the shared variable , where its write buffer contains no pending write to . The core reads directly from the shared memory, making a transition labelled by a read-event. Note that this read interacts with the shared memory earlier than any pending writes in the write buffer, despite being program-ordered-after the corresponding stores.
In (x86-read-buf), a core reading the shared variable finds a corresponding pending write to in its write buffer. Then the core locally takes the value from this (right-most) pending write. This “bypasses” the shared memory, and is a silent step with no interaction with the rest of the system.
In (x86-), a core can change from one local configuration to another via a silent transition, without modifying the write buffer. This occurs when the core itself makes a silent transition, for instance, via rule (skip).
In (x86-fence), a instruction can proceed, provided the write buffer is empty. This essentially forces all pending writes to be flushed (via (x86-write-mem)) to the shared memory, before further instructions in can be issued. A fence therefore ensures that reads after the fence occur strictly after the writes before the fence.
Combining the write-buffer rules (Fig.2) with the Sequential Consistency rules (Fig.1), we obtain an operational description of TSO-style memory models, including that of x86 architectures.
The built-in “delay” induced by the write buffers result in the weak behaviour observed in the store-buffer litmus test (Example 4.3). A standard view of weaker memory models is to have a single write buffer per variable, allowing independent writes to reorder within a single core, and to extend with a read buffer per variable, allowing reordering of reads; in essence, the pipeline itself serves as all of these facilities at once. The same set of rules can also be used to demonstrate forwarding; see Example 4.3.
In this presentation we have given each core its own write buffer; an alternative approach is to consider the write buffers to be part of a “storage subsystem”, separate to the cores; in that structure each core effectively operates sequentially, irrespective of the memory model. The memory model itself is wholly defined by the storage subsystem, which could include further buffers and cache mechanisms.
4.4. Operational Descriptions of More Complex Models
The write buffer concept neatly captures the essence of the TSO (x86) memory model – writes can be delayed in a buffer, keeping their relative order but allowing reads to proceed early. One way to extend the concept to related but weaker memory models, such as PSO (see Section 3.2), is to use one write buffer per variable. Then, writes to a variable are kept in order, but not necessarily with writes to another variable . In this setting, a fence must wait for all write buffers to be empty. Extending the concept further to individual load buffers (separate to the write buffers) may work for the weaker-again models of Arm and RISC-V.
Below, we outline how some of these more complex aspects of weak memory may be approached through an explicit model of a hardware pipeline as described in Section 2.1. However, a focus on microarchitectural details may become cumbersome to work with formally. For this reason many researchers focus on formal representations that abstract away from microarchitectural mechanisms, and focus instead on key relationships between operations across multiple processes in the system. This is typically referred to as an axiomatic approach, to which we turn our attention in Section 5. Operationally one of the best-known approaches is the Promising Semantics [165; 187; 241], as discussed in Section 6.2, which combines an operational semantics with global axioms.
Arm and RISC-V
The TSO memory model keeps the ordering amongst all stores, to any variables. A key relaxation of this is to allow stores to different variables within a process to execute out of order. Operationally, this can be instituted by allowing stores to be propagated to main memory even when not at the head of the buffer, as given by the following rule for Arm, which is similar to (x86-write-mem).
In this case we assume all instructions, not just writes, are fetched into the pipeline (generalising the buffer), similarly to (x86-write-buf). Provided that there are no writes to (of the form ), reads of (of the form ), or fences in the section of the pipeline preceding the write in question (), then the write can be issued. A similar rule can be constructed for reads. These constraints encode those of Definition 2.1, modified by fences as in Example 2.2, and allow the weaker behaviours shown in Example 2.1. A more complete operational description of pipelines that conform to the Arm memory model, and includes branches and jumps, is given in [90].
Release/Acquire
The release/acquire concepts [127], as introduced in (1) and Example 2.2, can also be described operationally. Continuing the generalisation of write buffers to pipelines, an operation with release semantics can be executed only when at the head of the pipeline; i.e., when all other prior operations (w.r.t the program order) have completed. A load with acquire semantics prevents later memory operations from completing earlier; hence we extend the constraints of (Arm-write-mem) to capture this possibility.
Rule (release-write) is similar to (x86-write-mem) in that a release write is issued only when at the head of the pipeline. Rule (RA-write-mem) extends (Arm-write-mem) with an extra condition, which we highlight with a box. This additional constraint means that an earlier acquire load of any variable prevents a later write from proceeding. For the stronger form of release/acquire consistency in [127] an acquire load additionally requires the absence of release stores in the earlier pipeline, which can be described syntactically similarly to the boxed constraint of (RA-write-mem).
Atomic operations
Instruction types such as swap, compare-and-swap, fetch-and-increment, etc., combine read and write operations atomically. These can be described using the concept of a global lock which prevents a core from performing a memory operation while another core holds the lock [261]. For instance, the atomic swap instruction, , updates the shared variable to hold the value in register , and updates to the initial value of ; this instruction can be defined as follows.
Conceptually, the operation sets a flag that prevents other cores from carrying out any read or write, until ed. The global can be specialised to per-variable locks, /, to allow the reordering between accesses of independent locations, as in Arm and RISC-V, while maintaining apparent atomicity per operation. The operation may also be weakened to allow concurrent cores to read while it is locked.
5. Introduction to Axiomatic Formalisms
The semantics of the previous section, based on write buffers, incorporates the definition of the memory model into the rules governing of the structures of the system. One of the more common approaches to formalising weak memory behaviours is to abstract from mechanisms and instead describe properties of traces of the system. In this section we provide an introduction to this common and effective style via a simplified version of the x86 (Total Store Order) memory model. The extension of these concepts to more complex models is outlined in Section 5.4; see also Section 6.3.
Fundamental notation and concepts
An axiomatic formalism specifies a memory model in terms of axioms that forbid certain executions of a concurrent program. The concurrent program is represented as an execution graph (Definition 5.7), which comprises the set of memory events, and a set of binary relations on . A memory event comes in the following types.
denotes an initialisation-event that initialises the location to the value .
denotes a read-event issued by Thread and reads the value from the location .
denotes a write-event issued by Thread and writes the value to the location .
A Fence event corresponds to a fence-instruction.
The binary relations are the subject of Section 5.1. Some of these relations are common to all memory models; some are specific to individual memory models; some are specific to instruction-types on particular hardware-platforms; some are derived from other relations using relational operators such as converse , composition , intersection , union , and transitive closure , among others.
Axiomatic formalisms also make use of various restrictions of the identity-relation. Given a set of memory events, denotes for every event in the execution graph.
Given an execution graph, an axiom on some relation takes one of the following forms, where the initialisation-events are excluded.
: The memory events can be arranged into a linear ordering that preserves .
Such a linear ordering is a “topological ordering” of the directed graph of ,
and it is well known that there exists a topological ordering if and only if the directed graph contains no directed cycle.
: No two memory events can be related by .
: No memory event can be related by to itself.
5.1. Fundamental Relations
Table 4 summarises the key relations that will be covered in the rest of this section. Some of the relations are derived, as indicated in Fig.3. In what follows, we use the variable convention where denotes a write- or initialisation-event, denotes a read-event, and denotes any type of memory event.
| standard | alternatives | ||||
| loc | same-location | Def. 5.1 | |||
| ext | external | Def. 5.1 | |||
| po | program order | Def. 5.2 | sb | sequenced-before | [50] |
| poloc | po per location | Def. 5.3 | |||
| fence | fence relation | Def. 5.4 | |||
| co | coherence order | Def. 5.5 | mo | modification order | [50] |
| ws | write serialization | [22] | |||
| rf | reads-from | Def. 5.6 | wt | writes-to | |
| fr | from-read | Def. 5.8 | rb | reads-before | |
| com | communication | Def. 5.9 | |||
| ca | coherence-after | ||||
| eco | extended coherence order | ||||
| hb | happens-before | ||||
| sw | synchronises-with | ||||
| rfe | external reads-from | Def. 5.13 | wte | external writes-to | |
TODO
Definition 5.1 (loc and ext).
The same-location relation loc consists of pairs where and access (i.e., read from or write to) the same memory location.
The external relation ext consists of pairs where and are issued by different threads.
Definition 5.2.
The program order po is a disjoint union of total orders on all events excluding initialisation-events.
It is specified by the program text, with the requirement that if and only if the memory events and are issued by the same process.
The meaning of is that the instruction corresponding to precedes the instruction corresponding to in the program text.
Definition 5.3.
The program order per-location is defined as . That is, it consists of pairs where and access the same memory location.
The preservation of is the one of the criteria of the “coherence” property (Definition 5.10), which corresponds to the dependencies in processor pipelines (Definition 2.1).
Definition 5.4.
The fence relation is then defined as , where is the set of all Fence events in an execution graph.
Recall that is the identity-relation restricted to the set , hence fence consists of pairs where there exists a fence such that and .
Definition 5.5.
The coherence order co is a disjoint union of total orders on initialisation- and write-events.
It is specified by the program execution (as opposed to po, which is specified by the program text), with the requirement that if and only if (1) is an initialisation- or write-event, (2) is a write-event, and (3) and write to the same location.
The meaning of is that takes place globally before .
Definition 5.6.
The relation rf is specified by the program execution, and consists of pairs where (an initialisation- or write-event) and (a read-event) accesses the same location and have the same value. (That is, reads the value written by .)
Note that the converse of a valid rf must be functional. Indeed, means that reads from , and every read-event must have a unique “source”. This property underpins Definition 5.8 later.
On the name “reads-from”, here we follow the recent work by Arm [27], where a typical pair means “ reads from ”, with occurring before . Some may find it more intuitive to think of the relation rf as writes-to. Indeed, means that writes to the value read by .
The relations po, co and rf are the “base relations”. The set of memory events and these base relations constitute an “execution graph”, also known as a “candidate execution” [27].
Definition 5.7.
An execution graph comprises the set of memory events, and po, co and rf.
Next we present the key derived relations, extracted from a given execution graph.
Definition 5.8.
The from-read relation fr is defined to be .
A typical pair in fr has the form , where and access the same location, and there exists another write such that and .
One intuition behind fr is the following. Whereas the reads-from relation rf places a read strictly after the write from which gets its value, the from-read relation fr places each read strictly before the subsequent write that changes the relevant location. It thus “fills-in” the order of operations on a single variable. Some may find it more intuitive to think of fr as reads-before, since essentially means that reads a location right before changes it.
For instance, let and be ordered writes to , and is a read of that takes its value from . Then, the coherence order on is , and the reads-from relation is . Hence, the derived from-read relation is . This makes explicit that must occur before , or else it would have read a different value. Combining all these, the order of the events on is thus .
Other Common Relations
Below are some relations that are common in the literature but are not used in the examples in the rest of this paper. We include them here for completeness.
Definition 5.9 (com and ca).
The relations co, rf and fr above are sometimes combined into the following abbreviations.
The communication order .
The coherence-after relation .
The extended coherence order – defined to be the transitive closure of com [183].
In addition, software memory models make use of the following two relations to enforce ordering.
The synchronises-with relation (sw) – induced by certain synchronisation constructs.
The happens-before relation (hb) – contains po and sw.
These are induced by specific language features, such as the “release/acquire” tags, and “volatile” variables in C and Java.
5.2. Axiomatic Definition of Sequential Consistency
Definition 5.10 (Coherence).
An execution graph yields a coherent execution if:
In other words, a possible execution is coherent if it preserves this relation.
A memory model is said to satisfy the coherence property if all possible executions of any execution graph is coherent. The coherence property is also known as the uniproc property [22], SC per location [36], SC per variable [26] or internal visibility [27].
The coherence property is satisfied by most memory models. It ensures that possible executions do not contain “invalid” pairs of events such as and then . Indeed, once Thread has written 2 to globally, Thread should not be able to read 3 from . (See also: the discussion following Definition 5.15.)
Definition 5.11 (SC Axiom).
An execution graph yields a sequentially consistent execution if:
Note that the coherence property follows from this definition because . Examples 5.2 and 5.2 now demonstrate how this definition is applied.
[t]
5.3. Axiomatic Definition of TSO
Compared to Sequential Consistency, the Total Store Order memory model allows two additional phenomena: (1) reads can take effect earlier than writes, making write-read pairs appear reordered; and (2) a process can read directly from its own writes before they are observed globally. This latter phenomenon is known as forwarding (see Section 2.3).
As instructions can appear reordered due to (1) above, fence instructions (see Section 2.2) are used to prevent any undesirable reordering. In this section, we introduce the axiomatic concepts for reasoning about write-read reordering, Fence instructions and forwarding.
To capture the phenomenon of write-read reordering, we define the following four relations to filter certain types of pairs from the program order.
Definition 5.12 (Local Event Relations).
| RR | RW | WW | WR |
Given a pair with and issued by the same thread, forwarding (Section 2.3) can be phrased as: reads the value written by , but appears to take effect earlier than in the global order (violating the principle of the reads-from relation). Hence, forwarding is described axiomatically by relaxing rf, and dropping those write-read pairs in rf whose events are issued by the same thread. This is described by the “external” reads-from relation.
Definition 5.13 (rfe).
The external writes-to relation, written rfe, consists of the pairs where and are issued by different threads.
Definition 5.14 (PPO of TSO).
The preserved program order of TSO is .
Definition 5.15 (TSO Axioms).
An execution graph yields a possible execution under TSO if:
The first axiom is the coherence axiom (Definition 5.10), and the second is the PPO axiom. TSO explicitly requires the coherence axiom, as illustrated by Example 5.3. Note that this is in contrast to SC (Definition 5.11) where the coherence axiom is implied.
[ht]
[t]
5.4. Axiomatic Descriptions of More Complex Models
For illustrative purposes, we have presented how axiomatic formalisms can define the straightforward SC memory model and a simplified version of TSO. In practice, axiomatic formalisms have been used to specify a wide range of behaviours related to more complex memory models and instruction types, e.g., [36; 240; 27], as surveyed in Section 6.3. Below we give an indication of how these more complex problems are tackled axiomatically.
Arm and RISC-V
Arm and RISC-V further weaken TSO by removing the RR, RW, and WW relations (Definition 5.12) from the preserved program order (Definition 5.15). This weakening corresponds to the principles of pipelining in Definition 2.1. This is straightforward to instrument and, because it is based on (subsets of) relations, clearly shows a weakening, for which approaches based on microarchitectural mechanisms (Section 4.4) would typically require more complex arguments.
Additional relations are required to handle aspects of assembler-level languages: data, where the value assigned to a variable is dependent on previous instructions; addr, where the calculation of an accessed address in memory is dependent on previous instructions; and ctrl, to handle branch instructions, where, on Arm and RISC-V architectures for example, later load instructions can be speculatively executed before a branch condition is resolved, though later stores cannot.
Substantial parts of the Arm and RISC-V memory models concern specific instruction-types; in particular, different types of fences and release-acquire constraints. For instance, RISC-V’s fence.r.r instruction enforces order between loads, similar to Arm’s load fence (dmb.ld).
Release/Acquire
Let the set include any operation with release semantics (e.g., a release fence, or a store designated as a release). Axiomatically, we require the preservation of the relation , so that any operation program-ordered before a release operation is guaranteed to have completed prior. This allows a release operation to act as a “flag” that some update has occurred in main memory (see, e.g., (1) and Example 2.2).
Similarly, let the set include any operation with acquire semantics (e.g., an acquire fence, or a load designated as an acquire). We then add , to the relations to preserve, meaning that any operation program-ordered after an acquire operation will occur after. In combination with a release constraint, this allows a form of message-passing communication via shared variables.
Finally, to maintain order between release and acquire locally, we require . (The Arm official model also includes requirements between release/acquire operations and RMWs.)
Atomic instructions
Most hardware platforms support RMW instructions such as compare-and-swap (CAS) and test-and-set (TAS). Such an instruction reads from a memory location, performs some local computation involving the read value, and finally writes to memory, all in one atomic step. When constructing an execution graph, an RMW instruction corresponds to both a read- and a write-event. The atomicity of such a pair of read and write can be enforced axiomatically, i.e., the rmw relation is a subset of po, and relates a read and a write that are to occur atomically according to some atomic instruction type.
Relations fr and co are used to filter rmw events using ext (Definition 5.1) to obtain the pairs whose events are issued by different threads.
Now, a pair belongs to the relation precisely when and are interleaved by a write issued by another thread. In other words, means that and are not atomic. Hence, to ensure that every pair in rmw is atomic, we require that no pair in rmw is in , that is, an execution graph enforces the RMW instructions if . Further details and mechanised tool support can be found in the online repository of the Herd tool [39; 27].
6. Survey of Memory Model Formalisms
In Sections 4 and 5 we gave detailed presentations for simplified versions of operational and axiomatic formalisms. In this section, we expand on the relevant literature, and also cover other approaches to representing memory models.
6.1. Mechanistic Formalisms
Mechanistic formalisms describe the effects of weak memory models with structures that operate in the manner of hardware mechanisms. The best-known of such mechanisms is the local write buffers elucidated in Section 4. Local write buffers describe the behaviour of TSO [263], on which the x86 memory model is based [227; 255; 260]. As local write buffers are relatively straightforward to encode, they often serve as the foundation for verification under TSO. Namely, when each write buffer is modelled as a separate process, one can analyse the entire system under Sequential Consistency using existing tools [68; 280; 281; 292; 282; 190; 191; 192; 273].
The mechanistic formalism of Boudol and Petri [63] is also based on write buffers that are not local but potentially hierarchical, making this formalism capable of expressing behaviours weaker than those of TSO. Another recent variant of write buffers is read buffers, proposed by Abdulla et al. [4] to improve various aspects of verification under TSO. Meanwhile, two other major hardware platforms – Power [254] and Arm [122; 240] – also have their own mechanistic formulations, albeit much more complex than x86-TSO and harder to encode for verification tools. A semantics based directly on pipelined and speculative execution, as outlined in Section 2.1, has been developed for Arm, RISC-V and x86 processors [90].
The main advantages of mechanistic formalisms is their intuition and compositionality: as they align closely with the actual mechanisms on processors, they give a local understanding of how weak memory behaviours arise, without needing to understand the system as a whole. On the other hand, the definition of the memory model itself is integrated in the operational rules, rather than being a stand-alone definition. It is also harder to show which behaviours cannot happen without a full analysis of all traces, in comparison with axiomatic approaches. Finally, a mechanistic description does not necessarily coincide with the behaviour of a compiler.
6.2. General Operational Formalisms: Timestamped Messages, Promises and Views
Hardware-like mechanisms are insufficient for describing the behaviours of compilers – the source of complexity for the C memory model especially. In order to reason about the C memory model operationally, many formulations make use of the notions of timestamped messages and views. Under weak memory models, a write does not become visible to other threads immediately. Thus, each write can be seen as a message that is broadcast by the issuing process, and it is up to the memory model that dictates when this message becomes visible to which other processes. Each process has a view of a subset of all the messages that have been sent out, and each message has a timestamp attached. By looking at the timestamps, a process updates its view according to the memory model. For instance, instead of just , a write-event is now a pair , where is the time at which the event took place. The timestamp can be some future time, meaning that any subsequent read of with return-value must occur at some later time , i.e., . An early use of timestamps is provided in [234]; we look at some other prominent formalisations in this vein below.
The “Promising Semantics”
In this framework, introduced in Kang et al. [165], each timestamped message is a promise to fulfil a later write. This semantics was later simplified by Lee et al. [187], and adapted to Arm and RISC-V by Pulte et al. [241]. The semantics has been used as the basis for verification of non-trivial systems [277].
A characteristic of the Promising Semantics is that it is a multi-execution formalism [218], as opposed to a single-execution formalism. In a single-execution formalism (e.g., that of Section 4), one behaviour, e.g., a trace of memory operations, corresponds to one execution of the code. In the Promising Semantics a single behaviour comprises more information that just the sequence memory events, and this extra information also covers multiple possible executions of the code. The multi-execution characteristic can also be applied to partial-order-based representations (see Section 6.4).
Views and Potentials
Formalisms with timestamped messages and similar notions have since been developed on several fronts, for example, Doherty et al. [111], with subsequent works leading to program logics [96; 295; 110; 98; 99; 45; 291]. Lahav and Boker [178] introduced the notion of potentials. In contrast to timestamped messages which reason about stores that have occurred in the past, potentials reason about possible stores in the future.
Timestamped messages, views, and potentials enable general frameworks for handling a wide range of weak memory effects, including the lack of multicopy atomicity (Section 2.4) and out-of-thin air behaviour (Section 3.3). This flexibility is not without disadvantages, however: these formalisms are perhaps the most complex of those surveyed, with the details of even small examples such as the store buffering litmus test being difficult to reason about [218].
6.3. Axiomatic Formalisms
One of the most successful approaches to formalising memory models is via constraints on the allowed behaviours. This is referred to as the axiomatic (or declarative) approach in the literature [36; 263]. Given a set of memory events (where each read has a return value), one starts with all possible permutations, and retains only those permutations that satisfy the axioms of the particular memory model (conceptually the opposite of first considering the set of behaviours allowed under Sequential Consistency and then expanding that set to capture the weak behaviours).
Due to its preciseness and relative conciseness, the axiomatic approach is widely adopted for the major memory models, e.g., x86-TSO [227], Arm [27], RISC-V [249], Java [208], and C [50].
Some of the most influential tools in this area are the domain-specific language Cat and its associated litmus-test simulator Herd, introduced in [36]. Cat and Herd are utilised by Arm to formalise its memory model [27], and Cat is also used in a number of model checkers for expressing other memory models (see Section 8.1). An axiomatic formalisation is leveraged by the Check suite of tools and techniques for investigating correctness and security of microarchitectural implementations [197; 204; 283; 205; 199; 284].
The downside of axiomatic formalisms is that they are based on global reasoning about completed executions, where every has a value, and the relations and are supplied (see Section 5). As such, axiomatic formalisms tend not to be suited for reasoning about programs compositionally; that is, one cannot form the set of behaviours of a system by considering individual processes in isolation. The purely axiomatic approach has less applicability to theorem proving (however, see [26; 143]), so operational semantics (see Section 6.2) are sometimes developed in tandem, often with an underlying axiomatic flavour [110; 165; 241; 95].
6.4. Event Structures and Pomsets
To express and reason about the subtle behaviours of more complex memory models such as the C memory model, there are works that use event structures and pomsets (partially ordered multisets), which can be viewed as mathematical structures that generalise execution graphs in Section 5.
Both event structures [294] and pomsets [237] were invented in the 1980s as mathematical models of general concurrency. An early use of event structures for weak memory models is by Cenciarelli et al. [77] in 2007, who described a small fragment of the Java memory model by combining event structures with operational and axiomatic techniques. Then from 2016 onwards, there have been many attempts to define memory models using related ideas [74; 155; 232; 79; 231; 153; 156; 218]. The Weakestmo Semantics and its extension [79; 218] are multi-execution, i.e., they encode multiple execution graphs in a single event structure (cf. the Promising Semantics Section 6.2).
6.5. Program Transformations and Instruction Reorderings
A number of early works [17; 210; 211] specified weak memory models in terms of which pairs of instructions can be reordered. This aligns with the workings of the processor pipeline (Definition 2.1), and is often the level at which processors are described informally. The first formal memory-model framework based on instruction reordering was developed by Arvind and Maessen [41], who included store-atomicity as an extra criterion.
Possible instruction reordering is important in the context of compiler-related aspects of weak memory models. Common program transformations (including instruction reordering) performed by optimising compilers are not always correct under weak memory models [258; 288]. To formulate what it means for a program transformation to be “correct” under a memory model, one typically develops some trace-based denotational semantics, and a transformation is then correct if the denotation of the transformed program is (a subset of) the denotation of the original program. Works in this line of research include Ševčík and Aspinall [258] and Burckhardt et al. [69] for the Java memory model, and Dodds et al. [109] and Dvir et al. [117] for the C memory model and its RA fragment.
Instruction reordering was also the basis in [85], where the memory model is encoded as a parameter to sequential composition on the level of the program text. Weak memory effects are explained in terms of the standard operators for (no-parameter) sequential composition and parallel composition, via a set of transformation rules on the instructions, in the spirit of the Concurrent Kleene Algebra [149; 146]. For example, instantiating the SB litmus test on an x86 (TSO) machine, one can derive the following.
Treating the programs on the right-hand side as under Sequential Consistency, one can observe that on the first line, the loads and can execute before the stores. Meanwhile, when fences are inserted as on the second line, instances of the x86-specific sequential composition are equivalent to standard sequential composition.
This reasoning framework has been used as the basis for addressing security vulnerabilities related to speculative execution [91] and information flow [267]. By transforming a program under a weak memory model to a program under Sequential Consistency, standard reasoning methods and tools become amenable, saving the need for specialised logics and techniques. However, in contrast to timestamped messages, these program-level transformations cannot explain non-multicopy-atomic behaviours [182].
7. Program Logics and mechanised support
There are three prominent families of program logics for concurrent programs that have been adapted to reason about weak memory effects: Owicki-Gries (OG) [228], Rely-Guarantee (RG) [157; 158], and Concurrent Separation Logic (CSL) [152; 247; 64]. This section surveys these adaptations; a summary is given in Table 5.
| Logic | Formalism | Memory models | Case studies | References |
| OG/RG | Axiomatic | C fragment | RCU | [181] |
| RG | Operational | x86-TSO | Simpson’s | [248] |
| RG | Operational | (parameterised) | litmus tests | [13] |
| OG | View-based | C fragment | Peterson’s, RCU | [96; 99; 110; 295] |
| RG | potential-based | Causal Consistency | litmus tests, Peterson’s | [179] |
| CSL | Axiomatic | C fragment | MS queue and others | [112; 113; 289; 286] |
| CSL | Operational | C fragment | MS queue and others | [164] |
| CSL | Operational | Rust | Rust library | [102] |
| CSL | Operational | OCaml | locks, bounded queue | [217; 216] |
| RG | Reordering | hardware models | Chase-Lev, Peterson’s | [93; 94] |
| Invariant | Axiomatic | (parameterised) | PostgreSQL | [30; 26] |
| Invariant | Operational | C fragment | Peterson’s | [111] |
Axiomatic OG and RG
Based on an axiomatic formulation of the release-acquire fragment of the C memory model, a rely-guarantee logic was developed and applied to the read-copy-update synchronisation mechanism [181].
Operational OG and RG
A rely-guarantee logic on the x86 assembly instructions was developed in [248], based on the operational formulation of x86-TSO [227]. This logic was applied to Simpson’s four-slot algorithm, and mechanised in HOL. The rely-guarantee logic of [13] works on general memory models that are specified operationally, with case-studies performed on litmus tests.
Another line of research [96; 99] devised a view-based formalisation (see Section 6.2) of the release-acquire-relaxed fragment of the C memory model, and built an Owicki-Gries logic based on this formulation. The main case-studies, carried out in Isabelle/HOL, are Peterson’s algorithm and the read-copy-update mechanism; related works in this line of research include [110; 295].
Similar to view-based semantics is the “potential-based” semantics, which has been used as the basis of a rely-guarantee logic for Causal Consistency [179].
Axiomatic CSLs
One line of research consists of the following three generations of logics: RSL (Relaxed Separation Logic) [289], GPS (Ghost states, Protocols, and Separation) [286], and FSL (Fenced Separation Logic) [112; 113]. All three are based on the axiomatic C memory model [50], and are mechanised in Rocq (formerly Coq). The case-studies of these logics include simple litmus tests, while GPS in particular was applied to more realistic concurrent data structures such as the Michael-Scott queue, the circular buffer, and the bounded ticket lock.
Operational CSLs
The Iris framework [163; 162] uses an underlying operational semantics to unify various concurrent separation logics from the setting of Sequential Consistency. Iris is mechanised in Rocq, which is also used by its derivatives listed below.
Iris was extended to the weak memory setting by combining with GPS. The results are iGPS [164] and iRC11 [102]; see also [275]. The case-studies of iGPS are the same as those of GPS, while the case-studies of iRC11 are drawn from a library of the Rust programming language.
Cosmo [217] is another CSL based on Iris for reasoning about multicore OCaml and its operationally defined memory model. The case-studies are a spin lock, a ticket lock, and Peterson’s algorithm. The sequel [216] then applied Cosmo to verify a concurrent bounded queue. See [121] for an alternative approach to CSL.
RG Based on Reorderings
There is a rely-guarantee logic [93] for multicopy-atomic hardware memory models, using a reordering formalism [89; 85]. The case-study is the Chase-Lev double-ended queue for Arm, with the proof mechanised in Isabelle/HOL. The same authors later extended their framework to handle memory models without multicopy atomicity [94], and verified Peterson’s algorithm in Isabelle/HOL. The CLH fine-grained lock, as used in the seL4 microkernel [169], is verified for Arm using a related rely-guarantee-based approach in [88].
Others
8. Model checking and computational problems
This section surveys the techniques and tools for model-checking programs that run under weak memory models. Section 8.1 covers model checkers. Section 8.2 concerns related algorithmic problems such as consistency checking, and summarises results on decidability and computational complexity.
8.1. Model Checkers
| Name | Type | Memory models |
| CBMC extended [31] | BMC | x86, Power |
| Herd [27] | simulator | parametric (Cat language) |
| SATCheck [104] | SMC | TSO |
| CDSCheck [222; 223] | SMC | C |
| Dartagnan [142] | BMC | parametric (Cat language) |
| GenMC [175] | SMC | parametric |
| Zord [119] | BMC | TSO, PSO |
| McSPIN [14] | BMC | parametric |
| Nidhugg [9; 2] | SMC | TSO, PSO, Power |
In the weak-memory realm, most model checkers employ either Stateless Model Checking [129] or Bounded Model Checking [83]. Here, we classify the prominent model checkers into those based on axiomatic formalisms (Section 5), and, less commonly, those based on operational formalisms (Section 6.2). Table 6 summarises these model checkers, together with the memory models they support. In the text below, we also cover the case-studies to which these model checkers have been applied.
Based on Axiomatic Formalisms
Model checking under weak memory models is mostly based on axiomatic formalisms. The most prominent axiomatic formalism is the Cat language, with the associated Herd tool [36]. Herd simulates the possible behaviours of assembly programs, under memory models specified in Cat. Herd and Cat have been applied to thousands of litmus tests to check whether a specification of a memory model does agree with the hardware behaviour. They were used by Arm as the basis for developing the specifications of its memory models [27]. Herd is included in the “diy tool-suite” [39], which includes facilities for automatically generating litmus tests and running them directly on hardware.
The CBMC bounded model checker [84] was extended to handle the memory models of x86 and Power [31]. Apart from litmus tests, the case-studies include components of system software, such as a queue mechanism of the Apache HTTP server, the worker synchronisation mechanism in PostgreSQL, and the read-copy-update mechanism of the Linux kernel.
Dartagnan [235; 125; 142] is a bounded model checker, where users specify memory models in the Cat language. The benchmarks are several mutual exclusion algorithms, including Dekker’s algorithm, Peterson’s algorithm and the qspinlock of the Linux kernel [141; 229].
Another recent bounded model checker is Zord [119], which is based on a novel axiomatic formalism that supports TSO and PSO, and has been applied to the SV-COMP 2020 benchmarks.
Based on Operational Formalisms
There are comparatively fewer model checkers that are based on operational formalisms to encode the weak memory models.
McSPIN [14] is an extension to the SPIN model checker, and users specify memory models by constraining the operational model in McSPIN. The case-studies include litmus tests from the documentation of Unified Parallel C and Itanium, as well as the double-checked locking algorithm and Dekker’s algorithm.
Nidhugg [9; 2] uses stateless model checking based on a specialised operational formalism of memory models. The memory models supported are TSO, PSO, Power, and a subset of Arm. Nidhugg accepts standard C input, and has been applied to classical benchmarks, such as Dekker’s, Lamport’s and Peterson’s mutual exclusion algorithms.
8.2. Algorithmic Problems
At the heart of model checkers are computational problems for deciding whether a program satisfies certain properties. The decidability and computational complexity of these problems are substantial topics in their own right [59]. This section surveys the works that address these problems under weak memory models, with the results summarised in Table 7.
| Model | Consistency | (data-ind.) | Robustness | Reach. | (param.) |
| SC | NP-com [128] | FPT [80] | |||
| Causal | NP-hard [61; 124] | P [61] | |||
| TSO | NP-hard [124] | FPT [80] | PSPACE-com [62] | dec [42] | PSPACE-com [12] |
| PSO | NP-hard [124] | FPT [80] | dec [42] | ||
| RMO | NP-hard [124] | FPT [80] | undec [42] | ||
| Power | PSPACE-com [105] | ||||
| RA | FTP [285] | undec [3] | PSPACE-com [176] |
Consistency Checking.
The problem of consistency checking is to decide whether a concurrent program yields an execution that conforms to the given memory model.
This problem is mostly NP-hard. Indeed, when phrased in terms of axiomatic formalisms, only the program order (po) is inferred from program text, and one needs to explore the different choices for the coherence order (co) and the reads-from (rf) relation (see Section 5).
Gibbons and Korach [128] first showed that consistency checking under Sequential Consistency is NP-complete. Later, Cantin et al. [72; 73] showed that it is NP-complete to check if a concurrent program yields a coherent execution (Definition 5.10). As a corollary, consistency checking is NP-hard under most memory models, since most memory models satisfy the coherence property. Indeed, Furbach et al. [124] established the NP-hardness of consistency checking under the memory models covered by Steinke and Nutt [272].
To circumvent the NP-hardness, one can aim for less general solutions. For instance, Hangal et al. [144] developed an incomplete but polynomial-time tool for consistency checking under TSO.
Another way to deal with the NP-hardness is to add extra premises to the problem, such as the data-independence assumption: every value that is ever written is unique. This means that the reads-from relation (Definition 5.6) is already supplied, as every read ever only has one possible write as its source in the reads-from relation. Hence, one only needs to assign a suitable coherence order before checking the axioms, which makes consistency checking more tractable at times. For instance, though consistency checking under Causal Consistency is NP-hard in general, Bouajjani et al. [61] showed that it is in P when data-independence is assumed.
NP-hardness can also be approached using more fined-grained notions of complexity, namely “fixed-parameter tractability” (FPT). Under SC, TSO and PSO and assuming data-independence, Chini and Saivasan [80] showed that consistency checking is FPT: when the number of writes is fixed, then the problem is polynomial in the number of reads.
One can also parameterise the problem into the number of memory events and the number of threads. Also assuming data-independence, Tunç et al. [285] found near-linear parameterised algorithms for consistency checking under several variants of the C memory model.
Robustness.
Given a memory model M, a program is said to be not robust under M if it exhibits any extra behaviour that is not possible under SC. The robustness problem for a memory model M takes a program as the input, and computes whether running under M yields any extra behaviour that is not possible when runs under Sequential Consistency.
Reachability.
Also related to model checking (Section 8.1) is the reachability problem, which is to decide whether a given program has an execution from a given starting state to a given final state.
Under RMO the reachability problem is undecidable, but under TSO and PSO it is decidable, albeit beyond the primitive recursive (PR) class and has a very high computational cost [42]. In light of this computational hardness, there were works that attempted to trade completeness with efficiency under TSO [60].
9. Future Directions: Common Properties of Weak Memory Models
As surveyed in Section 6, there are many different representations that may be used to define memory models. With some exceptions noted below, generally missing from the literature are agreed-upon results or properties that formalisms should exhibit; in this section we outline how the field could develop in this respect.
9.1. Empirically-Established Properties
The influential “herding cats” work of Alglave et al. [36] robustly and rigorously tests weak memory models against litmus tests. This work included testing real hardware across a range of architectures and vendors, and identified a bug in a processor that resulted in a “Programmer Advice Notice” [38], and has been used in the development of Arm processors [27] and the Linux kernel [32].
Based on this work Arm has led the way in formalising aspects of its architecture [27; 28; 123], including publishing the memory model [196, Chapter B2] and the herdtools suite of applications [39; 23]. Ideally other companies and committees would follow suit, publishing, preferably, a formal model, or at least a representative set of litmus tests (small concurrent programs with an associated reachability condition Section 1.2) that cover both allowed (weak) and forbidden behaviours. This would support easier comparison across models and formalisms, and aid cross-compilation.
However, when focussing purely on empirical evaluation, it is unclear how to achieve complete coverage of all behaviours. This is obviously a difficult problem: given say 10 individual instruction types (loads, stores, fence types, memory ordering constraints, branch points, etc.), across just 2 concurrent processes each with 5 lines of code, there are approximately 10 billion possible litmus tests. This number does not include any attempt to generate the expected results. Furthermore, litmus tests typically do not cover loops, or non-trivial conditional statements. A “minimal”, maintained set of tests and expected results could be kept by an architecture vendor; programmers and formalists together could access this database, with both small litmus tests and larger programs explaining weak behaviours that can arise. This would help formalists check that their definitions and representations are correct, and, with the help of a common notation that abstracts from processor-specific syntax (e.g., [25; 26]), would aid comparison and re-use of results across different architectures and languages.
9.2. Formally-Derived Properties
Even if a representative set of tests can be developed, however, this will not address all concerns for software development. Programmers want to know the required solution to their programming problem – where to place fences or constraints to maintain the logical integrity of their concurrent algorithm – without needing to search through a database of litmus tests or to understand complex memory model specifications.
One theoretical aspect upon which the language-design community has more or less reached a consensus is that “programs that contain no data-race should behave as if under Sequential Consistency” [16], a property known as Data-Race-Free implies Sequential Consistency, or the Data Race Freedom Guarantee (DRF). This relatively early concept [18; 19] provides a base-level property for programmers of sequential, non-interfering code, similar to the principals underlying the circumstances under which localised pipelined reordering can occur (Section 2.1). However, the result does not necessarily help programmers of low-level systems code that does include data races, such as system-wide locks [32; 88].
What would complement such properties and litmus test conformance is a set of more general properties of programs that can be understood locally [46]; for instance, if a fence is inserted between instructions i and j, can standard ordering be assumed? If we let ‘’ represent strict ordering, not subject to memory model effects, then this property of fences can be expressed as follows.
This is the sort of property software developers need to understand their code: it does not mention any particular representation, and does not require a global analysis of the behaviour of other threads. We may generalise such properties as follows, where represents the set of traces that program generates – in any formalism of memory model – and where ‘’ represents concatenation of traces and represents interleaving of traces.
Property (11) would provide the underlying justification for the earlier program-level understanding of property (10), giving assurance that the insertion of a (full) fence allows reasoning about the separated parts of a program separately (i.e., compositionality). Property (12), for some model-specific definition of “independent”, indicates that low-level parallelism is a factor despite the use of sequential composition. Property (13) states that, for a given formalism, instantiating the memory model by the simplest – Sequential Consistency– reduces the program to a standard trace semantics, represented by . Even this relatively simple property – a sanity check on the definitions and structures used in a formal semantics – is rarely formally proved. A general class of properties of programs at this level would make the task of comparing formalisms more straightforward, and provide a relatively stable basis for programmers and verifiers to understand their task. Reduction to standard operators also opens up the possibility for existing tools and techniques to be applied for reasoning, reducing the need for specialised inference systems [26; 181; 93; 111; 110; 295], and expanding the range of properties and domains that can be addressed.
10. Current and Future Directions: adapting and contributing to software development
We outline some other research topics related to weak memory models, where software developers must face the consequence of weak memory effects. This includes ongoing work into verification methods and tool support, along with model checkers (Sections 7 and 8); in addition, formalisation of weak memory models can and should also impact other aspects of formal concurrency research, such as verified compiler efforts (e.g. [259; 284]) and data race detection [20; 166]. These topics typically require some abstraction from the specific representation of weak memory to facilitate more standard analyses; we give an overview of some of these below where not previously discussed.
10.1. Properties of Programs
Linearisability
Linearisability is a widely-accepted correctness criterion for concurrent data structures. In contrast to reachability of final states, linearisability describes the continual interaction of processes operating on a shared data structure. Based on its original definition (which assumed Sequential Consistency), Travkin et al. [280; 281] developed techniques to reduce a TSO program into an SC program, enabling one to utilise existing tools such as the SPIN model checker to test linearisability. Other approaches to adapting linearisability for application to weak memory models propose alternative versions of linearisability [107; 106; 269; 268].
Progress/Liveness
Most existing formalisms and tools for weak memory models handle safety properties only. Verifying liveness properties (also called progress properties) is more challenging even in the Sequential Consistency setting, with one reason being the need to address various fairness assumptions. In the weak-memory setting, works that deal with liveness properties only started to emerge recently [180; 7; 8].
Security Vulnerabilities
These may emerge from microarchitectural features, including through the exploitation of speculative execution and other mechanisms that cause out-of-order execution. The intersection of vulnerabilities and weak memory models is thought to be a potentially important area to investigate [75], and has been addressed by several authors utilising a range of different representations [283; 236; 108; 139; 91; 87; 90]. The addition of security-focussed instructions, such as Capability Hardware Enhanced RISC Instructions (CHERI), also impacts vulnerability research [299], and will require richer pointer-level models to address in the presence of weak memory.
10.2. Adapting Existing Code
Development of Libraries and Portability.
The specification of concurrent libraries with weak memory is a crucial topic, as concurrent libraries are the building-blocks of real-world concurrent programs. Based on formalisms for defining memory models, much recent effort has extended and applied them to specifying and verifying concurrent libraries. These works focus on various memory models, from TSO [67] and Java [118] to fragments of the C memory model [47; 97; 103; 242; 264]. Large-scale migration of code between hardware platforms has been recently addressed by Beck et al. [51], who developed a tool for automatically porting x86 code to platforms with weaker memory models, such as Arm and RISC-V. Such tools utilises fence insertion and model checking techniques, and also employs heuristics to improve scalability.
Fence Insertion.
A fence insertion algorithm aims at automatically finding a minimum set of positions in an original program to insert fences, so that the resultant program does not have any extra behaviour when executed under a weak memory model. This is important in the context of compiler-design: given a program written under the assumption of Sequential Consistency, the compiler needs to ensure that this program still has the original intended behaviour after it is translated to the hardware-language and run under some weak memory model. This issue was first addressed by Shasha and Snir [262]. Lee and Padua [185] developed a fence insertion algorithm based on analysing dominators in dependence-graphs under three weak memory models: Weak Consistency, Processor Consistency and Release Consistency. Fang et al. [120] empirically applied the same method to the memory models of actual hardware platforms, namely IBM’s Power and Intel’s Pentium, which had different types of fences.
For TSO and PSO, there are fence insertion algorithms based on modelling the write-buffers [6; 190; 191; 192] and on predicate abstraction techniques [5; 10; 11]. In the case of structured programs, Taheri et al. [276] developed a polynomial-time greedy algorithm for fence insertion, and also demonstrated the NP-hardness of the minimum fence insertion problem with multiple types of fences. Many other approaches to fence insertion have been taken [29; 52; 66; 177; 161; 214; 215], with the most recent being [225; 265]. Finding the minimal set remains an open problem, however, as does finding the most computationally efficient set.
10.3. Programming Styles and Hardware Variants
Transactional Memory
Transactional memory [140] defines synchronisation mechanisms that allow the programmer to specify atomic “blocks” of instructions. While these blocks should appear atomic, on the hardware level they can still occur in parallel. When two atomic blocks attempt to change the same variable and a clash arises, the transactional memory system can safely revert to a consistent state. Naturally, transactional memory has interplay with memory consistency models [82; 115]. Most recently, Dalvandi and Dongol [98] devised a semantics of transactional memory under the release-acquire fragment of the C memory model, and developed a program logic based on their semantics.
Persistent Memory
Persistent memory [131] is a relatively new type of physical memory that maintains its state even after system failures, with the aim of avoiding the loss of data. Updates to memory are held in the volatile memory first, and they are flushed to the persistent memory in some order, during which the persistent memory may temporarily be in some inconsistent states. In the case of a power outage, one needs to ensure that the entire memory system can be reverted back to a consistent state. These issues, related to memory consistency and cache coherence, have been addressed by formalisations of weak-memory models [245; 243; 81; 168; 287; 170; 244].
11. Conclusions
Weak memory models are in a sense historical legacies from optimisations made by hardware and compilers for single-core architectures (Section 2). Computational efficiences can be gained in some cases by reordering the execution of individual statements of code, with a minimum requirement being simply that the programmer will never know that the reordering has taken place. However, these changes, already somewhat exposed to programmers of concurrent code with agressive compiler optimisations, became significant with the advent of multicore architectures, initiating the first wave of attempts at formalising these effects (Section 3).
Often this issue can be addressed by having programming abstractions between the assembler and the high-level code; however this is not possible for programmers at the interface of hardware and software, for instance, operating systems developers [32]. In this context safety, security and performance are all of high importance, and a clear place where a rigorous formal model of the behaviour of multicore systems is invaluable [75]. Since low-level details are almost unavoidable when addressing weak memory model concerns, further refinement of formalisms to include detailed instruction semantics [40] would assist binary-level analysis (e.g., [251; 138]).
Approaches to formalising weak memory effects range from modelling features of the architectures directly (e.g., Section 4) through to abstracting away from such details to axioms of global traces (the axiomatic approach, Section 5). Within these broad classifications many other representations are used (Section 6), commonly including a partial order on events: a convenient representation for defining order between events where necessary, while leaving order unspecified when weak memory effects manifest. Many are supported by verification techniques (Section 7) and model checking tools (Section 8); however what is lacking from the field are rigorous sets of common properties, whether they be determined by small litmus tests or program-level statements of behaviour (Section 9).
Researchers at Arm have adopted the axiomatic style to specify their model, which is made concrete via conformance to a large set of litmus tests. This essentially makes the axiomatic style the de facto standard in describing weak memory models. Typically new semantics styles are validated by showing conformance to the expected outcomes of a set of litmus tests, of which canonical sets exist for both Arm and RISC-V. The weakness of the axiomatic style is that it does not lend itself directly to a straightforward semantics for programs. While memory models can be compared, and general properties of memory models described [36], one cannot easily take an abstract program and apply a set of (denotational) rules to obtain the set of traces for a program written in some abstract syntax. This is addressed by, for instance, the Promising semantics [165], an operational semantics which includes axiomatic-style relations. However this semantics, partly due to capturing a wide-range of weak behaviours, is quite complex to apply to even simple programs. Explanations based on microarchitectural features are usually more straightforward to describe and implement, but less general and less transferrable to the software level.
This leaves quite a gap between hardware developers, compiler writers, and eventually, software programmers. Low level systems code needs clear guidelines to get the most out of modern hardware capabilities, since the overuse of fences may lead to performance downgrades. There are also many other abstractions which would benefit from representations that are amenable to further analysis (Section 10), not tied too directly to weak memory notions. The field would benefit greatly from language and hardware designers talking more in terms of desirable properties, rather than representations.
Acknowledgements.
The authors thank Scott Heiner, Graeme Smith and the anonymous reviewers for their detailed and insightful feedback on the paper. This project was funded by Department of Defence and administered through the Advanced Strategic Capabilities Accelerator.References
- [1]
- Abdulla et al. [2017] P. A. Abdulla, S. Aronis, M. F. Atig, B. Jonsson, C. Leonardsson, and K. Sagonas. 2017. Stateless model checking for TSO and PSO. Acta Inform. 54, 8, 789–818. https://doi.org/10.1007/s00236-016-0275-0
- Abdulla et al. [2019] P. A. Abdulla, J. Arora, M. F. Atig, and S. Krishna. 2019. Verification of programs under the release-acquire semantics. In Programming Language Design and Implementation (PLDI). ACM, 1117–1132. https://doi.org/10.1145/3314221.3314649
- Abdulla et al. [2018] P. A. Abdulla, M. F. Atig, A. Bouajjani, and T. P. Ngo. 2018. A load-buffer semantics for total store ordering. Log. Methods Comput. Sci. 14, 1, 1–46. https://doi.org/10.23638/LMCS-14(1:9)2018
- Abdulla et al. [2012a] P. A. Abdulla, M. F. Atig, Y.-F. Chen, C. Leonardsson, and A. Rezine. 2012a. Automatic fence insertion in integer programs via predicate abstraction. In Static Analysis (SAS). Springer, 164–180. https://doi.org/10.1007/978-3-642-33125-1_13
- Abdulla et al. [2012b] P. A. Abdulla, M. F. Atig, Y.-F. Chen, C. Leonardsson, and A. Rezine. 2012b. Counter-example guided fence insertion under TSO. In Tools and Algorithms for the Construction and Analysis of Systems (TACAS). 204–219. https://doi.org/10.1007/978-3-642-28756-5_15
- Abdulla et al. [2023] P. A. Abdulla, M. F. Atig, A. Godbole, S. Krishna, and M. Vahanwala. 2023. Overcoming memory weakness with unified fairness: Systematic verification of liveness in weak memory models. In Computer Aided Verification (CAV). 184–205. https://doi.org/10.1007/978-3-031-37706-8_10
- Abdulla et al. [2024] P. A. Abdulla, M. F. Atig, A. Godbole, K. Shankaranarayanan, and M. Vahanwala. 2024. Fairness and liveness under weak consistency. In Taming the Infinities of Concurrency: Essays Dedicated to Javier Esparza on the Occasion of His 60th Birthday, Stefan Kiefer, Jan Křetínský, and Antonín Kučera (Eds.). Springer, 1–21. https://doi.org/10.1007/978-3-031-56222-8_1
- Abdulla et al. [2016] P. A. Abdulla, M. F. Atig, B. Jonsson, and C. Leonardsson. 2016. Stateless model checking for POWER. In Computer Aided Verification (CAV). Springer, 134–156. https://doi.org/10.1007/978-3-319-41540-6_8
- Abdulla et al. [2015b] P. A. Abdulla, M. F. Atig, M. Lång, and T. P. Ngo. 2015b. Precise and sound automatic fence insertion procedure under PSO. In Networked Systems (NETYS). Springer, 32–47. https://doi.org/10.1007/978-3-319-26850-7_3
- Abdulla et al. [2015a] P. A. Abdulla, M. F. Atig, and T. P. Ngo. 2015a. The best of both worlds: trading efficiency and optimality in fence insertion for TSO. In Programming Languages and Systems (ESOP). Springer, 308–332. https://doi.org/10.1007/978-3-662-46669-8_13
- Abdulla et al. [2020] P. A. Abdulla, M. F. Atig, and R. Rezvan. 2020. Parameterized verification under TSO is PSPACE-complete. Proc. ACM Program. Lang. 4, POPL. https://doi.org/10.1145/3371094
- Abe and Maeda [2016] T. Abe and T. Maeda. 2016. Observation-based concurrent program logic for relaxed memory consistency models. In Programming Languages and Systems (APLAS). Springer, 63–84. https://doi.org/10.1007/978-3-319-47958-3_4
- Abe and Maeda [2017] T. Abe and T. Maeda. 2017. A general model checking framework for various memory consistency models. Int. J. Softw. Tools Technol. Transf. 19, 5, 623–647. https://doi.org/10.1007/s10009-016-0429-y
- Adir et al. [2003] A. Adir, H. Attiya, and G. Shurek. 2003. Information-flow models for shared memory with an application to the PowerPC architecture. IEEE Trans. Parallel Distrib. Syst. 14, 5, 502–515. https://doi.org/10.1109/TPDS.2003.1199067
- Adve and Boehm [2010] S. V. Adve and H.-J. Boehm. 2010. Memory models: A case for rethinking parallel languages and hardware. Commun. ACM 53, 8, 90–101. https://doi.org/10.1145/1787234.1787255
- Adve and Gharachorloo [1996] S. V. Adve and K. Gharachorloo. 1996. Shared memory consistency models: A tutorial. Computer 29, 12, 66–76. https://doi.org/10.1109/2.546611
- Adve and Hill [1992] S. V. Adve and M. D. Hill. 1992. Sufficient conditions for implementing the data-race-free-1 memory model. Technical Report 1107. Department of Computer Sciences, University of Wisconsin-Madison.
- Adve and Hill [1993] S. V. Adve and M. D. Hill. 1993. A unified formalization of four shared-memory models. IEEE Trans. Parallel Distr. Syst. 4, 6, 613–624.
- Adve et al. [1991] S. V. Adve, M. D. Hill, B. P. Miller, and R. H. B. Netzer. 1991. Detecting data races on weak memory systems. ACM SIGARCH Computer Architecture News 19, 3, 234–243.
- Ahamad et al. [1995] M. Ahamad, G. Neiger, J. E. Burns, P. Kohli, and P. W. Hutto. 1995. Causal memory: Definitions, implementation, and programming. Distrib. Comput. 9, 1, 37–49. https://doi.org/10.1007/BF01784241
- Alglave [2012] J. Alglave. 2012. A formal hierarchy of weak memory models. Form. Methods Syst. Des. 41, 178–210. https://doi.org/10.1007/s10703-012-0161-5
- Alglave [2020] J. Alglave. 2020. How to generate litmus tests automatically with the diy7 tool. https://community.arm.com/developer/ip-products/processors/b/processors-ip-blog/posts/generate-litmus-tests-automatically-diy7-tool (Accessed October 2024).
- Alglave et al. [2015] J. Alglave, M. Batty, A. F. Donaldson, G. Gopalakrishnan, J. Ketema, D. Poetzl, T. Sorensen, and J. Wickerson. 2015. GPU concurrency: Weak behaviours and programming assumptions. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 577–591. https://doi.org/10.1145/2694344.2694391
- Alglave and Cousot [2016] J. Alglave and P. Cousot. 2016. Syntax and analytic semantics of LISA. arXiv:1608.06583 [cs.PL]
- Alglave and Cousot [2017] J. Alglave and P. Cousot. 2017. Ogre and Pythia: An invariance proof method for weak consistency models. In Principles of Programming Languages (POPL). ACM, 3–18. https://doi.org/10.1145/3009837.3009883
- Alglave et al. [2021] J. Alglave, W. Deacon, R. Grisenthwaite, A. Hacquard, and L. Maranget. 2021. Armed Cats: Formal concurrency modelling at ARM. ACM Trans. Program. Lang. Syst. 43, 2. https://doi.org/10.1145/3458926
- Alglave et al. [2024] J. Alglave, R. Grisenthwaite, A. Khyzha, L. Maranget, and N. Nikoleris. 2024. Puss in boots: on formalizing arm’s virtual memory system architecture. IEEE Micro, 1–9. https://doi.org/10.1109/MM.2024.3422668
- Alglave et al. [2014a] J. Alglave, D. Kroening, V. Nimal, and D. Poetzl. 2014a. Don’t sit on the fence: A static analysis approach to automatic fence insertion. In Computer Aided Verification (CAV). Springer, 508–524. https://doi.org/10.1007/978-3-319-08867-9_33
- Alglave et al. [2013b] J. Alglave, D. Kroening, V. Nimal, and M. Tautschnig. 2013b. Software verification for weak memory via program transformation. In Programming Languages and Systems (ESOP). Springer, 512–532. https://doi.org/10.1007/978-3-642-37036-6_28
- Alglave et al. [2013a] J. Alglave, D. Kroening, and M. Tautschnig. 2013a. Partial orders for efficient bounded model checking of concurrent software. In Computer Aided Verification (CAV). Springer, 141–157. https://doi.org/10.1007/978-3-642-39799-8_9
- Alglave et al. [2018] J. Alglave, L. Maranget, P. McKenney, A. Parri, and A. Stern. 2018. Frightening small children and disconcerting grown-ups: Concurrency in the Linux kernel. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 405–418. https://doi.org/10.1145/3173162.3177156
- Alglave et al. [2010] J. Alglave, L. Maranget, S. Sarkar, and P. Sewell. 2010. Fences in weak memory models. In Computer Aided Verification (CAV). Springer, 258–272. https://doi.org/10.1007/978-3-642-14295-6_25
- Alglave et al. [2011] J. Alglave, L. Maranget, S. Sarkar, and P. Sewell. 2011. Litmus: Running tests against hardware. In Tools and Algorithms for the Construction and Analysis of Systems (TACAS). Springer, 41–44. https://doi.org/10.1007/978-3-642-19835-9_5
- Alglave et al. [2012] J. Alglave, L. Maranget, S. Sarkar, and P. Sewell. 2012. Fences in weak memory models (extended version). Form. Methods Syst. Des. 40, 170–205. https://doi.org/10.1007/s10703-011-0135-z
- Alglave et al. [2014b] J. Alglave, L. Maranget, and M. Tautschnig. 2014b. Herding Cats: Modelling, simulation, testing, and data mining for weak memory. ACM Trans. Program. Lang. Syst. 36, 2. https://doi.org/10.1145/2627752
- Anderson et al. [1967] D. W. Anderson, F. J. Sparacio, and R. M. Tomasulo. 1967. The IBM System/360 Model 91: Machine philosophy and instruction-handling. IBM J. Res. Dev. 11, 1, 8–24. https://doi.org/10.1147/rd.111.0008
- Arm Ltd. [2011] Arm Ltd. 2011. Cortex-A9 MPCore(TM) Programmer Advice Notice: Read-after-Read Hazards. https://developer.arm.com/documentation/uan0004/a/ ARM Reference 761319.
- Arm Ltd. [2024] Arm Ltd. 2024. Herd7 Simulator (online). https://developer.arm.com/herd7 (Accessed October 2024).
- Armstrong et al. [2019] A. Armstrong, T. Bauereiss, B. Campbell, A. Reid, K. E. Gray, R. M. Norton, P. Mundkur, M. Wassell, J. French, C. Pulte, S. Flur, I. Stark, N. Krishnaswami, and P. Sewell. 2019. ISA semantics for ARMv8-a, RISC-v, and CHERI-MIPS. Proc. ACM Program. Lang. 3, POPL. https://doi.org/10.1145/3290384
- Arvind and Maessen [2006] Arvind and J.-W. Maessen. 2006. Memory model = instruction reordering + store atomicity. In Proc. 33rd Annual International Symposium on Computer Architecture (ISCA). IEEE Computer Society, 29–40. https://doi.org/10.1109/ISCA.2006.26
- Atig et al. [2010] M. F. Atig, A. Bouajjani, S. Burckhardt, and M. Musuvathi. 2010. On the verification problem for weak memory models. In Principles of Programming Languages (POPL). ACM, 7–18. https://doi.org/10.1145/1706299.1706303
- Attali et al. [1998] I. Attali, D. Caromel, and M. Russo. 1998. A formal executable semantics for Java. In Proc. OOPSLA Workshop on Formal Underpinnings of Java. ACM.
- Awhad and Wallace [2003] V. Awhad and C. Wallace. 2003. A unified formal specification and analysis of the new Java memory models. In Abstract State Machines (ASM). Springer, 166–185. https://doi.org/10.1007/3-540-36498-6_9
- Bargmann and Wehrheim [2023] L. Bargmann and H. Wehrheim. 2023. View-based axiomatic reasoning for PSO. In Theoretical Aspects of Software Engineering (TASE). Springer, 286–304. https://doi.org/10.1007/978-3-031-35257-7_17
- Bargmann and Wehrheim [2024] L. Bargmann and H. Wehrheim. 2024. Lifting the reasoning level in generic weak memory verification. In Integrated Formal Methods (IFM). Springer, 175–192. https://doi.org/10.1007/978-3-031-47705-8_10
- Batty et al. [2013] M. Batty, M. Dodds, and A. Gotsman. 2013. Library abstraction for C/C++ concurrency. In Principles of programming languages (POPL). ACM, 235–248. https://doi.org/10.1145/2429069.2429099
- Batty et al. [2015] M. Batty, K. Memarian, K. Nienhuis, J. Pichon-Pharabod, and P. Sewell. 2015. The problem of programming language concurrency semantics. In Programming Languages and Systems (ESOP). Springer, 283–307. https://doi.org/10.1007/978-3-662-46669-8_12
- Batty et al. [2012] M. Batty, K. Memarian, S. Owens, S. Sarkar, and P. Sewell. 2012. Clarifying and compiling C/C++ concurrency: From C++11 to POWER. In Principles of Programming Languages (POPL). ACM, 509–520. https://doi.org/10.1145/2103656.2103717
- Batty et al. [2011] M. Batty, S. Owens, S. Sarkar, and P. Sewell. 2011. Mathematizing C++ concurrency. In Principles of Programming Languages (POPL). ACM, 55–66. https://doi.org/10.1145/1926385.1926394
- Beck et al. [2023] M. Beck, K. Bhat, L. Stričević, G. Chen, D. Behrens, M. Fu, V. Vafeiadis, H. Chen, and H. Härtig. 2023. AtoMig: Automatically migrating millions lines of code from TSO to WMM. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 61–73. https://doi.org/10.1145/3575693.3579849
- Bender et al. [2015] J. Bender, M. Lesani, and J. Palsberg. 2015. Declarative fence insertion. ACM SIGPLAN Not. 50, 10, 367–385. https://doi.org/10.1145/2858965.2814318
- Bender and Palsberg [2019] J. Bender and J. Palsberg. 2019. A formalization of Java’s concurrent access modes. Proc. ACM Program. Lang. 3, OOPSLA. https://doi.org/10.1145/3360568
- Boehm [2005] H.-J. Boehm. 2005. Threads cannot be implemented as a library. In Programming Language Design and Implementation (PLDI). ACM, 261–268. https://doi.org/10.1145/1065010.1065042
- Boehm [2007] H.-J. Boehm. 2007. Reordering constraints for pthread-style locks. In Principles and Practice of Parallel Programming (PPoPP). ACM, 173–182. https://doi.org/10.1145/1229428.1229470
- Boehm and Adve [2008] H.-J. Boehm and S. V. Adve. 2008. Foundations of the C++ concurrency memory model. In Programming Language Design and Implementation (PLDI). ACM, 68–78. https://doi.org/10.1145/1375581.1375591
- Boehm and Demsky [2014] H.-J. Boehm and B. Demsky. 2014. Outlawing ghosts: Avoiding out-of-thin-air results. In Proc. Workshop on Memory Systems Performance and Correctness (MSPC). ACM, Article 7, 6 pages. https://doi.org/10.1145/2618128.2618134
- Börger and Schulte [1999] E. Börger and W. Schulte. 1999. A programmer friendly modular definition of the semantics of Java. In Formal Syntax and Semantics of Java. Springer, 353–404. https://doi.org/10.1007/3-540-48737-9_10
- Bouajjani [2024] A. Bouajjani. 2024. On verifying concurrent programs under weak consistency models: Decidability and complexity. In Taming the Infinities of Concurrency: Essays Dedicated to Javier Esparza on the Occasion of His 60th Birthday, Stefan Kiefer, Jan Křetínský, and Antonín Kučera (Eds.). Springer, 133–147. https://doi.org/10.1007/978-3-031-56222-8_7
- Bouajjani et al. [2015] A. Bouajjani, G. Calin, E. Derevenetc, and R. Meyer. 2015. Lazy TSO reachability. In Fundamental Approaches to Software Engineering (FASE). 267–282. https://doi.org/10.1007/978-3-662-46675-9_18
- Bouajjani et al. [2017] A. Bouajjani, C. Enea, R. Guerraoui, and J. Hamza. 2017. On verifying causal consistency. In Principles of Programming Languages (POPL). ACM, 626–638. https://doi.org/10.1145/3009837.3009888
- Bouajjani et al. [2011] A. Bouajjani, R. Meyer, and E. Möhlmann. 2011. Deciding robustness against total store ordering. In Automata, Languages and Programming. Springer, 428–440. https://doi.org/10.1007/978-3-642-22012-8_34
- Boudol and Petri [2009] G. Boudol and G. Petri. 2009. Relaxed memory models: An operational approach. In Principles of Programming Languages (POPL). ACM, 392–403. https://doi.org/10.1145/1480881.1480930
- Brookes and O’Hearn [2016] S. Brookes and P. W. O’Hearn. 2016. Concurrent separation logic. ACM SIGLOG News 3, 3, 47–65. https://doi.org/10.1145/2984450.2984457
- Burch and Dill [1994] J. R. Burch and D. L. Dill. 1994. Automatic verification of pipelined microprocessor control. In Computer Aided Verification (CAV). Springer, 68–80. https://doi.org/10.1007/3-540-58179-0_44
- Burckhardt et al. [2007] S. Burckhardt, R. Alur, and M. M. K. Martin. 2007. CheckFence: Checking consistency of concurrent data types on relaxed memory models. In Programming Language Design and Implementation (PLDI). ACM, 12–21. https://doi.org/10.1145/1250734.1250737
- Burckhardt et al. [2012] S. Burckhardt, A. Gotsman, M. Musuvathi, and H. Yang. 2012. Concurrent library correctness on the TSO memory model. In Programming Languages and Systems (ESOP). Springer, 87–107. https://doi.org/10.1007/978-3-642-28869-2_5
- Burckhardt and Musuvathi [2008] S. Burckhardt and M. Musuvathi. 2008. Effective program verification for relaxed memory models. In Computer Aided Verification (CAV). Springer, 107–120. https://doi.org/10.1007/978-3-540-70545-1_12
- Burckhardt et al. [2010] S. Burckhardt, M. Musuvathi, and V. Singh. 2010. Verifying local transformations on relaxed memory models. In Compiler Construction (CC). Springer, 104–123. https://doi.org/10.1007/978-3-642-11970-5_7
- Burnim et al. [2011a] J. Burnim, K. Sen, and C. Stergiou. 2011a. Sound and complete monitoring of sequential consistency for relaxed memory models. In Tools and Algorithms for the Construction and Analysis of Systems (TACAS). Springer, 11–25. https://doi.org/10.1007/978-3-642-19835-9_3
- Burnim et al. [2011b] J. Burnim, K. Sen, and C. Stergiou. 2011b. Testing concurrent programs on relaxed memory models. In International Symposium on Software Testing and Analysis (ISSTA). ACM, 122–132. https://doi.org/10.1145/2001420.2001436
- Cantin et al. [2003] J. F. Cantin, M. H. Lipasti, and J. E. Smith. 2003. The complexity of verifying memory coherence. In Symposium on Parallel Algorithms and Architectures (SPAA). ACM, 254–255. https://doi.org/10.1145/777412.777457
- Cantin et al. [2005] J. F. Cantin, M. H. Lipasti, and J. E. Smith. 2005. The complexity of verifying memory coherence and consistency. IEEE Trans. Parallel Distrib. Syst. 16, 7, 663–671. https://doi.org/10.1109/TPDS.2005.86
- Castellan [2016] S. Castellan. 2016. Weak memory models using event structures. In Journées Francophones des Langages Applicatifs (JFLA). HAL-Inria, 39–54. https://hal.inria.fr/hal-01333582
- Cauligi et al. [2022] S. Cauligi, C. Disselkoen, D. Moghimi, G. Barthe, and D. Stefan. 2022. SoK: Practical foundations for software spectre defenses. In Symposium on Security and Privacy (SP). IEEE, 666–680. https://doi.org/10.1109/SP46214.2022.9833707
- Cenciarelli et al. [1999] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. 1999. An event-based structural operational semantics of multi-threaded Java. In Formal Syntax and Semantics of Java. Springer, 157–200. https://doi.org/10.1007/3-540-48737-9_5
- Cenciarelli et al. [2007] P. Cenciarelli, A. Knapp, and E. Sibilio. 2007. The Java memory model: Operationally, denotationally, axiomatically. In Programming Languages and Systems (ESOP). Springer, 331–346. https://doi.org/10.1007/978-3-540-71316-6_23
- Chakraborty and Vafeiadis [2017] S. Chakraborty and V. Vafeiadis. 2017. Formalizing the concurrency semantics of an LLVM fragment. In IEEE/ACM International Symposium on Code Generation and Optimization (CGO). IEEE, 100–110. https://doi.org/10.1109/CGO.2017.7863732
- Chakraborty and Vafeiadis [2019] S. Chakraborty and V. Vafeiadis. 2019. Grounding thin-air reads with event structures. Proc. ACM Program. Lang. 3, POPL. https://doi.org/10.1145/3290383
- Chini and Saivasan [2020] P. Chini and P. Saivasan. 2020. A framework for consistency algorithms. In IARCS Conference on Foundations of Software Technology and Theoretical Computer Science (FSTTCS). Leibniz-Zentrum für Informatik, 42:1–42:17. https://doi.org/10.4230/LIPIcs.FSTTCS.2020.42
- Cho et al. [2021] K. Cho, S.-H. Lee, A. Raad, and J. Kang. 2021. Revamping hardware persistency models: View-based and axiomatic persistency models for Intel-x86 and Armv8. In Programming Language Design and Implementation (PLDI). ACM, 16–31. https://doi.org/10.1145/3453483.3454027
- Chong et al. [2018] N. Chong, T. Sorensen, and J. Wickerson. 2018. The semantics of transactions and weak memory in x86, Power, ARM, and C++. In Programming Language Design and Implementation (PLDI). ACM, 211–225. https://doi.org/10.1145/3192366.3192373
- Clarke et al. [2001] E. Clarke, A. Biere, R. Raimi, and Y. Zhu. 2001. Bounded model checking using satisfiability solving. Form. Methods Syst. Des. 19, 7–34. https://doi.org/10.1023/A:1011276507260
- Clarke et al. [2004] E. Clarke, D. Kroening, and F. Lerda. 2004. A tool for checking ANSI-C programs. In Tools and Algorithms for the Construction and Analysis of Systems (TACAS). Springer, 168–176. https://doi.org/10.1007/978-3-540-24730-2_15
- Colvin [2021] R. J. Colvin. 2021. Parallelized sequential composition and hardware weak memory models. In Software Engineering and Formal Methods (ICFEM). Springer, 201–221. https://doi.org/10.1007/978-3-030-92124-8_12
- Colvin [2022] R. J. Colvin. 2022. Separation of concerning things: A simpler basis for defining and programming with the C/C++ memory model. In Formal Methods and Software Engineering. 71–89. https://doi.org/10.1007/978-3-031-17244-1_5
- Colvin [2023] R. J. Colvin. 2023. A fine-grained semantics for arrays and pointers under weak memory models. In Formal Methods (FM). Springer, 301–320. https://doi.org/10.1007/978-3-031-27481-7_18
- Colvin et al. [2024] R. J. Colvin, I. J. Hayes, S. Heiner, P. Höfner, L. Meinicke, and R. C. Su. 2024. Practical rely/guarantee verification of an efficient lock for sel4 on multicore architectures. In The Practice of Formal Methods: Essays in Honour of Cliff Jones, Part I, Ana Cavalcanti and James Baxter (Eds.). Springer, 65–87. https://doi.org/10.1007/978-3-031-66676-6_4
- Colvin and Smith [2018] R. J. Colvin and G. Smith. 2018. A wide-spectrum language for verification of programs on weak memory models. In Formal Methods (FM). Springer, 240–257. https://doi.org/10.1007/978-3-319-95582-7_14
- Colvin and Su [2025] R. J. Colvin and R. C. Su. 2025. Structural operational semantics for functional and security verification of pipelined processors. In Computer Aided Verification, Part I (CAV). Springer, 363–388. https://doi.org/10.1007/978-3-031-98668-0_18
- Colvin and Winter [2020] R. J. Colvin and K. Winter. 2020. An abstract semantics of speculative execution for reasoning about security vulnerabilities. In Formal Methods (FM). Springer, 323–341. https://doi.org/10.1007/978-3-030-54997-8_21
- Corella et al. [1993] F. Corella, J. M. Stone, and C. M. Barton. 1993. A formal specification of the PowerPC shared memory architecture. Technical Report RC18638. IBM Thomas J. Watson Research Division.
- Coughlin et al. [2021] N. Coughlin, K. Winter, and G. Smith. 2021. Rely/guarantee reasoning for multicopy atomic weak memory models. In Formal Methods (FM). Springer, 292–310. https://doi.org/10.1007/978-3-030-90870-6_16
- Coughlin et al. [2022] N. Coughlin, K. Winter, and G. Smith. 2022. Compositional reasoning for non-multicopy atomic architectures. Form. Asp. Comput. 35, 2, 1–30. https://doi.org/10.1145/3574137
- Crary and Sullivan [2015] K. Crary and M. J. Sullivan. 2015. A calculus for relaxed memory. In Principles of Programming Languages (POPL). ACM, 623–636. https://doi.org/10.1145/2676726.2676984
- Dalvandi et al. [2020] S. Dalvandi, S. Doherty, B. Dongol, and H. Wehrheim. 2020. Owicki-Gries reasoning for C11-RAR. In Proc. 34st European Conference on Object-Oriented Programming (ECOOP). Leibniz-Zentrum für Informatik, 11:1–11:26. https://doi.org/10.4230/LIPIcs.ECOOP.2020.11
- Dalvandi and Dongol [2021] S. Dalvandi and B. Dongol. 2021. Verifying C11-style weak memory libraries. In Principles and Practice of Parallel Programming (PPoPP). ACM, 451–453. https://doi.org/10.1145/3437801.3441619
- Dalvandi and Dongol [2022] S. Dalvandi and B. Dongol. 2022. Implementing and verifying release-acquire transactional memory in C11. Proc. ACM Program. Lang. 6, OOPSLA2. https://doi.org/10.1145/3563352
- Dalvandi et al. [2022] S. Dalvandi, B. Dongol, S. Doherty, and H. Wehrheim. 2022. Integrating Owicki-Gries for C11-style memory models into Isabelle/HOL. J. Autom. Reason. 66, 141–171. https://doi.org/10.1007/s10817-021-09610-2
- Dan et al. [2013] A. Dan, Y. Meshman, M. Vechev, and E. Yahav. 2013. Predicate abstraction for relaxed memory models. In Static Analysis (SAS). Springer, 84–104. https://doi.org/10.1007/978-3-642-38856-9_7
- Dan et al. [2017] A. Dan, Y. Meshman, M. Vechev, and E. Yahav. 2017. Effective abstractions for verification under relaxed memory models. Comput. Lang. Syst. Struct. 47, 1, 62–76. https://doi.org/10.1016/j.cl.2016.02.003
- Dang et al. [2019] H.-H. Dang, J.-H. Jourdan, J.-O. Kaiser, and D. Dreyer. 2019. RustBelt meets relaxed memory. Proc. ACM Program. Lang. 4, POPL. https://doi.org/10.1145/3371102
- Dang et al. [2022] H.-H. Dang, J. Jung, J. Choi, D.-T. Nguyen, W. Mansky, J. Kang, and D. Dreyer. 2022. Compass: Strong and compositional library specifications in relaxed memory separation logic. In Programming Language Design and Implementation (PLDI). ACM, 792–808. https://doi.org/10.1145/3519939.3523451
- Demsky and Lam [2015] B. Demsky and P. Lam. 2015. SATCheck: SAT-directed stateless model checking for SC and TSO. In Object-Oriented Programming, Systems, Languages, and Applications (OOPSLA). ACM, 20–36. https://doi.org/10.1145/2814270.2814297
- Derevenetc and Meyer [2014] E. Derevenetc and R. Meyer. 2014. Robustness against POWER is PSPACE-complete. In Automata, Languages, and Programming (ICALP). Springer, 158–170. https://doi.org/10.1007/978-3-662-43951-7_14
- Derrick and Smith [2017] J. Derrick and G. Smith. 2017. An observational approach to defining linearizability on weak memory models. In Formal Techniques for Distributed Objects, Components, and Systems (FORTE). Springer, 108–123. https://doi.org/10.1007/978-3-319-60225-7_8
- Derrick et al. [2014] J. Derrick, G. Smith, and B. Dongol. 2014. Verifying linearizability on TSO architectures. In Integrated Formal Methods (IFM). Springer, 341–356. https://doi.org/10.1007/978-3-319-10181-1_21
- Disselkoen et al. [2019] C. Disselkoen, R. Jagadeesan, A. Jeffrey, and J. Riely. 2019. The code that never ran: Modeling attacks on speculative evaluation. In Symposium on Security and Privacy (SP). IEEE, 1238–1255. https://doi.org/10.1109/SP.2019.00047
- Dodds et al. [2018] M. Dodds, M. Batty, and A. Gotsman. 2018. Compositional verification of compiler optimisations on relaxed memory. In Programming Languages and Systems (ESOP). Springer, 1027–1055. https://doi.org/10.1007/978-3-319-89884-1_36
- Doherty et al. [2022] S. Doherty, S. Dalvandi, B. Dongol, and H. Wehrheim. 2022. Unifying operational weak memory verification: An axiomatic approach. ACM Trans. Comput. Logic 23, 4. https://doi.org/10.1145/3545117
- Doherty et al. [2019] S. Doherty, B. Dongol, H. Wehrheim, and J. Derrick. 2019. Verifying C11 programs operationally. In Principles and Practice of Parallel Programming (PPoPP). ACM, 355–365. https://doi.org/10.1145/3293883.3295702
- Doko and Vafeiadis [2016] M. Doko and V. Vafeiadis. 2016. A program logic for C11 memory fences. In Verification, Model Checking, and Abstract Interpretation (VMCAI). Springer, 413–430. https://doi.org/10.1007/978-3-662-49122-5_20
- Doko and Vafeiadis [2017] M. Doko and V. Vafeiadis. 2017. Tackling real-life relaxed concurrency with FSL++. In Programming Languages and Systems (ESOP). Springer, 448–475. https://doi.org/10.1007/978-3-662-54434-1_17
- Dolan et al. [2018] S. Dolan, K. Sivaramakrishnan, and A. Madhavapeddy. 2018. Bounding data races in space and time. In Programming Language Design and Implementation (PLDI). ACM, 242–255. https://doi.org/10.1145/3192366.3192421
- Dongol et al. [2019] B. Dongol, R. Jagadeesan, and J. Riely. 2019. Modular transactions: bounding mixed races in space and time. In Principles and Practice of Parallel Programming (PPoPP). ACM, 82–93. https://doi.org/10.1145/3293883.3295708
- Dubois et al. [1986] M. Dubois, C. Scheurich, and F. Briggs. 1986. Memory access buffering in multiprocessors. SIGARCH Comput. Archit. News 14, 2, 434–442. https://doi.org/10.1145/17356.17406
- Dvir et al. [2024] Y. Dvir, O. Kammar, and O. Lahav. 2024. A denotational approach to release/acquire concurrency. In Programming Languages and Systems (ESOP). 121–149. https://doi.org/10.1007/978-3-031-57267-8_5
- Emmi and Enea [2019] M. Emmi and C. Enea. 2019. Weak-consistency specification via visibility relaxation. Proc. ACM Program. Lang. 3, POPL. https://doi.org/10.1145/3290373
- Fan et al. [2023] H. Fan, Z. Sun, and F. He. 2023. Satisfiability modulo ordering consistency theory for SC, TSO, and PSO memory models. ACM Trans. Program. Lang. Syst. 45, 1. https://doi.org/10.1145/3579835
- Fang et al. [2003] X. Fang, J. Lee, and S. P. Midkiff. 2003. Automatic fence insertion for shared memory multiprocessing. In International Conference on Supercomputing (ICS). ACM, 285–294. https://doi.org/10.1145/782814.782854
- Ferreira et al. [2010] R. Ferreira, X. Feng, and Z. Shao. 2010. Parameterized memory models and concurrent separation logic. In Programming Languages and Systems (ESOP). Springer, 267–286. https://doi.org/10.1007/978-3-642-11957-6_15
- Flur et al. [2016] S. Flur, K. E. Gray, C. Pulte, S. Sarkar, A. Sezgin, L. Maranget, W. Deacon, and P. Sewell. 2016. Modelling the ARMv8 architecture, operationally: Concurrency and ISA. In Principles of Programming Languages (POPL). ACM, 608–621. https://doi.org/10.1145/2837614.2837615
- Fox et al. [2023] A. C. J. Fox, G. Stockwell, S. Xiong, H. Becker, D. P. Mulligan, G. Petri, and N. Chong. 2023. A verification methodology for the Arm® confidential computing architecture: From a secure specification to safe implementations. Proc. ACM Program. Lang. 7, OOPSLA1. https://doi.org/10.1145/3586040
- Furbach et al. [2015] F. Furbach, R. Meyer, K. Schneider, and M. Senftleben. 2015. Memory-model-aware testing: A unified complexity analysis. ACM Trans. Embed. Comput. Syst. 14, 4. https://doi.org/10.1145/2753761
- Gavrilenko et al. [2019] N. Gavrilenko, H. Ponce de León, F. Furbach, K. Heljanko, and R. Meyer. 2019. BMC for weak memory models: Relation analysis for compact SMT encodings. In Computer Aided Verification (CAV). Springer, 355–365. https://doi.org/10.1007/978-3-030-25540-4_19
- Geeson and Smith [2024] L. Geeson and L. Smith. 2024. Compiler testing with relaxed memory models. In International Symposium on Code Generation and Optimization (CGO). IEEE, 334–348. https://doi.org/10.1109/CGO57630.2024.10444836
- Gharachorloo et al. [1990] K. Gharachorloo, D. Lenoski, J. Laudon, P. Gibbons, A. Gupta, and J. Hennessy. 1990. Memory consistency and event ordering in scalable shared-memory multiprocessors. In International Symposium on Computer Architecture (ISCA). ACM, 15–26. https://doi.org/10.1145/325164.325102
- Gibbons and Korach [1997] P. B. Gibbons and E. Korach. 1997. Testing shared memories. SIAM J. Comput. 26, 4, 1208–1244. https://doi.org/10.1137/S0097539794279614
- Godefroid [1997] P. Godefroid. 1997. Model checking for programming languages using VeriSoft. In Principles of Programming Languages (POPL). ACM, 174–186. https://doi.org/10.1145/263699.263717
- Goens et al. [2023] A. Goens, S. Chakraborty, S. Sarkar, S. Agarwal, N. Oswald, and V. Nagarajan. 2023. Compound memory models. Proc. ACM Program. Lang. 7, PLDI. https://doi.org/10.1145/3591267
- Gogte et al. [2022] V. Gogte, A. Kolli, and T. F. Wenisch. 2022. A Primer on Memory Persistency. Morgan and Claypool. https://doi.org/10.1007/978-3-031-79205-2
- Gontmakher and Schuster [1998] A. Gontmakher and A. Schuster. 1998. Characterizations for Java memory behavior. In Proc. First Merged International Parallel Processing Symposium and Symposium on Parallel and Distributed Processing. IEEE, 682–686. https://doi.org/10.1109/IPPS.1998.670000
- González et al. [2011] A. González, F. Latorre, and G. Magklis. 2011. Processor Microarchitecture: An Implementation Perspective. Morgan and Claypool. https://doi.org/10.1007/978-3-031-01729-2
- Goodman [1991] J. R. Goodman. 1991. Cache consistency and sequential consistency. Technical Report 1006. Computer Sciences Department, University of Wisconsin-Madison.
- Gopalakrishnan and Verbrugge [2020] A. Gopalakrishnan and C. Verbrugge. 2020. Reordering under the ECMAScript memory consistency model. In Languages and Compilers for Parallel Computing (LCPC). 198–214. https://doi.org/10.1007/978-3-030-95953-1_14
- Gopalakrishnan et al. [2023] A. Gopalakrishnan, C. Verbrugge, and M. Batty. 2023. Memory consistency models for program transformations: An intellectual abstract. In International Symposium on Memory Management (ISMM). ACM, 30–42. https://doi.org/10.1145/3591195.3595274
- Gosling et al. [1996] J. Gosling, B. Joy, and G. Steele. 1996. The Java Language Specification. Addison-Wesley.
- Gouicem et al. [2022] R. Gouicem, D. Sprokholt, J. Ruehl, R. C. O. Rocha, T. Spink, S. Chakraborty, and P. Bhatotia. 2022. Risotto: A dynamic binary translator for weak memory model architectures. In Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS). ACM, 107–122. https://doi.org/10.1145/3567955.3567962
- Griffin and Dongol [2021] M. Griffin and B. Dongol. 2021. Verifying secure speculation in Isabelle/HOL. In Formal Methods (FM). Springer, 43–60. https://doi.org/10.1007/978-3-030-90870-6_3
- Guerraoui and Kapatka [2010] R. Guerraoui and M. Kapatka. 2010. Principles of Transactional Memory. Morgan and Claypool. https://doi.org/10.1007/978-3-031-79205-2
- Haas et al. [2023] T. Haas, R. Maseli, R. Meyer, and H. Ponce de León. 2023. Static analysis of memory models for SMT encodings. Proc. ACM Program. Lang. 7, OOPSLA2. https://doi.org/10.1145/3622855
- Haas et al. [2022] T. Haas, R. Meyer, and H. Ponce de León. 2022. CAAT: Consistency as a theory. Proc. ACM Program. Lang. 6, OOPSLA2. https://doi.org/10.1145/3563292
- Hammond et al. [2024] A. Hammond, Z. Liu, T. Pérami, P. Sewell, L. Birkedal, and J. Pichon-Pharabod. 2024. An axiomatic basis for computer programming on the relaxed Arm-A architecture: The AxSL logic. Proc. ACM Program. Lang. 8, POPL. https://doi.org/10.1145/3632863
- Hangal et al. [2004] S. Hangal, D. Vahia, C. Manovit, and J.-Y. J. Lu. 2004. TSOtool: A program for verifying memory systems using the memory consistency model. In International Symposium on Computer Architecture (ISCA). IEEE, 114–124. https://doi.org/10.1109/ISCA.2004.1310768
- Harman [2007] N. A. Harman. 2007. Algebraic models of behaviour and correctness of smt and cmt processors. J. Log. Algebr. Methods Program. 74, 1, 32–56. https://doi.org/10.1016/j.jlap.2007.07.001
- Hayes et al. [2016] I. J. Hayes, R. J. Colvin, L. A. Meinicke, K. Winter, and A. Velykis. 2016. An algebra of synchronous atomic steps. In Formal Methods (FM). Springer, 352–369. https://doi.org/10.1007/978-3-319-48989-6_22
- Hennessy and Patterson [2011] J. L. Hennessy and D. A. Patterson. 2011. Computer architecture: a quantitative approach. Elsevier.
- Higham and Kawash [1998] L. Higham and J. Kawash. 1998. Java: Memory consistency and process coordination. In Distributed Computing. Springer, 201–215.
- Hoare et al. [2009] C. A. R. T. Hoare, B. Möller, G. Struth, and I. Wehrman. 2009. Concurrent Kleene algebra. In Concurrency Theory (CONCUR). Springer, 399–414. https://doi.org/10.1007/978-3-642-04081-8_27
- Hóu et al. [2021] Z. Hóu, D. Sanan, A. Tiu, Y. Liu, K. C. Hoa, and J. S. Dong. 2021. An Isabelle/HOL formalisation of the SPARC instruction set architecture and the TSO memory model. J. Autom. Reasoning 65, 569–598. https://doi.org/10.1007/s10817-020-09579-4
- Iorga et al. [2021] D. Iorga, A. Donaldson, T. Sorensen, and J. Wickerson. 2021. The semantics of shared memory in Intel CPU/FPGA systems. Proc. ACM Program. Lang. 5, OOPSLA. https://doi.org/10.1145/3485497
- Ishtiaq and O’Hearn [2001] S. S. Ishtiaq and P. W. O’Hearn. 2001. BI as an assertion language for mutable data structures. In Principles of Programming Languages (POPL). ACM, 14–26.
- Jagadeesan et al. [2020] R. Jagadeesan, A. Jeffrey, and J. Riely. 2020. Pomsets with preconditions: A simple model of relaxed memory. Proc. ACM Program. Lang. 4, OOPSLA. https://doi.org/10.1145/3428262
- Jagadeesan et al. [2010] R. Jagadeesan, C. Pitcher, and J. Riely. 2010. Generative operational semantics for relaxed memory models. In Programming Languages and Systems (ESOP). Springer, 307–326. https://doi.org/10.1007/978-3-642-11957-6_17
- Jeffrey and Riely [2016] A. Jeffrey and J. Riely. 2016. On thin air reads: towards an event structures model of relaxed memory. In Logic in Computer Science (LICS). ACM, 759–767. https://doi.org/10.1145/2933575.2934536
- Jeffrey et al. [2022] A. Jeffrey, J. Riely, M. Batty, S. Cooksey, I. Kaysin, and A. Podkopaev. 2022. The leaky semicolon: Compositional semantic dependencies for relaxed-memory concurrency. Proc. ACM Program. Lang. 6, POPL. https://doi.org/10.1145/3498716
- Jones [1983a] C. B. Jones. 1983a. Specification and design of (parallel) programs. In IFIP Congress. 321–332.
- Jones [1983b] C. B. Jones. 1983b. Tentative steps toward a development method for interfering programs. ACM Trans. Program. Lang. Syst. 5, 4, 596–619. https://doi.org/10.1145/69575.69577
- Jones and Astarte [2018] C. B. Jones and T. K. Astarte. 2018. Challenges for formal semantic description: Responses from the main approaches. In Engineering Trustworthy Software Systems. Springer, 176–217.
- Jones et al. [2002] R. B. Jones, J. U. Skakkebæk, and D. L. Dill. 2002. Formal verification of out-of-order execution with incremental flushing. Form. Methods Syst. Des. 20, 139–158. https://doi.org/10.1023/A:1014118529369
- Joshi and Kroening [2015] S. Joshi and D. Kroening. 2015. Property-driven fence insertion using reorder bounded model checking. In Formal Methods (FM). Springer, 291–307. https://doi.org/10.1007/978-3-319-19249-9_19
- Jung et al. [2018] R. Jung, R. Krebbers, J.-H. Jourdan, A. Bizjak, L. Birkedal, and D. Dreyer. 2018. Iris from the gound up: A modular foundation for higher-order concurrent separation logic. J. Funct. Program. 28. https://doi.org/10.1017/S0956796818000151
- Jung et al. [2015] R. Jung, D. Swasey, F. Sieczkowski, K. Svendsen, A. Turon, L. Birkedal, and D. Dreyer. 2015. Iris: Monoids and invariants as an orthogonal basis for concurrent reasoning. In Principles of Programming Languages (POPL). ACM, 637–650. https://doi.org/10.1145/2775051.2676980
- Kaiser et al. [2017] J.-O. Kaiser, H.-H. Dang, D. Dreyer, O. Lahav, and V. Vafeiadis. 2017. Strong logic for weak memory: Reasoning about release-acquire consistency in Iris. In European Conference on Object-Oriented Programming (ECOOP). Leibniz-Zentrum für Informatik, Article 17, 29 pages. https://doi.org/10.4230/LIPIcs.ECOOP.2017.17
- Kang et al. [2017] J. Kang, C.-K. Hur, O. Lahav, V. Vafeiadis, and D. Dreyer. 2017. A promising semantics for relaxed-memory concurrency. In Principles of Programming Languages (POPL). ACM, 175–189. https://doi.org/10.1145/3009837.3009850
- Kasikci et al. [2015] B. Kasikci, C. Zamfir, and G. Candea. 2015. Automated classification of data races under both strong and weak memory models. ACM Trans. Program. Lang. Syst. 37, 3. https://doi.org/10.1145/2734118
- Khronos OpenCL Working Group [2020] Khronos OpenCL Working Group. 2020. The OpenCL Specification (3rd ed.). https://registry.khronos.org/OpenCL
- Khyzha and Lahav [2021] A. Khyzha and O. Lahav. 2021. Taming x86-TSO persistency. Proc. ACM Program. Lang. 5, POPL. https://doi.org/10.1145/3434328
- Klein et al. [2009] G. Klein, K. Elphinstone, G. Heiser, J. Andronick, D. Cock, P. Derrin, D. Elkaduwe, K. Engelhardt, R. Kolanski, M. Norrish, T. Sewell, H. Tuch, and S. Winwood. 2009. seL4: Formal verification of an OS kernel. In Symposium on Operating Systems Principles (SOSP). ACM, 207–220. https://doi.org/10.1145/1629575.1629596
- Klimis et al. [2024] V. Klimis, A. F. Donaldson, V. Vafeiadis, J. Wickerson, and A. Raad. 2024. Challenges in empirically testing memory persistency models. In International Conference on Software Engineering: New Ideas and Emerging Results (ICSE-NIER). ACM, 82–86. https://doi.org/10.1145/3639476.3639765
- Kocher et al. [2019] P. Kocher, J. Horn, A. Fogh, D. Genkin, D. Gruss, W. Haas, M. Hamburg, M. Lipp, S. Mangard, T. Prescher, M. Schwarz, and Y. Yarom. 2019. Spectre attacks: Exploiting speculative execution. In Security and Privacy (SP). IEEE, 1–19.
- Kokologiannakis et al. [2023] M. Kokologiannakis, O. Lahav, and V. Vafeiadis. 2023. Kater: Automating weak memory model metatheory and consistency checking. Proc. ACM Program. Lang. 7, POPL. https://doi.org/10.1145/3571212
- Kokologiannakis et al. [2018] M. Kokologiannakis, A. Raad, and V. Vafeiadis. 2018. Effective stateless model checking for C/C++ concurrency. Proc. ACM Program. Lang. 2, POPL. https://doi.org/10.1145/3158105
- Kokologiannakis and Vafeiadis [2020] M. Kokologiannakis and V. Vafeiadis. 2020. HMC: Model checking for hardware memory models. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 1157–1171. https://doi.org/10.1145/3373376.3378480
- Kokologiannakis and Vafeiadis [2021] M. Kokologiannakis and V. Vafeiadis. 2021. GenMC: A model checker for weak memory models. In Computer Aided Verification (CAV). Springer, 427–440. https://doi.org/10.1007/978-3-030-81685-8_20
- Krishna et al. [2022] S. Krishna, A. Godbole, R. Meyer, and S. Chakraborty. 2022. Parameterized verification under release acquire is PSPACE-complete. In Principles of Distributed Computing (PODC). ACM, 482–492. https://doi.org/10.1145/3519270.3538445
- Kuperstein et al. [2012] M. Kuperstein, M. Vechev, and E. Yahav. 2012. Automatic inference of memory fences. ACM SIGACT News 43, 108–123. Issue 2. https://doi.org/10.1145/2261417.2261438
- Lahav and Boker [2022] O. Lahav and U. Boker. 2022. What’s decidable about causally consistent shared memory? ACM Trans. Program. Lang. Syst. 44, 2. https://doi.org/10.1145/3505273
- Lahav et al. [2023] O. Lahav, B. Dongol, and H. Wehrheim. 2023. Rely-guarantee reasoning for causally consistent shared memory. In Computer Aided Verification (CAV). Springer, 206–229. https://doi.org/10.1007/978-3-031-37706-8_11
- Lahav et al. [2021] O. Lahav, E. Namakonov, J. Oberhauser, A. Podkopaev, and V. Vafeiadis. 2021. Making weak memory models fair. Proc. ACM Program. Lang. 5, OOPSLA. https://doi.org/10.1145/3485475
- Lahav and Vafeiadis [2015] O. Lahav and V. Vafeiadis. 2015. Owicki-Gries reasoning for weak memory models. In Automata, Languages, and Programming (ICALP). Springer, 311–323. https://doi.org/10.1007/978-3-662-47666-6_25
- Lahav and Vafeiadis [2016] O. Lahav and V. Vafeiadis. 2016. Explaining relaxed memory models with program transformations. In Formal Methods (FM). Springer, 479–495. https://doi.org/10.1007/978-3-319-48989-6_29
- Lahav et al. [2017] O. Lahav, V. Vafeiadis, J. Kang, C.-K. Hur, and D. Dreyer. 2017. Repairing sequential consistency in C/C++11. In Programming Language Design and Implementation (PLDI). ACM, 618–632. https://doi.org/10.1145/3062341.3062352
- Lamport [1979] Lamport. 1979. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Trans. Comput. C-28, 9, 690–691.
- Lee and Padua [2000] J. Lee and D. A. Padua. 2000. Hiding relaxed memory consistency with compilers. In Parallel Architectures and Compilation Techniques (PACT). IEEE, 111. https://doi.org/10.1109/PACT.2000.888336
- Lee et al. [2023] S.-H. Lee, M. Cho, R. Margalit, C.-K. Hur, and O. Lahav. 2023. Putting weak memory in order via a promising intermediate representation. Proc. ACM Program. Lang. 7, PLDI. https://doi.org/10.1145/3591297
- Lee et al. [2020] S.-H. Lee, M. Cho, A. Podkopaev, S. Chakraborty, C.-K. Hur, O. Lahav, and V. Vafeiadis. 2020. Promising 2.0: Global optimizations in relaxed memory concurrency. In Programming Language Design and Implementation. ACM, 362–376. https://doi.org/10.1145/3385412.3386010
- Lenoski et al. [1992] D. Lenoski, J. Laudon, K. Gharachorloo, W. . Weber, A. Gupta, J. Hennessy, M. Horowitz, and M. S. Lam. 1992. The Stanford Dash multiprocessor. Computer 25, 3, 63–79.
- Lepigre et al. [2022] R. Lepigre, M. Sammler, K. Memarian, R. Krebbers, D. Dreyer, and P. Sewell. 2022. VIP: Verifying real-world C idioms with integer-pointer casts. Proc. ACM Program. Lang. 6, POPL. https://doi.org/10.1145/3498681
- Linden and Wolper [2010] A. Linden and P. Wolper. 2010. An automata-based symbolic approach for verifying programs on relaxed memory models. In Model Checking Software (SPIN). Springer, 212–226. https://doi.org/10.1007/978-3-642-16164-3_16
- Linden and Wolper [2011] A. Linden and P. Wolper. 2011. A verification-based approach to memory fence insertion in relaxed memory systems. In Model Checking Software (SPIN). Springer, 144–160. https://doi.org/10.1007/978-3-642-22306-8_10
- Linden and Wolper [2013] A. Linden and P. Wolper. 2013. A verification-based approach to memory fence insertion in PSO memory systems. In Tools and Algorithms for the Construction and Analysis of Systems (TACAS). Springer, 339–353. https://doi.org/10.1007/978-3-642-36742-7_24
- Lipp et al. [2018] M. Lipp, M. Schwarz, D. Gruss, T. Prescher, W. Haas, A. Fogh, J. Horn, S. Mangard, P. Kocher, D. Genkin, Y. Yarom, and M. Hamburg. 2018. Meltdown: Reading kernel memory from user space. In 27th USENIX Security Symposium (USENIX Security 18).
- Lipton and Sandberg [1988] R. J. Lipton and J. S. Sandberg. 1988. PRAM: A Scalable Shared Memory. Technical Report 180-88. Department of Computer Science, Princeton University.
- Lochbihler [2013] A. Lochbihler. 2013. Making the Java memory model safe. ACM Trans. Program. Lang. Syst. 35, 4. https://doi.org/10.1145/2518191
- Ltd. [2024] A. Ltd. 2024. Arm® Architecture Reference Manual (for A-profile architecture). https://developer.arm.com/documentation/ddi0487/la/ Document number: ARM DDI 0487.
- Lustig et al. [2014] D. Lustig, M. Pellauer, and M. Martonosi. 2014. PipeCheck: Specifying and verifying microarchitectural enforcement of memory consistency models. In Proc. 47th International Symposium on Microarchitecture. 635–646.
- Lustig et al. [2019] D. Lustig, S. Sahasrabuddhe, and O. Giroux. 2019. A formal analysis of the NVIDIA PTX memory consistency model. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 257–270. https://doi.org/10.1145/3297858.3304043
- Lustig et al. [2016] D. Lustig, G. Sethi, M. Martonosi, and A. Bhattacharjee. 2016. COATCheck: Verifying memory ordering at the hardware-OS interface. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 233–247. https://doi.org/10.1145/2872362.2872399
- Lustig et al. [2017] D. Lustig, A. Wright, A. Papakonstantinou, and O. Giroux. 2017. Automated synthesis of comprehensive memory model litmus test suites. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 661–675. https://doi.org/10.1145/3037697.3037723
- Mador-Haim et al. [2010] S. Mador-Haim, R. Alur, and M. M. K. Martin. 2010. Generating litmus tests for contrasting memory consistency models. In Computer Aided Verification (CAV). Springer, 273–287. https://doi.org/10.1007/978-3-642-14295-6_26
- Mador-Haim et al. [2011] S. Mador-Haim, R. Alur, and M. M. K. Martin. 2011. Litmus tests for comparing memory consistency models: How long do they need to be?. In Proc. 48th Design Automation Conference (DAC). ACM, 504–509. https://doi.org/10.1145/2024724.2024842
- Maessen et al. [2000] J.-W. Maessen, Arvind, and X. Shen. 2000. Improving the Java memory model using CRF. In OOPSLA. ACM, 1–12. https://doi.org/10.1145/354222.353172
- Manerkar et al. [2018] Y. A. Manerkar, D. Lustig, M. Martonosi, and A. Gupta. 2018. PipeProof: Automated memory consistency proofs for microarchitectural specifications. In Proc. 51st International Symposium on Microarchitecture. 788–801. https://doi.org/10.1109/MICRO.2018.00069
- Manerkar et al. [2015] Y. A. Manerkar, D. Lustig, M. Pellauer, and M. Martonosi. 2015. CCICheck: using hb graphs to verify the coherence-consistency interface. In Proc. 48th International Symposium on Microarchitecture. ACM, 26–37. https://doi.org/10.1145/2830772.2830782
- Manson and Pugh [2001] J. Manson and W. Pugh. 2001. Core semantics of multithreaded Java. In Proc. 2001 Joint ACM-ISCOPE Conference on Java Grande (JGI). ACM, 29–38. https://doi.org/10.1145/376656.376806
- Manson and Pugh [2004] J. Manson and W. Pugh. 2004. Requirements for programming language memory models. In PODC Workshop on Concurrency and Synchronization in Java Programs. ACM, 36–45.
- Manson et al. [2005] J. Manson, W. Pugh, and S. V. Adve. 2005. The Java memory model. In Principles of Programming Languages (POPL). ACM, 378–391. https://doi.org/10.1145/1040305.1040336
- Maranget et al. [2012] L. Maranget, S. Sarkar, and P. Sewell. 2012. A Tutorial Introduction to the ARM and POWER Relaxed Memory Models. https://www.cl.cam.ac.uk/~pes20/ppc-supplemental/test7.pdf
- McKenney [2005a] P. E. McKenney. 2005a. Memory Ordering in Modern Microprocessors, Part I. Linux Journal. Retrieved May 10, 2023 from https://www.linuxjournal.com/article/8211
- McKenney [2005b] P. E. McKenney. 2005b. Memory Ordering in Modern Microprocessors, Part II. Linux Journal. Retrieved May 10, 2023 from https://www.linuxjournal.com/article/8212
- McKenney [2010] P. E. McKenney. 2010. Memory barriers: a hardware view for software hackers. Linux Technology Center, IBM Beaverton.
- Memarian et al. [2016] K. Memarian, J. Matthiesen, J. Lingard, K. Nienhuis, D. Chisnall, R. Watson, and P. Sewell. 2016. Into the depths of C: Elaborating the de factor standards. In Programming Language Design and Implementation (PLDI). ACM, 1–15. https://doi.org/10.1145/2908080.2908081
- Meshman et al. [2014] Y. Meshman, A. Dan, M. Vechev, and E. Yahav. 2014. Synthesis of memory fences via refinement propagation. In Static Analysis (SAS). Springer, 237–252. https://doi.org/10.1007/978-3-319-10936-7_15
- Meshman et al. [2015] Y. Meshman, N. Rinetzky, and E. Yahav. 2015. Pattern-based synthesis of synchronization for the C++ memory model. In Formal Methods in Computer-Aided Design (FMCAD). 120–127. https://doi.org/10.1109/FMCAD.2015.7542261
- Mével and Jourdan [2021] G. Mével and J.-H. Jourdan. 2021. Formal verification of a concurrent bounded queue in a weak memory model. Proc. ACM Program. Lang. 5, ICFP. https://doi.org/10.1145/3473571
- Mével et al. [2020] G. Mével, J.-H. Jourdan, and F. Pottier. 2020. Cosmo: a concurrent separation logic for multicore OCaml. Proc. ACM Program. Lang. 4, ICFP. https://doi.org/10.1145/3408978
- Moiseenko et al. [2022] E. Moiseenko, M. Kokologiannakis, and V. Vafeiadis. 2022. Model checking for a multi-execution memory model. Proc. ACM Program. Lang. 6, OOPSLA2. https://doi.org/10.1145/3563315
- Moiseenko et al. [2021] E. Moiseenko, A. Podkopaev, and D. Koznov. 2021. A survey of programming language memory models. Programming and Computer Software 47, 439–456. https://doi.org/10.1134/S0361768821060050
- Nagarajan et al. [2020] V. Nagarajan, D. J. Sorin, M. D. Hill, and D. A. Wood. 2020. A Primer on Memory Consistency and Cache Coherence (2nd ed.). Morgan and Claypool. https://doi.org/10.1007/978-3-031-01733-9
- Nienhuis et al. [2016] K. Nienhuis, K. Memarian, and P. Sewell. 2016. An operational semantics for C/C++ 11 concurrency. In OOPSLA. ACM, 111–128. https://doi.org/10.1145/2983990.2983997
- Norris and Demsky [2013] B. Norris and B. Demsky. 2013. CDSChecker: Checking concurrent data structures written with C/C++ atomics. In OOPSLA. ACM, 131–150. https://doi.org/10.1145/2509136.2509514
- Norris and Demsky [2016] B. Norris and B. Demsky. 2016. A practical approach for model checking C/C++11 code. ACM Trans. Program. Lang. Syst. 38, 3. https://doi.org/10.1145/2806886
- NVIDIA Ltd [2024] NVIDIA Ltd. 2024. Parallel Thread Execution ISA Version 8.5 (Memory Consistency Model). https://docs.nvidia.com/cuda/parallel-thread-execution/index.html#memory-consistency-model
- Oberhauser et al. [2021] J. Oberhauser, R. L. de Lima Chehab, D. Behrens, M. Fu, A. Paolillo, L. Oberhauser, K. Bhat, Y. Wen, H. Chen, J. Kim, and V. Vafeiadis. 2021. VSync: push-button verification and optimization for synchronization primitives on weak memory models. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 530–545. https://doi.org/10.1145/3445814.3446748
- Ou and Demsky [2018] P. Ou and B. Demsky. 2018. Towards understanding the costs of avoiding out-of-thin-air results. Proc. ACM Program. Lang. 2, OOPSLA. https://doi.org/10.1145/3276506
- Owens et al. [2009] S. Owens, S. Sarkar, and P. Sewell. 2009. A better x86 memory model: x86-TSO. In Theorem Proving in Higher Order Logics (TPHOL). Springer, 391–407. https://doi.org/10.1007/978-3-642-03359-9_27
- Owicki and Gries [1976] S. Owicki and D. Gries. 1976. An axiomatic proof technique for parallel programs I. Acta Inform. 6, 4, 319–340. https://doi.org/10.1007/BF00268134
- Paolillo et al. [2021] A. Paolillo, H. Ponce de León, T. Haas, D. Behrens, R. Chehab, M. Fu, and R. Meyer. 2021. Verifying and optimizing compact NUMA-aware locks on weak memory models. arXiv:2111.15240 [cs.OS]
- Patterson and Hennessy [2016] D. A. Patterson and J. L. Hennessy. 2016. Computer organization and design ARM edition: the hardware software interface. Morgan Kaufmann.
- Paviotti et al. [2020] M. Paviotti, S. Cooksey, A. Paradis, D. Wright, S. Owens, and M. Batty. 2020. Modular relaxed dependencies in weak memory concurrency. In Programming Languages and Systems (ESOP). Springer, 599–625. https://doi.org/10.1007/978-3-030-44914-8_22
- Pichon-Pharabod and Sewell [2016] J. Pichon-Pharabod and P. Sewell. 2016. A concurrency semantics for relaxed atomics that permits optimisation and avoids thin-air executions. In Principles of Programming Languages (POPL). ACM, 622–633. https://doi.org/10.1145/2837614.2837616
- Podkopaev et al. [2019] A. Podkopaev, O. Lahav, and V. Vafeiadis. 2019. Bridging the gap between programming languages and hardware memory models. Proc. ACM Program. Lang. 3, POPL. https://doi.org/10.1145/3290382
- Podkopaev et al. [2016] A. Podkopaev, I. Sergey, and A. Nanevski. 2016. Operational aspects of C/C++ concurrency. arXiv:1606.01400 [cs.PL]
- Ponce de León et al. [2018] H. Ponce de León, F. Furbach, K. Heljanko, and R. Meyer. 2018. BMC with memory models as modules. In Formal Methods in Computer-Aided Design (FMCAD). IEEE, 1–9. https://doi.org/10.23919/FMCAD.2018.8603021
- Ponce de León and Kinder [2022] H. Ponce de León and J. Kinder. 2022. Cats vs. Spectre: An axiomatic approach to modeling speculative execution attacks. In Security and Privacy (SP). IEEE, 235–248. https://doi.org/10.1109/SP46214.2022.9833774
- Pratt [1986] V. R. Pratt. 1986. Modelling in concurrency with partial orders. Int. J. Parallel Prog. 15, 33–71. https://doi.org/10.1007/BF01379149
- Pugh [1999] W. Pugh. 1999. Fixing the Java memory model. In Proc. ACM 1999 Conference on Java Grande (JAVA). ACM, 89–98. https://doi.org/10.1145/304065.304106
- Pugh [2000] W. Pugh. 2000. The Java memory model is fatally flawed. Concurrency: Practice and Experience 12, 6, 445–455. https://doi.org/10.1002/1096-9128
- Pulte et al. [2017] C. Pulte, S. Flur, W. Deacon, J. French, S. Sarkar, and P. Sewell. 2017. Simplifying ARM concurrency: Multicopy-atomic axiomatic and operational models for ARMv8. Proc. ACM Program. Lang. 2, POPL. https://doi.org/10.1145/3158107
- Pulte et al. [2019] C. Pulte, J. Pichon-Pharabod, J. Kang, S.-H. Lee, and C.-K. Hur. 2019. Promising-ARM/RISC-V: A simpler and faster operational concurrency model. In Programming Language Design and Implementation (PLDI). ACM, 1–15. https://doi.org/10.1145/3314221.3314624
- Raad et al. [2019a] A. Raad, M. Doko, L. Rožić, O. Lahav, and V. Vafeiadis. 2019a. On library correctness under weak memory consistency: Specifying and verifying concurrent libraries under declarative consistency models. POPL. https://doi.org/10.1145/3290381
- Raad et al. [2020] A. Raad, O. Lahav, and V. Vafeiadis. 2020. Persistent Owicki-Gries reasoning: A program logic for reasoning about persistent programs on Intel-x86. Proc. ACM Program. Lang. 4, OOPSLA. https://doi.org/10.1145/3428219
- Raad et al. [2022] A. Raad, L. Maranget, and V. Vafeiadis. 2022. Extending intel-x86 consistency and persistency: formalising the semantics of intel-x86 memory types and non-temporal stores. Proc. ACM Program. Lang. 6, POPL. https://doi.org/10.1145/3498683
- Raad et al. [2019b] A. Raad, J. Wickerson, G. Neiger, and V. Vafeiadis. 2019b. Persistency semantics of the Intel-x86 architecture. Proc. ACM Program. Lang. 4, POPL. https://doi.org/10.1145/3371079
- Rau and Fisher [1993] B. R. Rau and J. A. Fisher. 1993. Instruction-level parallel processing: History, overview, and perspective. In Instruction-Level Parallelism: A Special Issue of The Journal of Supercomputing. Springer, 9–50. https://doi.org/10.1007/978-1-4615-3200-2_3
- Reynolds [2002] J. C. Reynolds. 2002. Separation logic: A logic for shared mutable data structures. In Logic in Computer Science (LICS). IEEE, 55–74. https://doi.org/10.1109/LICS.2002.1029817
- Ridge [2010] T. Ridge. 2010. A rely-guarantee proof system for x86-TSO. In Verified Software: Theories, Tools, Experiments (VSTTE). Springer, 55–70. https://doi.org/10.1007/978-3-642-15057-9_4
- RISC-V Foundation [2019] RISC-V Foundation. 2019. The RISC-V Instruction Set Manual, Volume I: User-Level ISA, Document Version 20191213.
- Ritson and Owens [2016] C. G. Ritson and S. Owens. 2016. Benchmarking weak memory models. In Principles and Practice of Parallel Programming (PPoPP). ACM, 1–11. https://doi.org/10.1145/2851141.2851150
- Rocha et al. [2022] R. C. O. Rocha, D. Sprokholt, M. Fink, R. Gouicem, T. Spink, S. Chakraborty, and P. Bhatotia. 2022. Lasagne: a static binary translator for weak memory model architectures. In Programming Language Design and Implementation (PLDI). ACM, 888–902. https://doi.org/10.1145/3519939.3523719
- Roychoudhury and Mitra [2002] A. Roychoudhury and T. Mitra. 2002. Specifying multithreaded Java semantics for program verification. In Proc. 24th International Conference on Software Engineering (ICSE). ACM, 489–499. https://doi.org/10.1145/581339.581399
- Sarkar et al. [2012] S. Sarkar, K. Memarian, S. Owens, M. Batty, P. Sewell, L. Maranget, J. Alglave, and D. Williams. 2012. Synchronising C/C++ and POWER. In Programming Language Design and Implementation (PLDI). ACM, 311–322. https://doi.org/10.1145/2254064.2254102
- Sarkar et al. [2011] S. Sarkar, P. Sewell, J. Alglave, L. Maranget, and D. Williams. 2011. Understanding POWER multiprocessors. In Programming Language Design and Implementation (PLDI). ACM, 175–186. https://doi.org/10.1145/1993498.1993520
- Sarkar et al. [2009] S. Sarkar, P. Sewell, F. Zappa Nardelli, S. Owens, T. Ridge, T. Braibant, M. Myreen, and J. Alglave. 2009. The semantics of x86-CC multiprocessor machine code. In Principles of Programming Languages (POPL). ACM, 379–391. https://doi.org/10.1145/1480881.1480929
- Senftleben and Schneider [2018a] M. Senftleben and K. Schneider. 2018a. Operational characterization of weak memory consistency models. In Architecture of Computing Systems (ARCS). Springer, 195–208. https://doi.org/10.1007/978-3-319-77610-1_15
- Senftleben and Schneider [2018b] M. Senftleben and K. Schneider. 2018b. Using temporal logics for specifying weak memory consistency models. Int. J. Crit. Comput.-Based Syst. 8, 2, 214–229. https://doi.org/10.1504/IJCCBS.2018.096187
- Ševčík and Aspinall [2008] J. Ševčík and D. Aspinall. 2008. On validity of program transformations in the Java memory model. In ECOOP 2008 – Object-Oriented Programming. Springer, 27–51. https://doi.org/10.1007/978-3-540-70592-5_3
- Ševčík et al. [2013] J. Ševčík, V. Vafeiadis, F. Zappa Nardelli, S. Jagannathan, and P. Sewell. 2013. Compcerttso: A verified compiler for relaxed-memory concurrency. J. ACM 60, 3. https://doi.org/10.1145/2487241.2487248
- Sewell et al. [2010a] P. Sewell, S. Sarkar, S. Owens, F. Zappa Nardelli, and M. Myreen. 2010a. x86-TSO: A rigorous and usable programmer’s model for x86 multiprocessors. Commun. ACM 53, 7, 89–97. https://doi.org/10.1145/1785414.1785443
- Sewell et al. [2010b] P. Sewell, S. Sarkar, S. Owens, F. Zappa Nardelli, and M. O. Myreen. 2010b. x86-TSO: A rigorous and usable programmer’s model for x86 multiprocessors. Commun. ACM 53, 7, 89–97. https://doi.org/10.1145/1785414.1785443
- Shasha and Snir [1988] D. Shasha and M. Snir. 1988. Efficient and correct execution of parallel programs that share memory. ACM Trans. Program. Lang. Syst. 10, 2, 282–312. https://doi.org/10.1145/42190.42277
- Sindhu et al. [1992] P. S. Sindhu, J.-M. Frailong, and M. Cekleov. 1992. Formal specification of memory models. In Scalable Shared Memory Multiprocessors. Springer, 25–41. https://doi.org/10.1007/978-1-4615-3604-8_2
- Singh and Lahav [2023] A. K. Singh and O. Lahav. 2023. An operational approach to library abstraction under relaxed memory concurrency. Proc. ACM Program. Lang. 7, POPL, 1542–1572. https://doi.org/10.1145/3571246
- Singh et al. [2022] S. Singh, D. Sharma, I. Jaju, and S. Sharma. 2022. Fence synthesis under the C11 memory model. In Automated Technology for Verification and Analysis (ATVA). Springer, 83–99. https://doi.org/10.1007/978-3-031-19992-9_6
- Singh et al. [2021] S. Singh, D. Sharma, and S. Sharma. 2021. Dynamic verification of C11 concurrency over multi copy atomics. In Theoretical Aspects of Software Engineering (TASE). IEEE, 39–46. https://doi.org/10.1109/TASE52547.2021.00023
- Smith et al. [2019] G. Smith, N. Coughlin, and T. Murray. 2019. Value-dependent information-flow security on weak memory models. In Formal Methods (FM). Springer, 539–555.
- Smith and Groves [2019] G. Smith and L. Groves. 2019. Weakening correctness and linearizability for concurrent objects on multicore processors. In Formal Methods (FM). Springer, 342–357. https://doi.org/10.1007/978-3-030-54997-8_22
- Smith et al. [2020] G. Smith, K. Winter, and R. Colvin. 2020. Linearizability on hardware weak memory models. Form. Asp. Comput. 32, 1–32. https://doi.org/10.1007/s00165-019-00499-8
- Sorensen and Donaldson [2016] T. Sorensen and A. F. Donaldson. 2016. Exposing errors related to weak memory in GPU applications. In Programming Language Design and Implementation (PLDI). ACM, 100–113. https://doi.org/10.1145/2908080.2908114
- SPARC International Inc. [1994] SPARC International Inc. 1994. The SPARC Architecture Manual (Version 9).
- Steinke and Nutt [2004] R. Steinke and G. Nutt. 2004. A unified theory of shared memory consistency. J. ACM 51, 5, 800–849. https://doi.org/10.1145/1017460.1017464
- Štill et al. [2016] V. Štill, P. Ročkai, and J. Barnat. 2016. Weak memory models as LLVM-to-LLVM transformations. In Mathematical and Engineering Methods in Computer Science (MEMICS). Springer, 144–155. https://doi.org/10.1007/978-3-319-29817-7_13
- Suzanne and Miné [2016] T. Suzanne and A. Miné. 2016. From array domains to abstract interpretation under store-buffer-based memory models. In Static Analysis (SAS). Springer, 469–488. https://doi.org/10.1007/978-3-662-53413-7_23
- Svendsen et al. [2018] K. Svendsen, J. Pichon-Pharabod, M. Doko, O. Lahav, and V. Vafeiadis. 2018. A separation logic for a promising semantics. In Programming Languages and Systems (ESOP). Springer, 357–384. https://doi.org/10.1007/978-3-319-89884-1_13
- Taheri et al. [2019] M. Taheri, A. Pourdamghani, and M. Lesani. 2019. Polynomial-time fence insertion for structured programs. In Proc. 33rd International Symposium on Distributed Computing (DISC). Leibniz-Zentrum für Informatik, 34:1–34:17. https://doi.org/10.4230/LIPIcs.DISC.2019.34
- Tao et al. [2021] R. Tao, J. Yao, X. Li, S.-W. Li, J. Nieh, and R. Gu. 2021. Formal verification of a multiprocessor hypervisor on arm relaxed memory hardware. In Symposium on Operating Systems Principles (SOSP). ACM, 866–881. https://doi.org/10.1145/3477132.3483560
- Thornton [1964] J. E. Thornton. 1964. Parallel operation in the Control Data 6600. In Proceedings of the October 27-29, 1964, Fall Joint Computer Conference, Part II: Very High Speed Computer Systems (AFIPS ’64). 33–40. https://doi.org/10.1145/1464039.1464045
- Tomasulo [1967] R. M. Tomasulo. 1967. An efficient algorithm for exploiting multiple arithmetic units. IBM Journal of Research and Development 11, 1, 25–33. https://doi.org/10.1147/rd.111.0025
- Travkin et al. [2013] O. Travkin, A. Mütze, and H. Wehrheim. 2013. SPIN as a linearizability checker under weak memory models. In Hardware and Software: Verification and Testing (HVC). 311–326. https://doi.org/10.1007/978-3-319-03077-7_21
- Travkin and Wehrheim [2014] O. Travkin and H. Wehrheim. 2014. Handling TSO in mechanized linearizability proofs. In Hardware and Software: Verification and Testing (HVC). 132–147. https://doi.org/10.1007/978-3-319-13338-6_11
- Travkin and Wehrheim [2016] O. Travkin and H. Wehrheim. 2016. Verification of concurrent programs on weak memory models. In Theoretical Aspects of Computing (ICTAC). Springer, 3–24. https://doi.org/10.1007/978-3-319-46750-4_1
- Trippel et al. [2019] C. Trippel, D. Lustig, and M. Martonosi. 2019. Security verification via automatic hardware-aware exploit synthesis: The CheckMate approach. IEEE Micro 39, 3, 84–93. https://doi.org/10.1109/MM.2019.2910010
- Trippel et al. [2018] C. Trippel, Y. A. Manerkar, D. Lustig, M. Pellauer, and M. Martonosi. 2018. Full-stack memory model verification with TriCheck. IEEE Micro 38, 3, 58–68. https://doi.org/10.1109/MM.2018.032271062
- Tunç et al. [2023] H. C. Tunç, P. A. Abdulla, S. Chakraborty, S. Krishna, U. Mathur, and A. Pavlogiannis. 2023. Optimal reads-from consistency checking for C11-style memory models. Proc. ACM Program. Lang. 7, PLDI. https://doi.org/10.1145/3591251
- Turon et al. [2014] A. Turon, V. Vafeiadis, and D. Dreyer. 2014. GPS: Navigating weak memory with ghosts, protocols, and separation. In OOPSLA. ACM, 691–707. https://doi.org/10.1145/2660193.2660243
- Vafeiadi Bila et al. [2022] E. Vafeiadi Bila, B. Dongol, O. Lahav, A. Raad, and J. Wickerson. 2022. View-based Owicki-Gries reasoning for persistent x86-TSO. In Programming Languages and Systems (ESOP). Springer, 234–261. https://doi.org/10.1007/978-3-030-99336-8_9
- Vafeiadis et al. [2015] V. Vafeiadis, T. Balabonski, S. Chakraborty, R. Morisset, and F. Zappa Nardelli. 2015. Common compiler optimisations are invalid in the C11 memory model and what we can do about it. In Principles of Programming Languages (POPL). ACM, 209–220. https://doi.org/10.1145/2676726.2676995
- Vafeiadis and Narayan [2013] V. Vafeiadis and C. Narayan. 2013. Relaxed separation logic: a program logic for C11 concurrency. In OOPSLA. ACM, 867–884. https://doi.org/10.1145/2509136.2509532
- Watt et al. [2020] C. Watt, C. Pulte, A. Podkopaev, G. Barbier, S. Dolan, S. Flur, J. Pichon-Pharabod, and S. Guo. 2020. Repairing and mechanising the JavaScript relaxed memory model. In Programming Language Design and Implementation (PLDI). ACM, 346–361. https://doi.org/10.1145/3385412.3385973
- Wehrheim et al. [2023] H. Wehrheim, L. Bargmann, and B. Dongol. 2023. Reasoning about promises in weak memory models with event structures. In Formal Methods (FM). Springer, 282–300. https://doi.org/10.1007/978-3-031-27481-7_17
- Wehrheim and Travkin [2015] H. Wehrheim and O. Travkin. 2015. TSO to SC via symbolic execution. In Hardware and Software: Verification and Testing (HVC). 104–119. https://doi.org/10.1007/978-3-319-26287-1_7
- Wickerson et al. [2017] J. Wickerson, M. Batty, T. Sorensen, and G. A. Constantinides. 2017. Automatically comparing memory consistency models. In Principles of Programming Languages (POPL). ACM, 190–204. https://doi.org/10.1145/3009837.3009838
- Winskel [1987] G. Winskel. 1987. Event structures. In Petri Nets: Applications and Relationships to Other Models of Concurrency. 325–392. https://doi.org/10.1007/3-540-17906-2_31
- Wright et al. [2021] D. Wright, M. Batty, and B. Dongol. 2021. Owicki-Gries reasoning for C11 programs with relaxed dependencies. In Formal Methods (FM). Springer, 237–254. https://doi.org/10.1007/978-3-030-90870-6_13
- Yang et al. [2001] Y. Yang, G. Gopalakrishnan, and G. Lindstrom. 2001. Analyzing the CRF Java memory model. In Proc. 8th Asia-Pacific Software Engineering Conference (APSEC). IEEE, 21–28. https://doi.org/10.1109/APSEC.2001.991455
- Yang et al. [2002] Y. Yang, G. Gopalakrishnan, and G. Lindstrom. 2002. Specifying Java thread semantics using a uniform memory model. In Proc. 2002 Joint ACM-ISCOPE Conference on Java Grande (JGI). ACM, 192–201. https://doi.org/10.1145/583810.583832
- Yang et al. [2005] Y. Yang, G. Gopalakrishnan, and G. Lindstrom. 2005. UMM: an operational memory model specification framework with integrated model checking capability. Concurrency Comput. Pract. Exper. 17, 465–487. Issue 5‐6. https://doi.org/10.1002/cpe.837
- Zaliva et al. [2024] V. Zaliva, K. Memarian, R. Almeida, J. Clarke, B. Davis, A. Richardson, D. Chisnall, B. Campbell, I. Stark, R. N. M. Watson, and P. Sewell. 2024. Formal mechanised semantics of CHERI C: Capabilities, undefined behaviour, and provenance. In Architectural Support for Programming Languages and Operating Systems (ASPLOS). ACM, 181–196. https://doi.org/10.1145/3617232.3624859

### v2

@@
 If those constraints are too weak to enforce the logic of an algorithm, programmers may inject “fences” or related mechanisms that can restore the standard, sequential behaviour of a (segment of a) program.
-For instance, the use of release/acquire constraints is a well-known programming pattern for addressing weak memory model effects [127].
+For instance, the use of release/acquire constraints is a well-known programming pattern for addressing weak memory model effects [Gharachorloo et al., 1990].
 A release fence, ‘’, (in C, atomic_thread_fence(memory_order_release)) can be injected into the first process of (1), .
 In tandem with an acquire fence after the loop of the second process these ensure the intended value for will be read.
-However, excessive use of such mechanisms comes with performance penalties, which can be a significant issue for low-level systems code [250].
+However, excessive use of such mechanisms comes with performance penalties, which can be a significant issue for low-level systems code [Ritson and Owens, 2016].
 Since the early days of multiprocessors (a system consisting of multiple processors) to the present age of multicore processors (each processor consisting of multiple cores), the literature on weak memory models has developed in academia, the commercial sector, and standardisation committees.
-The issue affects hardware designers, compiler writers, application programmers, and language designers [17; 16].
+The issue affects hardware designers, compiler writers, application programmers, and language designers [Adve and Gharachorloo, 1996; Adve and Boehm, 2010].
 As yet, no consensus on a particular formal representation has been settled upon, leaving programmers – in particular those requiring formal assurance for the code – with many choices.
@@
 We aim to trace the historical development of processor design from the 1960s, through theoretical developments in the 1990s and up to the present day, which includes regular contributions in major publishing venues and refinements of commercial specifications.
-As a result, we were unable to find a methodical sequence of search terms of the type used in a recent survey on programming language memory models [219].
+As a result, we were unable to find a methodical sequence of search terms of the type used in a recent survey on programming language memory models [Moiseenko et al., 2021].
 Hence, we do not claim full coverage, but intend that the range of works surveyed is indicative of the state of the field.
 We do not specifically consider issues related to hardware or compiler construction: we assume that microarchitectural implementations of instruction set architectures faithfully match their specifications, and that compiler transformations correctly preserve their intended software memory model in the underlying architecture.
-Of course, this is not always the case, with problems being discovered in hardware [36; 38] and potential compiler mappings [284].
+Of course, this is not always the case, with problems being discovered in hardware [Alglave et al., 2014b; Arm Ltd., 2011] and potential compiler mappings [Trippel et al., 2018].
 Outline.
@@
 We then briefly review aspects of processors that give rise to weak memory effects (Section 2), and give an overview of prominent weak memory models, past and present (Section 3).
-As an introduction to formalisations and the range of representations that have been considered we present an operational (Section 4) representation based on write buffers (as defined in [261]), and and an axiomatic (Section 5) view that abstracts from architectural details and instead focusses on properties of traces of behaviours (based on work best exemplified in [36]).
+As an introduction to formalisations and the range of representations that have been considered we present an operational (Section 4) representation based on write buffers (as defined in [Sewell et al., 2010b]), and and an axiomatic (Section 5) view that abstracts from architectural details and instead focusses on properties of traces of behaviours (based on work best exemplified in [Alglave et al., 2014b]).
 We use a representative fragment of Intel’s x86 as a running example of a memory model, and use the store buffering litmus test (Example 1.2) as the running example program.
@@
 The word “model” is heavily overloaded and potentially ambiguous.111For example, a “memory model” can mean the structure of the memory address space.
-This usage, unrelated to concurrency, occurs in microarchitectures [133, p.72] as well as in C programming [189; 213].
-In the setting of C, a memory model in this sense is sometimes called a “memory object model.” In this paper, a memory model222A memory model is also called memory consistency, consistency model and memory consistency model [220].
+This usage, unrelated to concurrency, occurs in microarchitectures [González et al., 2011, p.72] as well as in C programming [Lepigre et al., 2022; Memarian et al., 2016].
+In the setting of C, a memory model in this sense is sometimes called a “memory object model.” In this paper, a memory model222A memory model is also called memory consistency, consistency model and memory consistency model [Nagarajan et al., 2020].
 denotes the underlying set of possible behaviours of every concurrent program.
@@
 Given some initial state of the shared and local variables, one can “probe” a hardware or software memory model by checking whether a final state is reachable for a given litmus test.
-Examples 1.2–1.2 give several litmus tests, including the store buffering litmus test, which we use as a running example, and is one of the most common litmus tests referred to in the literature [209; 202].
+Examples 1.2–1.2 give several litmus tests, including the store buffering litmus test, which we use as a running example, and is one of the most common litmus tests referred to in the literature [Maranget et al., 2012; Mador-Haim et al., 2011].
 Table 1 summarises the relationships between the key litmus tests and the common memory models.
@@
 | LB+ctrls | (Ex.3.3, p.3.3) |
-[t]
 2. The Development of Processors
@@
 Milestones in processor design relevant to the development of weak memory models are outlined in Table 2.
-| 1963 | The CDC 6600 [278] significantly improved processor speed by using a “scoreboard” to track which registers and locations are being accessed by instructions in the pipeline; instructions that have no dependencies may be parallelised.
+| 1963 | The CDC 6600 [Thornton, 1964] significantly improved processor speed by using a “scoreboard” to track which registers and locations are being accessed by instructions in the pipeline; instructions that have no dependencies may be parallelised.
 |
-| 1967 | The first implementation of Tomasulo’s algorithm [279] was Model 91 in IBM’s System/360 family [37], which improved the utilisation of multiple arithmetic units. |
-| 1980s | The Stanford DASH multiprocessor [188] brought together much existing research and influenced later processors. |
+| 1967 | The first implementation of Tomasulo’s algorithm [Tomasulo, 1967] was Model 91 in IBM’s System/360 family [Anderson et al., 1967], which improved the utilisation of multiple arithmetic units.
+|
+| 1980s | The Stanford DASH multiprocessor [Lenoski et al., 1992] brought together much existing research and influenced later processors. |
 | 1990s | Speculative execution was introduced by Intel’s Pentium family of processors (a type of x86 architecture). |
@@
 |
-| 2011 | RISC-V was released as an openly available specification [249], with a memory model based on Release Consistency (see Section 3.1). |
-| 2017 | Based on research outcomes, as described in [36], the Arm specification explicitly became multicopy-atomic [240]. |
+| 2011 | RISC-V was released as an openly available specification [RISC-V Foundation, 2019], with a memory model based on Release Consistency (see Section 3.1). |
+| 2017 | Based on research outcomes, as described in [Alglave et al., 2014b], the Arm specification explicitly became multicopy-atomic [Pulte et al., 2017]. |
 2.1. Instruction-Level Parallelism
@@
 These latter micro-steps, which we call “memory operations”, can include complex interactions with the cache system, or across PCI devices.
-Any one of these or many other aspects of computer hardware [147; 230; 212] could potentially influence the actual order of execution of the original sequence of compiled instructions.
+Any one of these or many other aspects of computer hardware [Hennessy and Patterson, 2011; Patterson and Hennessy, 2016; McKenney, 2010] could potentially influence the actual order of execution of the original sequence of compiled instructions.
 Focussing at the local level, the performance of a processor can be improved in part by executing independent micro-steps in parallel.
-Techniques in this vein are collectively termed instruction-level parallelism, and have been developed since the early days of computing [246].
+Techniques in this vein are collectively termed instruction-level parallelism, and have been developed since the early days of computing [Rau and Fisher, 1993].
 Straightforward cases include parallelising loads of distinct locations in memory, and parallelising independent calculations to utilise multiple arithmetic and logic units (ALUs) on a single chip.
 The decision whether or not to parallelise is guided by the observable effect on the sequential semantics – if the programmer could potentially observe the parallelism on a single chip, then it should not be allowed (see Example 2.1).
-Early examples include the “scoreboarding” technique of the CDC 6600 computer [278] and Tomasulo’s algorithm [279].
+Early examples include the “scoreboarding” technique of the CDC 6600 computer [Thornton, 1964] and Tomasulo’s algorithm [Tomasulo, 1967].
 Parallelism can also be employed when accessing distinct locations in memory (see Example 2.1).
-[t]
 As an example of how microarchitectural features can result in weak memory effects we outline below a standard view of processor pipelines.
@@
 As different instructions may take different lengths of time in the issue and complete phases, they may be parallelised.
-[t]
 These “reordering” principles were extended by Intel’s Pentium processors in the 1990s to include speculative execution: while waiting for a branch condition to be resolved to true or false (e.g., while waiting for a value to return from memory), a processor may predict the result and begin transiently executing down the corresponding path.
@@
 Where processor speeds are much faster than memory latency, speculative execution provides a substantial performance gain.
-Such optimisations have led to significant security vulnerabilities, exposed by attacks such as Spectre [171] and Meltdown [193], which were only discovered decades after speculative execution and other optimisation mechanisms had become standard.
+Such optimisations have led to significant security vulnerabilities, exposed by attacks such as Spectre [Kocher et al., 2019] and Meltdown [Lipp et al., 2018], which were only discovered decades after speculative execution and other optimisation mechanisms had become standard.
 The out-of-order executions outlined above pose no issue on single-core processors, as they have been designed to preserve the sequential semantics of programs.
@@
 Some processors come with more fine-grained fences; such a specialised fence may only affect specific types of memory operations (loads or stores), or may only affect instructions before or after it.
-Memory ordering constraints may also be used to enforce ordering between otherwise independent instructions; the classic examples are the release and acquire constraints pioneered by Release Consistency [127] (see Example 2.2 and Section 3.1: as an example, Arm provides both LDAR and LDAPR instructions for acquire loads).
-[t]
+Memory ordering constraints may also be used to enforce ordering between otherwise independent instructions; the classic examples are the release and acquire constraints pioneered by Release Consistency [Gharachorloo et al., 1990] (see Example 2.2 and Section 3.1: as an example, Arm provides both LDAR and LDAPR instructions for acquire loads).
 2.3. Forwarding
@@
 In contrast, on a system that lacks multicopy atomicity, Processes B and C may disagree on the order of modifications, i.e., Process B may observe the modification to followed by the modification to , but Process C can simultaneously observe the modification to before the modification to (see Example 2.4).
-[ht]
-Currently, IBM’s Power series is the only commercially available CPU architecture that lacks multicopy atomicity (the NVIDIA PTX GPU also lacks multicopy atomicity [224]).
+Currently, IBM’s Power series is the only commercially available CPU architecture that lacks multicopy atomicity (the NVIDIA PTX GPU also lacks multicopy atomicity [NVIDIA Ltd, 2024]).
 This is due to its cache coherence system, which effectively allows individual cores to communicate point-to-point via shared caches, before writes are flushed to the shared memory.
 The lack of multicopy atomicity creates an extra level of complexity for describing memory models, as well as understanding their effects on code – both at the assembler-level as well as in high-level languages to be compiled to the hardware.
-To this end, Arm officially stipulates multicopy atomicity for their architectures from Version 8 in 2017 [240].
+To this end, Arm officially stipulates multicopy atomicity for their architectures from Version 8 in 2017 [Pulte et al., 2017].
 (In fact, Arm processors effectively have always been multicopy-atomic.
-Although multicopy atomicity was not explicitly stipulated by earlier versions of Arm, no manufacturer of Arm processors ever produced an implementation that lacked multicopy atomicity [240].) Similarly, RISC-V insists on multicopy atomicity, and Intel’s x86 architectures are also multicopy-atomic.
+Although multicopy atomicity was not explicitly stipulated by earlier versions of Arm, no manufacturer of Arm processors ever produced an implementation that lacked multicopy atomicity [Pulte et al., 2017].) Similarly, RISC-V insists on multicopy atomicity, and Intel’s x86 architectures are also multicopy-atomic.
 In practice, the complexity of dealing with an architecture that lacks multicopy atomicity leads programmers to insert enough fences to restore multicopy atomicity (or indeed Sequential Consistency).
-We therefore do not consider the lack of multicopy atomicity in this paper; a description of the behaviour of Power is given in [122], and the difficulties of verification under Power’s memory model are discussed in [94].
+We therefore do not consider the lack of multicopy atomicity in this paper; a description of the behaviour of Power is given in [Flur et al., 2016], and the difficulties of verification under Power’s memory model are discussed in [Coughlin et al., 2022].
 3. The Development of Memory Models
@@
 | SPARC | Power | x86 | Arm | RISC-V | Java | C | |
-| 1990s | [263] | [92] | |||||
-| 2000s | [227] | [208] | [55] | ||||
-| 2010s | [122] | [249] | [50] | ||||
-| 2020s | [27] |
+| 1990s | [Sindhu et al., 1992] | [Corella et al., 1993] | |||||
+| 2000s | [Owens et al., 2009] | [Manson et al., 2005] | [Boehm, 2007] | ||||
+| 2010s | [Flur et al., 2016] | [RISC-V Foundation, 2019] | [Batty et al., 2011] | ||||
+| 2020s | [Alglave et al., 2021] |
 3.1. Early Memory Models
 Lamport [1979] made the first exploration on the notion of the “intuitive correctness” of concurrent programs in the 1970s, and formalised the concept of what later came to be known as Sequential Consistency.
-From the late 1980s onwards there were considerable efforts in defining the semantics of various instruction-level parallelism techniques, such as pipelining and out-of-order execution (see Section 2) [65; 145; 160; 262].
+From the late 1980s onwards there were considerable efforts in defining the semantics of various instruction-level parallelism techniques, such as pipelining and out-of-order execution (see Section 2) [Burch and Dill, 1994; Harman, 2007; Jones et al., 2002; Shasha and Snir, 1988].
 This saw the proposal of many theoretical memory models, aiming to categorise and unify observed and potential behaviours.
-Examples include Weak Consistency [116], PRAM Consistency [194], Release Consistency [127], Cache Consistency [134], Processor Consistency [134], and Causal Consistency [21].
+Examples include Weak Consistency [Dubois et al., 1986], PRAM Consistency [Lipton and Sandberg, 1988], Release Consistency [Gharachorloo et al., 1990], Cache Consistency [Goodman, 1991], Processor Consistency [Goodman, 1991], and Causal Consistency [Ahamad et al., 1995].
 These memory models were unified in 2004 by Steinke and Nutt [2004], who used a mathematical formalism to compare these memory models and to arrange them in a lattice.
-(In 2018, these theoretical models were again unified and compared [256; 257].)
+(In 2018, these theoretical models were again unified and compared [Senftleben and Schneider, 2018a, b].)
 Before the advent of multicore processors, computers that had multiprocessing were mostly large-scale machines at scientific institutions, and only the programmers at these institutions needed to handle weak memory behaviours.
@@
 SPARC
-The Total Store Order (TSO) and Partial Store Order (PSO) memory models were defined for the SPARC architecture of Sun Microsystems [263].
+The Total Store Order (TSO) and Partial Store Order (PSO) memory models were defined for the SPARC architecture of Sun Microsystems [Sindhu et al., 1992].
 As the name suggests, TSO keeps the writes of each thread in program order, but a read that accesses some location can complete before an earlier write of the same thread to a location different from .
 PSO relaxes this further to maintain order only on writes to the same variable.
-SPARC also defines a Relaxed Memory Order (RMO) model, which “places no ordering constraints on memory references beyond those required for processor self-consistency” [271].
-The TSO model of SPARC is formalised in Isabelle/HOL [150].
+SPARC also defines a Relaxed Memory Order (RMO) model, which “places no ordering constraints on memory references beyond those required for processor self-consistency” [SPARC International Inc., 1994].
+The TSO model of SPARC is formalised in Isabelle/HOL [Hóu et al., 2021].
 x86
@@
 To this date, x86-TSO is widely accepted as the correct description of the behaviour of x86 architectures.
-In [227], x86-TSO was introduced in terms of two formulations, one operational and one axiomatic.
+In [Owens et al., 2009], x86-TSO was introduced in terms of two formulations, one operational and one axiomatic.
 The operational formulation uses a write buffer to model the weak memory effects, and is influential on many later formalisations of memory models; we provide a version of it in Section 4.
-The two key papers [227; 255] that developed x86-TSO are summarised by Sewell et al.
-[260].
+The two key papers [Owens et al., 2009; Sarkar et al., 2009] that developed x86-TSO are summarised by Sewell et al.
+[2010a].
 Power
@@
 Alglave et al.
-[33, 35, 36] also studied the Power memory model based on an axiomatic definition.
+[2010, 2012, 2014b] also studied the Power memory model based on an axiomatic definition.
 An alternative, operational approach to formalisation of Power was given by Sarkar et al.
@@
 Whereas the desktop and laptop markets are dominated by x86, as of 2025 the mobile device market is dominated by Arm architectures.
-Arm architectures previously did not forbid non-multicopy-atomic behaviours, but after a major revision in 2017, Arm has mandated multicopy atomicity from Version 8 onwards [240; 27].
-There have been numerous works on specifying versions of the Arm memory model using different formalisms, such as [122].
-The development history is included in the latest official paper on the Arm memory model [27].
+Arm architectures previously did not forbid non-multicopy-atomic behaviours, but after a major revision in 2017, Arm has mandated multicopy atomicity from Version 8 onwards [Pulte et al., 2017; Alglave et al., 2021].
+There have been numerous works on specifying versions of the Arm memory model using different formalisms, such as [Flur et al., 2016].
+The development history is included in the latest official paper on the Arm memory model [Alglave et al., 2021].
 RISC-V
-The RISC-V architecture [249], first released in 2010, is an open-source architecture that attempts to remove some of the problems associated with commercial-in-confidence manufacturing.
+The RISC-V architecture [RISC-V Foundation, 2019], first released in 2010, is an open-source architecture that attempts to remove some of the problems associated with commercial-in-confidence manufacturing.
 Its memory model is based on experiences at Arm, and includes many recent developments in weak memory theory, including support for release-acquire primitives.
 Specialised Hardware
-Recent research effort has focussed on GPUs [24; 270; 198] and FPGAs [151], and their combination with CPUs [130].
-GPUs in particular focus on significant parallelisation of processing power, and hence would benefit from formalisation; NVIDIA has provided a rigorous description of its memory model [224], which lacks multicopy atomicity (see Section 2.4).
-Earlier formalisations determined a type of total-store-order model [24], while more recent formalisations suggest a weaker model again [198].
+Recent research effort has focussed on GPUs [Alglave et al., 2015; Sorensen and Donaldson, 2016; Lustig et al., 2019] and FPGAs [Iorga et al., 2021], and their combination with CPUs [Goens et al., 2023].
+GPUs in particular focus on significant parallelisation of processing power, and hence would benefit from formalisation; NVIDIA has provided a rigorous description of its memory model [NVIDIA Ltd, 2024], which lacks multicopy atomicity (see Section 2.4).
+Earlier formalisations determined a type of total-store-order model [Alglave et al., 2015], while more recent formalisations suggest a weaker model again [Lustig et al., 2019].
 Earlier Hardware Memory Models
-Intel’s Itanium processors [14] never became as commercially popular as other processors, and have since ceased production.
+Intel’s Itanium processors [Abe and Maeda, 2017] never became as commercially popular as other processors, and have since ceased production.
 Similarly, DEC’s Alpha processors were prominent for a time, but have also ceased production.
@@
 Java.
-Sun Microsystems released the Java programming language in 1996, with multi-threading and shared-memory behaviour informally described in [137].
-This was immediately followed by multiple independent efforts to formalise the memory model of Java [43; 132; 148].
-There is a formulation that explicitly used local working memories [58], and a more abstract formulation with event structures [76].
-In 1999, however, the original prose memory model turned out to contain flaws [238; 239].
-Subsequent work [203; 206; 252; 296; 297; 44; 207] attempted to revise the Java memory model, culminating in [208], which describes the new Java memory model as part of Java 5.0.
-After the formulation in [208] became accepted as the official Java model, a few alternative specifications continued to emerge [298; 154; 195; 53].
+Sun Microsystems released the Java programming language in 1996, with multi-threading and shared-memory behaviour informally described in [Gosling et al., 1996].
+This was immediately followed by multiple independent efforts to formalise the memory model of Java [Attali et al., 1998; Gontmakher and Schuster, 1998; Higham and Kawash, 1998].
+There is a formulation that explicitly used local working memories [Börger and Schulte, 1999], and a more abstract formulation with event structures [Cenciarelli et al., 1999].
+In 1999, however, the original prose memory model turned out to contain flaws [Pugh, 1999, 2000].
+Subsequent work [Maessen et al., 2000; Manson and Pugh, 2001; Roychoudhury and Mitra, 2002; Yang et al., 2001, 2002; Awhad and Wallace, 2003; Manson and Pugh, 2004] attempted to revise the Java memory model, culminating in [Manson et al., 2005], which describes the new Java memory model as part of Java 5.0.
+After the formulation in [Manson et al., 2005] became accepted as the official Java model, a few alternative specifications continued to emerge [Yang et al., 2005; Jagadeesan et al., 2010; Lochbihler, 2013; Bender and Palsberg, 2019].
 On the compiler-related side, Ševčík and Aspinall [2008] identified common optimisations performed Java compilers that do not comply with the Java memory model.
@@
 However, as Boehm [2005] pointed out in 2005, “threads cannot be implemented as a library”: the correctness of concurrent code cannot be guaranteed when the concurrency mechanisms are designed separately from the core language and its compiler.
-Boehm and others worked as part of the International Organisation for Standardisation (ISO) on the formal semantics of concurrent C and C++ programs [55; 56].
+Boehm and others worked as part of the International Organisation for Standardisation (ISO) on the formal semantics of concurrent C and C++ programs [Boehm, 2007; Boehm and Adve, 2008].
 The resulting 2011 ISO standards of the C and C++ programming languages were the first time that multi-threading in both languages was formally defined, and the memory model defined by these ISO standards is commonly referred to as the C memory model (or C11 in the literature, owing to its inception in 2011).
@@
 [2016].
-The prominence of the C language and its memory model has drawn much attention in the context of verification, and many recent publications are devoted to program logics [295], model checkers [173], or other problems [266; 265; 86]; more of these are covered in Section 8.
+The prominence of the C language and its memory model has drawn much attention in the context of verification, and many recent publications are devoted to program logics [Wright et al., 2021], model checkers [Kokologiannakis et al., 2018], or other problems [Singh et al., 2021, 2022; Colvin, 2022]; more of these are covered in Section 8.
 Software memory models are often subject to more potential issues, due to the wider range of transformations open to a compiler, on top of the weak memory behaviours on the target hardware.
 These issues include the “out of thin air” (see Example 3.3) and “read from untaken branch” behaviours.
-These have been particularly addressed in the context of C, and the question about how to forbid out-of-thin air (and similar) behaviours without overly constraining standard compiler optimisations is a subject of ongoing debate in the standardisation committees and the academic community [57; 48; 155; 232; 226; 79].
+These have been particularly addressed in the context of C, and the question about how to forbid out-of-thin air (and similar) behaviours without overly constraining standard compiler optimisations is a subject of ongoing debate in the standardisation committees and the academic community [Boehm and Demsky, 2014; Batty et al., 2015; Jeffrey and Riely, 2016; Pichon-Pharabod and Sewell, 2016; Ou and Demsky, 2018; Chakraborty and Vafeiadis, 2019].
 The most recent C reference explicitly gives a version of out-of-thin-air behaviour, stating that it is a theoretically allowed behaviour, but with a caveat that no compiler should ever implement it.
-[ht]
-With the memory model of Power lacking multicopy atomicity and hence being harder to handle (see Section 3.2), there have been works that focus on the compilation from C to Power while respecting the memory models of both ends [49; 253; 183], and more general compiler optimisations under the C memory model [233; 109; 126].
+With the memory model of Power lacking multicopy atomicity and hence being harder to handle (see Section 3.2), there have been works that focus on the compilation from C to Power while respecting the memory models of both ends [Batty et al., 2012; Sarkar et al., 2012; Lahav et al., 2017], and more general compiler optimisations under the C memory model [Podkopaev et al., 2019; Dodds et al., 2018; Geeson and Smith, 2024].
 The Linux Kernel.
 Alglave et al.
-[32] provide a formalisation of the memory model of the Linux kernel.
+[Alglave et al., 2018] provide a formalisation of the memory model of the Linux kernel.
 This model is to further codify the complex interaction of the wide variety of hardware memory models and the C language memory model, supporting some of the common patterns specific to Linux kernel programming.
@@
 Every programming language that supports shared-memory concurrency has interplay with weak memory models.
-For programming languages such as JavaScript [135; 290], OCaml [114], OpenCL [167] and Rust [102], there have been works that study their own memory models, or how they are affected by the weak memory models of the languages to which they may be compiled.
-A potential simplification for this effort for many languages is via a memory model for an intermediate language such as LLVM [78; 186], and through the investigation of general properties related to safe optimisation strategies [136].
+For programming languages such as JavaScript [Gopalakrishnan and Verbrugge, 2020; Watt et al., 2020], OCaml [Dolan et al., 2018], OpenCL [Khronos OpenCL Working Group, 2020] and Rust [Dang et al., 2019], there have been works that study their own memory models, or how they are affected by the weak memory models of the languages to which they may be compiled.
+A potential simplification for this effort for many languages is via a memory model for an intermediate language such as LLVM [Chakraborty and Vafeiadis, 2017; Lee et al., 2023], and through the investigation of general properties related to safe optimisation strategies [Gopalakrishnan et al., 2023].
 4. Introduction to Operational Formalisms
 Before giving a general survey of the field we provide an example formalisation: a simplification of the formalisation of x86 given by the influential work by Owens et al.
-[227; 260].
-This formalisation is operational, meaning described by way of individual steps of the system, and as such is often easier to follow than other types of semantics [159].
+[Owens et al., 2009; Sewell et al., 2010a].
+This formalisation is operational, meaning described by way of individual steps of the system, and as such is often easier to follow than other types of semantics [Jones and Astarte, 2018].
 We also choose it as it is based around a microarchitectural feature, and as such provides a contrast to axiomatic descriptions, which purposefully abstract away from such considerations.
-In this formalisation, weak memory effects (specifically, loads appearing to take effect before earlier stores) emerge by the use of a write buffer in each core (acting similarly to a pipeline (Section 2.1) or as part of the cache system [212]).
+In this formalisation, weak memory effects (specifically, loads appearing to take effect before earlier stores) emerge by the use of a write buffer in each core (acting similarly to a pipeline (Section 2.1) or as part of the cache system [McKenney, 2010]).
 We first give an operational semantics for a standard sequential language, as if instructions were executed in the order specified by the program text; after extending that semantics with extra rules covering write buffers the weaker behaviours of x86 (Total Store Order) become apparent.
@@
 System level
-TODO
 (seq) states that a sequential composition of commands may always take a step of . If has terminated (i.e., is ), then may be executed via (skip).
@@
 A system is said to terminate if all processes terminate, and a system state, , is said to be reachable if the system from some starting state can terminate in via a sequence of top-level transitions of (mem-read), (mem-write), and (mem-).
-[ht]
 4.3. Write Buffers for x86
-We now augment the system of Sections 4.1 and 4.2 with an explicit write buffer for each core, following [227].
+We now augment the system of Sections 4.1 and 4.2 with an explicit write buffer for each core, following [Owens et al., 2009].
 A write buffer, of type , is a list of pairs , where is a shared variable and is a value being written to , representing a pending write of to .
@@
 All previous transitions in Fig.1 still apply as-is, with the combination of (x86-write-buf) and (x86-read-mem) allowing loads to take effect before stores.
-TODO
-[t]
 In (x86-write-buf), when a core issues a write via rule (store), rather than interacting directly with the shared memory it is appended to the end of the write buffer.
@@
 This is typically referred to as an axiomatic approach, to which we turn our attention in Section 5.
-Operationally one of the best-known approaches is the Promising Semantics [165; 187; 241], as discussed in Section 6.2, which combines an operational semantics with global axioms.
+Operationally one of the best-known approaches is the Promising Semantics [Kang et al., 2017; Lee et al., 2020; Pulte et al., 2019], as discussed in Section 6.2, which combines an operational semantics with global axioms.
 Arm and RISC-V
@@
 These constraints encode those of Definition 2.1, modified by fences as in Example 2.2, and allow the weaker behaviours shown in Example 2.1.
-A more complete operational description of pipelines that conform to the Arm memory model, and includes branches and jumps, is given in [90].
+A more complete operational description of pipelines that conform to the Arm memory model, and includes branches and jumps, is given in [Colvin and Su, 2025].
 Release/Acquire
-The release/acquire concepts [127], as introduced in (1) and Example 2.2, can also be described operationally.
+The release/acquire concepts [Gharachorloo et al., 1990], as introduced in (1) and Example 2.2, can also be described operationally.
 Continuing the generalisation of write buffers to pipelines, an operation with release semantics can be executed only when at the head of the pipeline; i.e., when all other prior operations (w.r.t the program order) have completed.
@@
 This additional constraint means that an earlier acquire load of any variable prevents a later write from proceeding.
-For the stronger form of release/acquire consistency in [127] an acquire load additionally requires the absence of release stores in the earlier pipeline, which can be described syntactically similarly to the boxed constraint of (RA-write-mem).
+For the stronger form of release/acquire consistency in [Gharachorloo et al., 1990] an acquire load additionally requires the absence of release stores in the earlier pipeline, which can be described syntactically similarly to the boxed constraint of (RA-write-mem).
 Atomic operations
 Instruction types such as swap, compare-and-swap, fetch-and-increment, etc., combine read and write operations atomically.
-These can be described using the concept of a global lock which prevents a core from performing a memory operation while another core holds the lock [261].
+These can be described using the concept of a global lock which prevents a core from performing a memory operation while another core holds the lock [Sewell et al., 2010b].
 For instance, the atomic swap instruction, , updates the shared variable to hold the value in register , and updates to the initial value of ; this instruction can be defined as follows.
@@
 | ext | external | Def. 5.1 | |||
-| po | program order | Def. 5.2 | sb | sequenced-before | [50] |
+| po | program order | Def. 5.2 | sb | sequenced-before | [Batty et al., 2011] |
 | poloc | po per location | Def. 5.3 | |||
 | fence | fence relation | Def. 5.4 | |||
-| co | coherence order | Def. 5.5 | mo | modification order | [50] |
-| ws | write serialization | [22] | |||
+| co | coherence order | Def. 5.5 | mo | modification order | [Batty et al., 2011] |
+| ws | write serialization | [Alglave, 2012] | |||
 | rf | reads-from | Def. 5.6 | wt | writes-to | |
@@
 | rfe | external reads-from | Def. 5.13 | wte | external writes-to | |
-TODO
 Definition 5.1 (loc and ext).
@@
 Note that the converse of a valid rf must be functional. Indeed, means that reads from , and every read-event must have a unique “source”. This property underpins Definition 5.8 later.
-On the name “reads-from”, here we follow the recent work by Arm [27], where a typical pair means “ reads from ”, with occurring before .
+On the name “reads-from”, here we follow the recent work by Arm [Alglave et al., 2021], where a typical pair means “ reads from ”, with occurring before .
 Some may find it more intuitive to think of the relation rf as writes-to.
 Indeed, means that writes to the value read by .
-The relations po, co and rf are the “base relations”. The set of memory events and these base relations constitute an “execution graph”, also known as a “candidate execution” [27].
+The relations po, co and rf are the “base relations”. The set of memory events and these base relations constitute an “execution graph”, also known as a “candidate execution” [Alglave et al., 2021].
 Definition 5.7.
@@
 The coherence-after relation .
-The extended coherence order – defined to be the transitive closure of com [183].
+The extended coherence order – defined to be the transitive closure of com [Lahav et al., 2017].
 In addition, software memory models make use of the following two relations to enforce ordering.
@@
 A memory model is said to satisfy the coherence property if all possible executions of any execution graph is coherent.
-The coherence property is also known as the uniproc property [22], SC per location [36], SC per variable [26] or internal visibility [27].
+The coherence property is also known as the uniproc property [Alglave, 2012], SC per location [Alglave et al., 2014b], SC per variable [Alglave and Cousot, 2017] or internal visibility [Alglave et al., 2021].
 The coherence property is satisfied by most memory models.
@@
 Note that the coherence property follows from this definition because . Examples 5.2 and 5.2 now demonstrate how this definition is applied.
-[t]
 5.3. Axiomatic Definition of TSO
@@
 Note that this is in contrast to SC (Definition 5.11) where the coherence axiom is implied.
-[ht]
-[t]
 5.4. Axiomatic Descriptions of More Complex Models
 For illustrative purposes, we have presented how axiomatic formalisms can define the straightforward SC memory model and a simplified version of TSO.
-In practice, axiomatic formalisms have been used to specify a wide range of behaviours related to more complex memory models and instruction types, e.g., [36; 240; 27], as surveyed in Section 6.3.
+In practice, axiomatic formalisms have been used to specify a wide range of behaviours related to more complex memory models and instruction types, e.g., [Alglave et al., 2014b; Pulte et al., 2017; Alglave et al., 2021], as surveyed in Section 6.3.
 Below we give an indication of how these more complex problems are tackled axiomatically.
@@
 Hence, to ensure that every pair in rmw is atomic, we require that no pair in rmw is in , that is, an execution graph enforces the RMW instructions if .
-Further details and mechanised tool support can be found in the online repository of the Herd tool [39; 27].
+Further details and mechanised tool support can be found in the online repository of the Herd tool [Arm Ltd., 2024; Alglave et al., 2021].
 6. Survey of Memory Model Formalisms
@@
 The best-known of such mechanisms is the local write buffers elucidated in Section 4.
-Local write buffers describe the behaviour of TSO [263], on which the x86 memory model is based [227; 255; 260].
+Local write buffers describe the behaviour of TSO [Sindhu et al., 1992], on which the x86 memory model is based [Owens et al., 2009; Sarkar et al., 2009; Sewell et al., 2010a].
 As local write buffers are relatively straightforward to encode, they often serve as the foundation for verification under TSO.
-Namely, when each write buffer is modelled as a separate process, one can analyse the entire system under Sequential Consistency using existing tools [68; 280; 281; 292; 282; 190; 191; 192; 273].
+Namely, when each write buffer is modelled as a separate process, one can analyse the entire system under Sequential Consistency using existing tools [Burckhardt and Musuvathi, 2008; Travkin et al., 2013; Travkin and Wehrheim, 2014; Wehrheim and Travkin, 2015; Travkin and Wehrheim, 2016; Linden and Wolper, 2010, 2011, 2013; Štill et al., 2016].
 The mechanistic formalism of Boudol and Petri [2009] is also based on write buffers that are not local but potentially hierarchical, making this formalism capable of expressing behaviours weaker than those of TSO.
@@
 [2018] to improve various aspects of verification under TSO.
-Meanwhile, two other major hardware platforms – Power [254] and Arm [122; 240] – also have their own mechanistic formulations, albeit much more complex than x86-TSO and harder to encode for verification tools.
-A semantics based directly on pipelined and speculative execution, as outlined in Section 2.1, has been developed for Arm, RISC-V and x86 processors [90].
+Meanwhile, two other major hardware platforms – Power [Sarkar et al., 2011] and Arm [Flur et al., 2016; Pulte et al., 2017] – also have their own mechanistic formulations, albeit much more complex than x86-TSO and harder to encode for verification tools.
+A semantics based directly on pipelined and speculative execution, as outlined in Section 2.1, has been developed for Arm, RISC-V and x86 processors [Colvin and Su, 2025].
 The main advantages of mechanistic formalisms is their intuition and compositionality: as they align closely with the actual mechanisms on processors, they give a local understanding of how weak memory behaviours arise, without needing to understand the system as a whole.
@@
 The timestamp can be some future time, meaning that any subsequent read of with return-value must occur at some later time , i.e., .
-An early use of timestamps is provided in [234]; we look at some other prominent formalisations in this vein below.
+An early use of timestamps is provided in [Podkopaev et al., 2016]; we look at some other prominent formalisations in this vein below.
 The “Promising Semantics”
@@
 [2019].
-The semantics has been used as the basis for verification of non-trivial systems [277].
-A characteristic of the Promising Semantics is that it is a multi-execution formalism [218], as opposed to a single-execution formalism.
+The semantics has been used as the basis for verification of non-trivial systems [Tao et al., 2021].
+A characteristic of the Promising Semantics is that it is a multi-execution formalism [Moiseenko et al., 2022], as opposed to a single-execution formalism.
 In a single-execution formalism (e.g., that of Section 4), one behaviour, e.g., a trace of memory operations, corresponds to one execution of the code.
@@
 Formalisms with timestamped messages and similar notions have since been developed on several fronts, for example, Doherty et al.
-[111], with subsequent works leading to program logics [96; 295; 110; 98; 99; 45; 291].
+[2019], with subsequent works leading to program logics [Dalvandi et al., 2020; Wright et al., 2021; Doherty et al., 2022; Dalvandi and Dongol, 2022; Dalvandi et al., 2022; Bargmann and Wehrheim, 2023; Wehrheim et al., 2023].
 Lahav and Boker [2022] introduced the notion of potentials.
@@
 Timestamped messages, views, and potentials enable general frameworks for handling a wide range of weak memory effects, including the lack of multicopy atomicity (Section 2.4) and out-of-thin air behaviour (Section 3.3).
-This flexibility is not without disadvantages, however: these formalisms are perhaps the most complex of those surveyed, with the details of even small examples such as the store buffering litmus test being difficult to reason about [218].
+This flexibility is not without disadvantages, however: these formalisms are perhaps the most complex of those surveyed, with the details of even small examples such as the store buffering litmus test being difficult to reason about [Moiseenko et al., 2022].
 6.3. Axiomatic Formalisms
 One of the most successful approaches to formalising memory models is via constraints on the allowed behaviours.
-This is referred to as the axiomatic (or declarative) approach in the literature [36; 263].
+This is referred to as the axiomatic (or declarative) approach in the literature [Alglave et al., 2014b; Sindhu et al., 1992].
 Given a set of memory events (where each read has a return value), one starts with all possible permutations, and retains only those permutations that satisfy the axioms of the particular memory model (conceptually the opposite of first considering the set of behaviours allowed under Sequential Consistency and then expanding that set to capture the weak behaviours).
-Due to its preciseness and relative conciseness, the axiomatic approach is widely adopted for the major memory models, e.g., x86-TSO [227], Arm [27], RISC-V [249], Java [208], and C [50].
-Some of the most influential tools in this area are the domain-specific language Cat and its associated litmus-test simulator Herd, introduced in [36].
-Cat and Herd are utilised by Arm to formalise its memory model [27], and Cat is also used in a number of model checkers for expressing other memory models (see Section 8.1).
-An axiomatic formalisation is leveraged by the Check suite of tools and techniques for investigating correctness and security of microarchitectural implementations [197; 204; 283; 205; 199; 284].
+Due to its preciseness and relative conciseness, the axiomatic approach is widely adopted for the major memory models, e.g., x86-TSO [Owens et al., 2009], Arm [Alglave et al., 2021], RISC-V [RISC-V Foundation, 2019], Java [Manson et al., 2005], and C [Batty et al., 2011].
+Some of the most influential tools in this area are the domain-specific language Cat and its associated litmus-test simulator Herd, introduced in [Alglave et al., 2014b].
+Cat and Herd are utilised by Arm to formalise its memory model [Alglave et al., 2021], and Cat is also used in a number of model checkers for expressing other memory models (see Section 8.1).
+An axiomatic formalisation is leveraged by the Check suite of tools and techniques for investigating correctness and security of microarchitectural implementations [Lustig et al., 2014; Manerkar et al., 2018; Trippel et al., 2019; Manerkar et al., 2015; Lustig et al., 2016; Trippel et al., 2018].
 The downside of axiomatic formalisms is that they are based on global reasoning about completed executions, where every has a value, and the relations and are supplied (see Section 5).
 As such, axiomatic formalisms tend not to be suited for reasoning about programs compositionally; that is, one cannot form the set of behaviours of a system by considering individual processes in isolation.
-The purely axiomatic approach has less applicability to theorem proving (however, see [26; 143]), so operational semantics (see Section 6.2) are sometimes developed in tandem, often with an underlying axiomatic flavour [110; 165; 241; 95].
+The purely axiomatic approach has less applicability to theorem proving (however, see [Alglave and Cousot, 2017; Hammond et al., 2024]), so operational semantics (see Section 6.2) are sometimes developed in tandem, often with an underlying axiomatic flavour [Doherty et al., 2022; Kang et al., 2017; Pulte et al., 2019; Crary and Sullivan, 2015].
 6.4. Event Structures and Pomsets
 To express and reason about the subtle behaviours of more complex memory models such as the C memory model, there are works that use event structures and pomsets (partially ordered multisets), which can be viewed as mathematical structures that generalise execution graphs in Section 5.
-Both event structures [294] and pomsets [237] were invented in the 1980s as mathematical models of general concurrency.
+Both event structures [Winskel, 1987] and pomsets [Pratt, 1986] were invented in the 1980s as mathematical models of general concurrency.
 An early use of event structures for weak memory models is by Cenciarelli et al.
 [2007] in 2007, who described a small fragment of the Java memory model by combining event structures with operational and axiomatic techniques.
-Then from 2016 onwards, there have been many attempts to define memory models using related ideas [74; 155; 232; 79; 231; 153; 156; 218].
-The Weakestmo Semantics and its extension [79; 218] are multi-execution, i.e., they encode multiple execution graphs in a single event structure (cf.
+Then from 2016 onwards, there have been many attempts to define memory models using related ideas [Castellan, 2016; Jeffrey and Riely, 2016; Pichon-Pharabod and Sewell, 2016; Chakraborty and Vafeiadis, 2019; Paviotti et al., 2020; Jagadeesan et al., 2020; Jeffrey et al., 2022; Moiseenko et al., 2022].
+The Weakestmo Semantics and its extension [Chakraborty and Vafeiadis, 2019; Moiseenko et al., 2022] are multi-execution, i.e., they encode multiple execution graphs in a single event structure (cf.
 the Promising Semantics Section 6.2).
 6.5. Program Transformations and Instruction Reorderings
-A number of early works [17; 210; 211] specified weak memory models in terms of which pairs of instructions can be reordered.
+A number of early works [Adve and Gharachorloo, 1996; McKenney, 2005a, b] specified weak memory models in terms of which pairs of instructions can be reordered.
 This aligns with the workings of the processor pipeline (Definition 2.1), and is often the level at which processors are described informally.
@@
 Possible instruction reordering is important in the context of compiler-related aspects of weak memory models.
-Common program transformations (including instruction reordering) performed by optimising compilers are not always correct under weak memory models [258; 288].
+Common program transformations (including instruction reordering) performed by optimising compilers are not always correct under weak memory models [Ševčík and Aspinall, 2008; Vafeiadis et al., 2015].
 To formulate what it means for a program transformation to be “correct” under a memory model, one typically develops some trace-based denotational semantics, and a transformation is then correct if the denotation of the transformed program is (a subset of) the denotation of the original program.
@@
 [2024] for the C memory model and its RA fragment.
-Instruction reordering was also the basis in [85], where the memory model is encoded as a parameter to sequential composition on the level of the program text.
-Weak memory effects are explained in terms of the standard operators for (no-parameter) sequential composition and parallel composition, via a set of transformation rules on the instructions, in the spirit of the Concurrent Kleene Algebra [149; 146].
+Instruction reordering was also the basis in [Colvin, 2021], where the memory model is encoded as a parameter to sequential composition on the level of the program text.
+Weak memory effects are explained in terms of the standard operators for (no-parameter) sequential composition and parallel composition, via a set of transformation rules on the instructions, in the spirit of the Concurrent Kleene Algebra [Hoare et al., 2009; Hayes et al., 2016].
 For example, instantiating the SB litmus test on an x86 (TSO) machine, one can derive the following.
@@
 Meanwhile, when fences are inserted as on the second line, instances of the x86-specific sequential composition are equivalent to standard sequential composition.
-This reasoning framework has been used as the basis for addressing security vulnerabilities related to speculative execution [91] and information flow [267].
+This reasoning framework has been used as the basis for addressing security vulnerabilities related to speculative execution [Colvin and Winter, 2020] and information flow [Smith et al., 2019].
 By transforming a program under a weak memory model to a program under Sequential Consistency, standard reasoning methods and tools become amenable, saving the need for specialised logics and techniques.
-However, in contrast to timestamped messages, these program-level transformations cannot explain non-multicopy-atomic behaviours [182].
+However, in contrast to timestamped messages, these program-level transformations cannot explain non-multicopy-atomic behaviours [Lahav and Vafeiadis, 2016].
 7. Program Logics and mechanised support
-There are three prominent families of program logics for concurrent programs that have been adapted to reason about weak memory effects: Owicki-Gries (OG) [228], Rely-Guarantee (RG) [157; 158], and Concurrent Separation Logic (CSL) [152; 247; 64].
+There are three prominent families of program logics for concurrent programs that have been adapted to reason about weak memory effects: Owicki-Gries (OG) [Owicki and Gries, 1976], Rely-Guarantee (RG) [Jones, 1983a, b], and Concurrent Separation Logic (CSL) [Ishtiaq and O’Hearn, 2001; Reynolds, 2002; Brookes and O’Hearn, 2016].
 This section surveys these adaptations; a summary is given in Table 5.
 | Logic | Formalism | Memory models | Case studies | References |
-| OG/RG | Axiomatic | C fragment | RCU | [181] |
-| RG | Operational | x86-TSO | Simpson’s | [248] |
-| RG | Operational | (parameterised) | litmus tests | [13] |
-| OG | View-based | C fragment | Peterson’s, RCU | [96; 99; 110; 295] |
-| RG | potential-based | Causal Consistency | litmus tests, Peterson’s | [179] |
-| CSL | Axiomatic | C fragment | MS queue and others | [112; 113; 289; 286] |
-| CSL | Operational | C fragment | MS queue and others | [164] |
-| CSL | Operational | Rust | Rust library | [102] |
-| CSL | Operational | OCaml | locks, bounded queue | [217; 216] |
-| RG | Reordering | hardware models | Chase-Lev, Peterson’s | [93; 94] |
-| Invariant | Axiomatic | (parameterised) | PostgreSQL | [30; 26] |
-| Invariant | Operational | C fragment | Peterson’s | [111] |
+| OG/RG | Axiomatic | C fragment | RCU | [Lahav and Vafeiadis, 2015] |
+| RG | Operational | x86-TSO | Simpson’s | [Ridge, 2010] |
+| RG | Operational | (parameterised) | litmus tests | [Abe and Maeda, 2016] |
+| OG | View-based | C fragment | Peterson’s, RCU | [Dalvandi et al., 2020, 2022; Doherty et al., 2022; Wright et al., 2021] |
+| RG | potential-based | Causal Consistency | litmus tests, Peterson’s | [Lahav et al., 2023] |
+| CSL | Axiomatic | C fragment | MS queue and others | [Doko and Vafeiadis, 2016, 2017; Vafeiadis and Narayan, 2013; Turon et al., 2014] |
+| CSL | Operational | C fragment | MS queue and others | [Kaiser et al., 2017] |
+| CSL | Operational | Rust | Rust library | [Dang et al., 2019] |
+| CSL | Operational | OCaml | locks, bounded queue | [Mével et al., 2020; Mével and Jourdan, 2021] |
+| RG | Reordering | hardware models | Chase-Lev, Peterson’s | [Coughlin et al., 2021, 2022] |
+| Invariant | Axiomatic | (parameterised) | PostgreSQL | [Alglave et al., 2013b; Alglave and Cousot, 2017] |
+| Invariant | Operational | C fragment | Peterson’s | [Doherty et al., 2019] |
 Axiomatic OG and RG
-Based on an axiomatic formulation of the release-acquire fragment of the C memory model, a rely-guarantee logic was developed and applied to the read-copy-update synchronisation mechanism [181].
+Based on an axiomatic formulation of the release-acquire fragment of the C memory model, a rely-guarantee logic was developed and applied to the read-copy-update synchronisation mechanism [Lahav and Vafeiadis, 2015].
 Operational OG and RG
-A rely-guarantee logic on the x86 assembly instructions was developed in [248], based on the operational formulation of x86-TSO [227].
+A rely-guarantee logic on the x86 assembly instructions was developed in [Ridge, 2010], based on the operational formulation of x86-TSO [Owens et al., 2009].
 This logic was applied to Simpson’s four-slot algorithm, and mechanised in HOL.
-The rely-guarantee logic of [13] works on general memory models that are specified operationally, with case-studies performed on litmus tests.
-Another line of research [96; 99] devised a view-based formalisation (see Section 6.2) of the release-acquire-relaxed fragment of the C memory model, and built an Owicki-Gries logic based on this formulation.
-The main case-studies, carried out in Isabelle/HOL, are Peterson’s algorithm and the read-copy-update mechanism; related works in this line of research include [110; 295].
-Similar to view-based semantics is the “potential-based” semantics, which has been used as the basis of a rely-guarantee logic for Causal Consistency [179].
+The rely-guarantee logic of [Abe and Maeda, 2016] works on general memory models that are specified operationally, with case-studies performed on litmus tests.
+Another line of research [Dalvandi et al., 2020, 2022] devised a view-based formalisation (see Section 6.2) of the release-acquire-relaxed fragment of the C memory model, and built an Owicki-Gries logic based on this formulation.
+The main case-studies, carried out in Isabelle/HOL, are Peterson’s algorithm and the read-copy-update mechanism; related works in this line of research include [Doherty et al., 2022; Wright et al., 2021].
+Similar to view-based semantics is the “potential-based” semantics, which has been used as the basis of a rely-guarantee logic for Causal Consistency [Lahav et al., 2023].
 Axiomatic CSLs
-One line of research consists of the following three generations of logics: RSL (Relaxed Separation Logic) [289], GPS (Ghost states, Protocols, and Separation) [286], and FSL (Fenced Separation Logic) [112; 113].
-All three are based on the axiomatic C memory model [50], and are mechanised in Rocq (formerly Coq).
+One line of research consists of the following three generations of logics: RSL (Relaxed Separation Logic) [Vafeiadis and Narayan, 2013], GPS (Ghost states, Protocols, and Separation) [Turon et al., 2014], and FSL (Fenced Separation Logic) [Doko and Vafeiadis, 2016, 2017].
+All three are based on the axiomatic C memory model [Batty et al., 2011], and are mechanised in Rocq (formerly Coq).
 The case-studies of these logics include simple litmus tests, while GPS in particular was applied to more realistic concurrent data structures such as the Michael-Scott queue, the circular buffer, and the bounded ticket lock.
 Operational CSLs
-The Iris framework [163; 162] uses an underlying operational semantics to unify various concurrent separation logics from the setting of Sequential Consistency.
+The Iris framework [Jung et al., 2015, 2018] uses an underlying operational semantics to unify various concurrent separation logics from the setting of Sequential Consistency.
 Iris is mechanised in Rocq, which is also used by its derivatives listed below.
 Iris was extended to the weak memory setting by combining with GPS.
-The results are iGPS [164] and iRC11 [102]; see also [275].
+The results are iGPS [Kaiser et al., 2017] and iRC11 [Dang et al., 2019]; see also [Svendsen et al., 2018].
 The case-studies of iGPS are the same as those of GPS, while the case-studies of iRC11 are drawn from a library of the Rust programming language.
-Cosmo [217] is another CSL based on Iris for reasoning about multicore OCaml and its operationally defined memory model.
+Cosmo [Mével et al., 2020] is another CSL based on Iris for reasoning about multicore OCaml and its operationally defined memory model.
 The case-studies are a spin lock, a ticket lock, and Peterson’s algorithm.
-The sequel [216] then applied Cosmo to verify a concurrent bounded queue.
-See [121] for an alternative approach to CSL.
+The sequel [Mével and Jourdan, 2021] then applied Cosmo to verify a concurrent bounded queue.
+See [Ferreira et al., 2010] for an alternative approach to CSL.
 RG Based on Reorderings
-There is a rely-guarantee logic [93] for multicopy-atomic hardware memory models, using a reordering formalism [89; 85].
+There is a rely-guarantee logic [Coughlin et al., 2021] for multicopy-atomic hardware memory models, using a reordering formalism [Colvin and Smith, 2018; Colvin, 2021].
 The case-study is the Chase-Lev double-ended queue for Arm, with the proof mechanised in Isabelle/HOL.
-The same authors later extended their framework to handle memory models without multicopy atomicity [94], and verified Peterson’s algorithm in Isabelle/HOL.
-The CLH fine-grained lock, as used in the seL4 microkernel [169], is verified for Arm using a related rely-guarantee-based approach in [88].
+The same authors later extended their framework to handle memory models without multicopy atomicity [Coughlin et al., 2022], and verified Peterson’s algorithm in Isabelle/HOL.
+The CLH fine-grained lock, as used in the seL4 microkernel [Klein et al., 2009], is verified for Arm using a related rely-guarantee-based approach in [Colvin et al., 2024].
 Others
+There are other proof methods for invariant-based reasoning [Alglave et al., 2013b; Alglave and Cousot, 2017; Doherty et al., 2019].
+The case studies in [Alglave et al., 2013b; Alglave and Cousot, 2017] are drawn from PostgreSQL, whereas in [Doherty et al., 2019] the case study is Peterson’s mutual exclusion algorithm.
 8. Model checking and computational problems
@@
 | Name | Type | Memory models |
-| CBMC extended [31] | BMC | x86, Power |
-| Herd [27] | simulator | parametric (Cat language) |
-| SATCheck [104] | SMC | TSO |
-| CDSCheck [222; 223] | SMC | C |
-| Dartagnan [142] | BMC | parametric (Cat language) |
-| GenMC [175] | SMC | parametric |
-| Zord [119] | BMC | TSO, PSO |
-| McSPIN [14] | BMC | parametric |
-| Nidhugg [9; 2] | SMC | TSO, PSO, Power |
-In the weak-memory realm, most model checkers employ either Stateless Model Checking [129] or Bounded Model Checking [83].
+| CBMC extended [Alglave et al., 2013a] | BMC | x86, Power |
+| Herd [Alglave et al., 2021] | simulator | parametric (Cat language) |
+| SATCheck [Demsky and Lam, 2015] | SMC | TSO |
+| CDSCheck [Norris and Demsky, 2013, 2016] | SMC | C |
+| Dartagnan [Haas et al., 2022] | BMC | parametric (Cat language) |
+| GenMC [Kokologiannakis and Vafeiadis, 2021] | SMC | parametric |
+| Zord [Fan et al., 2023] | BMC | TSO, PSO |
+| McSPIN [Abe and Maeda, 2017] | BMC | parametric |
+| Nidhugg [Abdulla et al., 2016, 2017] | SMC | TSO, PSO, Power |
+In the weak-memory realm, most model checkers employ either Stateless Model Checking [Godefroid, 1997] or Bounded Model Checking [Clarke et al., 2001].
 Here, we classify the prominent model checkers into those based on axiomatic formalisms (Section 5), and, less commonly, those based on operational formalisms (Section 6.2).
@@
 Model checking under weak memory models is mostly based on axiomatic formalisms.
-The most prominent axiomatic formalism is the Cat language, with the associated Herd tool [36].
+The most prominent axiomatic formalism is the Cat language, with the associated Herd tool [Alglave et al., 2014b].
 Herd simulates the possible behaviours of assembly programs, under memory models specified in Cat.
 Herd and Cat have been applied to thousands of litmus tests to check whether a specification of a memory model does agree with the hardware behaviour.
-They were used by Arm as the basis for developing the specifications of its memory models [27].
-Herd is included in the “diy tool-suite” [39], which includes facilities for automatically generating litmus tests and running them directly on hardware.
-The CBMC bounded model checker [84] was extended to handle the memory models of x86 and Power [31].
+They were used by Arm as the basis for developing the specifications of its memory models [Alglave et al., 2021].
+Herd is included in the “diy tool-suite” [Arm Ltd., 2024], which includes facilities for automatically generating litmus tests and running them directly on hardware.
+The CBMC bounded model checker [Clarke et al., 2004] was extended to handle the memory models of x86 and Power [Alglave et al., 2013a].
 Apart from litmus tests, the case-studies include components of system software, such as a queue mechanism of the Apache HTTP server, the worker synchronisation mechanism in PostgreSQL, and the read-copy-update mechanism of the Linux kernel.
-Dartagnan [235; 125; 142] is a bounded model checker, where users specify memory models in the Cat language.
-The benchmarks are several mutual exclusion algorithms, including Dekker’s algorithm, Peterson’s algorithm and the qspinlock of the Linux kernel [141; 229].
-Another recent bounded model checker is Zord [119], which is based on a novel axiomatic formalism that supports TSO and PSO, and has been applied to the SV-COMP 2020 benchmarks.
+Dartagnan [Ponce de León et al., 2018; Gavrilenko et al., 2019; Haas et al., 2022] is a bounded model checker, where users specify memory models in the Cat language.
+The benchmarks are several mutual exclusion algorithms, including Dekker’s algorithm, Peterson’s algorithm and the qspinlock of the Linux kernel [Haas et al., 2023; Paolillo et al., 2021].
+Another recent bounded model checker is Zord [Fan et al., 2023], which is based on a novel axiomatic formalism that supports TSO and PSO, and has been applied to the SV-COMP 2020 benchmarks.
+Meanwhile, there are a few stateless model checkers that are based on axiomatic formalisms.
+For instance, CDSCheck [Norris and Demsky, 2013, 2016] is based on the axiomatic C memory model [Batty et al., 2011], and uncovered an error in an implementation of the Chase-Lev double-ended queue.
+There is also the line of stateless model checkers HMC [Kokologiannakis and Vafeiadis, 2020], GenMC [Kokologiannakis and Vafeiadis, 2021] and WMC [Moiseenko et al., 2022].
+They have been applied to synthetic litmus tests, mutual exclusion algorithms, and concurrent data structures (Chase-Lev double-ended queue, Treiber stack, and Michael-Scott queue).
 Based on Operational Formalisms
 There are comparatively fewer model checkers that are based on operational formalisms to encode the weak memory models.
-McSPIN [14] is an extension to the SPIN model checker, and users specify memory models by constraining the operational model in McSPIN.
+McSPIN [Abe and Maeda, 2017] is an extension to the SPIN model checker, and users specify memory models by constraining the operational model in McSPIN.
 The case-studies include litmus tests from the documentation of Unified Parallel C and Itanium, as well as the double-checked locking algorithm and Dekker’s algorithm.
-Nidhugg [9; 2] uses stateless model checking based on a specialised operational formalism of memory models.
+Nidhugg [Abdulla et al., 2016, 2017] uses stateless model checking based on a specialised operational formalism of memory models.
 The memory models supported are TSO, PSO, Power, and a subset of Arm.
@@
 At the heart of model checkers are computational problems for deciding whether a program satisfies certain properties.
-The decidability and computational complexity of these problems are substantial topics in their own right [59].
+The decidability and computational complexity of these problems are substantial topics in their own right [Bouajjani, 2024].
 This section surveys the works that address these problems under weak memory models, with the results summarised in Table 7.
 | Model | Consistency | (data-ind.) | Robustness | Reach. | (param.) |
-| SC | NP-com [128] | FPT [80] | |||
-| Causal | NP-hard [61; 124] | P [61] | |||
-| TSO | NP-hard [124] | FPT [80] | PSPACE-com [62] | dec [42] | PSPACE-com [12] |
-| PSO | NP-hard [124] | FPT [80] | dec [42] | ||
-| RMO | NP-hard [124] | FPT [80] | undec [42] | ||
-| Power | PSPACE-com [105] | ||||
-| RA | FTP [285] | undec [3] | PSPACE-com [176] |
+| SC | NP-com [Gibbons and Korach, 1997] | FPT [Chini and Saivasan, 2020] | |||
+| Causal | NP-hard [Bouajjani et al., 2017; Furbach et al., 2015] | P [Bouajjani et al., 2017] | |||
+| TSO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | PSPACE-com [Bouajjani et al., 2011] | dec [Atig et al., 2010] | PSPACE-com [Abdulla et al., 2020] |
+| PSO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | dec [Atig et al., 2010] | ||
+| RMO | NP-hard [Furbach et al., 2015] | FPT [Chini and Saivasan, 2020] | undec [Atig et al., 2010] | ||
+| Power | PSPACE-com [Derevenetc and Meyer, 2014] | ||||
+| RA | FTP [Tunç et al., 2023] | undec [Abdulla et al., 2019] | PSPACE-com [Krishna et al., 2022] |
 Consistency Checking.
@@
 Later, Cantin et al.
-[72; 73] showed that it is NP-complete to check if a concurrent program yields a coherent execution (Definition 5.10).
+[Cantin et al., 2003, 2005] showed that it is NP-complete to check if a concurrent program yields a coherent execution (Definition 5.10).
 As a corollary, consistency checking is NP-hard under most memory models, since most memory models satisfy the coherence property.
@@
 Also related to model checking (Section 8.1) is the reachability problem, which is to decide whether a given program has an execution from a given starting state to a given final state.
-Under RMO the reachability problem is undecidable, but under TSO and PSO it is decidable, albeit beyond the primitive recursive (PR) class and has a very high computational cost [42].
-In light of this computational hardness, there were works that attempted to trade completeness with efficiency under TSO [60].
+Under RMO the reachability problem is undecidable, but under TSO and PSO it is decidable, albeit beyond the primitive recursive (PR) class and has a very high computational cost [Atig et al., 2010].
+In light of this computational hardness, there were works that attempted to trade completeness with efficiency under TSO [Bouajjani et al., 2015].
 9. Future Directions: Common Properties of Weak Memory Models
@@
 The influential “herding cats” work of Alglave et al.
-[36] robustly and rigorously tests weak memory models against litmus tests.
-This work included testing real hardware across a range of architectures and vendors, and identified a bug in a processor that resulted in a “Programmer Advice Notice” [38], and has been used in the development of Arm processors [27] and the Linux kernel [32].
-Based on this work Arm has led the way in formalising aspects of its architecture [27; 28; 123], including publishing the memory model [196, Chapter B2] and the herdtools suite of applications [39; 23].
+[2014b] robustly and rigorously tests weak memory models against litmus tests.
+This work included testing real hardware across a range of architectures and vendors, and identified a bug in a processor that resulted in a “Programmer Advice Notice” [Arm Ltd., 2011], and has been used in the development of Arm processors [Alglave et al., 2021] and the Linux kernel [Alglave et al., 2018].
+Based on this work Arm has led the way in formalising aspects of its architecture [Alglave et al., 2021, 2024; Fox et al., 2023], including publishing the memory model [Ltd., 2024, Chapter B2] and the herdtools suite of applications [Arm Ltd., 2024; Alglave, 2020].
 Ideally other companies and committees would follow suit, publishing, preferably, a formal model, or at least a representative set of litmus tests (small concurrent programs with an associated reachability condition Section 1.2) that cover both allowed (weak) and forbidden behaviours.
@@
 A “minimal”, maintained set of tests and expected results could be kept by an architecture vendor; programmers and formalists together could access this database, with both small litmus tests and larger programs explaining weak behaviours that can arise.
-This would help formalists check that their definitions and representations are correct, and, with the help of a common notation that abstracts from processor-specific syntax (e.g., [25; 26]), would aid comparison and re-use of results across different architectures and languages.
+This would help formalists check that their definitions and representations are correct, and, with the help of a common notation that abstracts from processor-specific syntax (e.g., [Alglave and Cousot, 2016, 2017]), would aid comparison and re-use of results across different architectures and languages.
 9.2. Formally-Derived Properties
@@
 Programmers want to know the required solution to their programming problem – where to place fences or constraints to maintain the logical integrity of their concurrent algorithm – without needing to search through a database of litmus tests or to understand complex memory model specifications.
-One theoretical aspect upon which the language-design community has more or less reached a consensus is that “programs that contain no data-race should behave as if under Sequential Consistency” [16], a property known as Data-Race-Free implies Sequential Consistency, or the Data Race Freedom Guarantee (DRF).
-This relatively early concept [18; 19] provides a base-level property for programmers of sequential, non-interfering code, similar to the principals underlying the circumstances under which localised pipelined reordering can occur (Section 2.1).
-However, the result does not necessarily help programmers of low-level systems code that does include data races, such as system-wide locks [32; 88].
-What would complement such properties and litmus test conformance is a set of more general properties of programs that can be understood locally [46]; for instance, if a fence is inserted between instructions i and j, can standard ordering be assumed?
+One theoretical aspect upon which the language-design community has more or less reached a consensus is that “programs that contain no data-race should behave as if under Sequential Consistency” [Adve and Boehm, 2010], a property known as Data-Race-Free implies Sequential Consistency, or the Data Race Freedom Guarantee (DRF).
+This relatively early concept [Adve and Hill, 1992, 1993] provides a base-level property for programmers of sequential, non-interfering code, similar to the principals underlying the circumstances under which localised pipelined reordering can occur (Section 2.1).
+However, the result does not necessarily help programmers of low-level systems code that does include data races, such as system-wide locks [Alglave et al., 2018; Colvin et al., 2024].
+What would complement such properties and litmus test conformance is a set of more general properties of programs that can be understood locally [Bargmann and Wehrheim, 2024]; for instance, if a fence is inserted between instructions i and j, can standard ordering be assumed?
 If we let ‘’ represent strict ordering, not subject to memory model effects, then this property of fences can be expressed as follows.
@@
 A general class of properties of programs at this level would make the task of comparing formalisms more straightforward, and provide a relatively stable basis for programmers and verifiers to understand their task.
-Reduction to standard operators also opens up the possibility for existing tools and techniques to be applied for reasoning, reducing the need for specialised inference systems [26; 181; 93; 111; 110; 295], and expanding the range of properties and domains that can be addressed.
+Reduction to standard operators also opens up the possibility for existing tools and techniques to be applied for reasoning, reducing the need for specialised inference systems [Alglave and Cousot, 2017; Lahav and Vafeiadis, 2015; Coughlin et al., 2021; Doherty et al., 2019, 2022; Wright et al., 2021], and expanding the range of properties and domains that can be addressed.
 10. Current and Future Directions: adapting and contributing to software development
@@
 This includes ongoing work into verification methods and tool support, along with model checkers (Sections 7 and 8); in addition, formalisation of weak memory models can and should also impact other aspects of formal concurrency research, such as verified compiler efforts (e.g.
-[259; 284]) and data race detection [20; 166].
+[Ševčík et al., 2013; Trippel et al., 2018]) and data race detection [Adve et al., 1991; Kasikci et al., 2015].
 These topics typically require some abstraction from the specific representation of weak memory to facilitate more standard analyses; we give an overview of some of these below where not previously discussed.
@@
 Based on its original definition (which assumed Sequential Consistency), Travkin et al.
-[280; 281] developed techniques to reduce a TSO program into an SC program, enabling one to utilise existing tools such as the SPIN model checker to test linearisability.
-Other approaches to adapting linearisability for application to weak memory models propose alternative versions of linearisability [107; 106; 269; 268].
+[Travkin et al., 2013; Travkin and Wehrheim, 2014] developed techniques to reduce a TSO program into an SC program, enabling one to utilise existing tools such as the SPIN model checker to test linearisability.
+Other approaches to adapting linearisability for application to weak memory models propose alternative versions of linearisability [Derrick et al., 2014; Derrick and Smith, 2017; Smith et al., 2020; Smith and Groves, 2019].
 Progress/Liveness
@@
 Verifying liveness properties (also called progress properties) is more challenging even in the Sequential Consistency setting, with one reason being the need to address various fairness assumptions.
-In the weak-memory setting, works that deal with liveness properties only started to emerge recently [180; 7; 8].
+In the weak-memory setting, works that deal with liveness properties only started to emerge recently [Lahav et al., 2021; Abdulla et al., 2023, 2024].
 Security Vulnerabilities
 These may emerge from microarchitectural features, including through the exploitation of speculative execution and other mechanisms that cause out-of-order execution.
-The intersection of vulnerabilities and weak memory models is thought to be a potentially important area to investigate [75], and has been addressed by several authors utilising a range of different representations [283; 236; 108; 139; 91; 87; 90].
-The addition of security-focussed instructions, such as Capability Hardware Enhanced RISC Instructions (CHERI), also impacts vulnerability research [299], and will require richer pointer-level models to address in the presence of weak memory.
+The intersection of vulnerabilities and weak memory models is thought to be a potentially important area to investigate [Cauligi et al., 2022], and has been addressed by several authors utilising a range of different representations [Trippel et al., 2019; Ponce de León and Kinder, 2022; Disselkoen et al., 2019; Griffin and Dongol, 2021; Colvin and Winter, 2020; Colvin, 2023; Colvin and Su, 2025].
+The addition of security-focussed instructions, such as Capability Hardware Enhanced RISC Instructions (CHERI), also impacts vulnerability research [Zaliva et al., 2024], and will require richer pointer-level models to address in the presence of weak memory.
 10.2. Adapting Existing Code
@@
 Based on formalisms for defining memory models, much recent effort has extended and applied them to specifying and verifying concurrent libraries.
-These works focus on various memory models, from TSO [67] and Java [118] to fragments of the C memory model [47; 97; 103; 242; 264].
+These works focus on various memory models, from TSO [Burckhardt et al., 2012] and Java [Emmi and Enea, 2019] to fragments of the C memory model [Batty et al., 2013; Dalvandi and Dongol, 2021; Dang et al., 2022; Raad et al., 2019a; Singh and Lahav, 2023].
 Large-scale migration of code between hardware platforms has been recently addressed by Beck et al.
@@
 [2003] empirically applied the same method to the memory models of actual hardware platforms, namely IBM’s Power and Intel’s Pentium, which had different types of fences.
-For TSO and PSO, there are fence insertion algorithms based on modelling the write-buffers [6; 190; 191; 192] and on predicate abstraction techniques [5; 10; 11].
+For TSO and PSO, there are fence insertion algorithms based on modelling the write-buffers [Abdulla et al., 2012b; Linden and Wolper, 2010, 2011, 2013] and on predicate abstraction techniques [Abdulla et al., 2012a, 2015b, 2015a].
 In the case of structured programs, Taheri et al.
 [2019] developed a polynomial-time greedy algorithm for fence insertion, and also demonstrated the NP-hardness of the minimum fence insertion problem with multiple types of fences.
-Many other approaches to fence insertion have been taken [29; 52; 66; 177; 161; 214; 215], with the most recent being [225; 265].
+Many other approaches to fence insertion have been taken [Alglave et al., 2014a; Bender et al., 2015; Burckhardt et al., 2007; Kuperstein et al., 2012; Joshi and Kroening, 2015; Meshman et al., 2014, 2015], with the most recent being [Oberhauser et al., 2021; Singh et al., 2022].
 Finding the minimal set remains an open problem, however, as does finding the most computationally efficient set.
@@
 Transactional Memory
-Transactional memory [140] defines synchronisation mechanisms that allow the programmer to specify atomic “blocks” of instructions.
+Transactional memory [Guerraoui and Kapatka, 2010] defines synchronisation mechanisms that allow the programmer to specify atomic “blocks” of instructions.
 While these blocks should appear atomic, on the hardware level they can still occur in parallel.
 When two atomic blocks attempt to change the same variable and a clash arises, the transactional memory system can safely revert to a consistent state.
-Naturally, transactional memory has interplay with memory consistency models [82; 115].
+Naturally, transactional memory has interplay with memory consistency models [Chong et al., 2018; Dongol et al., 2019].
 Most recently, Dalvandi and Dongol [2022] devised a semantics of transactional memory under the release-acquire fragment of the C memory model, and developed a program logic based on their semantics.
 Persistent Memory
-Persistent memory [131] is a relatively new type of physical memory that maintains its state even after system failures, with the aim of avoiding the loss of data.
+Persistent memory [Gogte et al., 2022] is a relatively new type of physical memory that maintains its state even after system failures, with the aim of avoiding the loss of data.
 Updates to memory are held in the volatile memory first, and they are flushed to the persistent memory in some order, during which the persistent memory may temporarily be in some inconsistent states.
 In the case of a power outage, one needs to ensure that the entire memory system can be reverted back to a consistent state.
-These issues, related to memory consistency and cache coherence, have been addressed by formalisations of weak-memory models [245; 243; 81; 168; 287; 170; 244].
+These issues, related to memory consistency and cache coherence, have been addressed by formalisations of weak-memory models [Raad et al., 2019b, 2020; Cho et al., 2021; Khyzha and Lahav, 2021; Vafeiadi Bila et al., 2022; Klimis et al., 2024; Raad et al., 2022].
 11. Conclusions
@@
 However, these changes, already somewhat exposed to programmers of concurrent code with agressive compiler optimisations, became significant with the advent of multicore architectures, initiating the first wave of attempts at formalising these effects (Section 3).
-Often this issue can be addressed by having programming abstractions between the assembler and the high-level code; however this is not possible for programmers at the interface of hardware and software, for instance, operating systems developers [32].
-In this context safety, security and performance are all of high importance, and a clear place where a rigorous formal model of the behaviour of multicore systems is invaluable [75].
-Since low-level details are almost unavoidable when addressing weak memory model concerns, further refinement of formalisms to include detailed instruction semantics [40] would assist binary-level analysis (e.g., [251; 138]).
+Often this issue can be addressed by having programming abstractions between the assembler and the high-level code; however this is not possible for programmers at the interface of hardware and software, for instance, operating systems developers [Alglave et al., 2018].
+In this context safety, security and performance are all of high importance, and a clear place where a rigorous formal model of the behaviour of multicore systems is invaluable [Cauligi et al., 2022].
+Since low-level details are almost unavoidable when addressing weak memory model concerns, further refinement of formalisms to include detailed instruction semantics [Armstrong et al., 2019] would assist binary-level analysis (e.g., [Rocha et al., 2022; Gouicem et al., 2022]).
 Approaches to formalising weak memory effects range from modelling features of the architectures directly (e.g., Section 4) through to abstracting away from such details to axioms of global traces (the axiomatic approach, Section 5).
@@
 The weakness of the axiomatic style is that it does not lend itself directly to a straightforward semantics for programs.
-While memory models can be compared, and general properties of memory models described [36], one cannot easily take an abstract program and apply a set of (denotational) rules to obtain the set of traces for a program written in some abstract syntax.
-This is addressed by, for instance, the Promising semantics [165], an operational semantics which includes axiomatic-style relations.
+While memory models can be compared, and general properties of memory models described [Alglave et al., 2014b], one cannot easily take an abstract program and apply a set of (denotational) rules to obtain the set of traces for a program written in some abstract syntax.
+This is addressed by, for instance, the Promising semantics [Kang et al., 2017], an operational semantics which includes axiomatic-style relations.
 However this semantics, partly due to capturing a wide-range of weak behaviours, is quite complex to apply to even simple programs.
@@
 https://doi.org/10.1145/3062341.3062352
-- Lamport [1979] Lamport. 1979. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Trans. Comput. C-28, 9, 690–691.
+- Lamport [1979] L. Lamport. 1979. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Trans. Comput. C-28, 9, 690–691.
 - Lee and Padua [2000] J.